### v1

# Dynamical Tidal Response of Schwarzschild Black Holes

Abstract
Dynamical Love numbers capture the conservative response of an object to a time-dependent external tidal gravitational field. We compute the dynamical Love numbers of Schwarzschild black holes in general relativity within a point-particle effective field theory framework. In addition to the known logarithmic running, we compute the finite scheme-dependent contributions to the Love number couplings. We do this by matching the renormalized one-point function in the effective theory to the classical field profile computed in general relativity. On the general relativity side, we solve the Regge–Wheeler and Zerilli equations perturbatively in a small frequency expansion. In order to match on the effective field theory side we include gravitational interactions using the Born series and employ dimensional regularization to obtain a renormalized field profile.
Contents
1 Introduction
We often figure out what is inside things by shaking them. When presented with objects as diverse as a wrapped present, a piggy bank, or a cereal box, our first instinct is give them a shake it to get an idea of what is inside. It should therefore come as no surprise that this is also an effective strategy to learn about the most mysterious objects in nature: black holes. Unfortunately, black holes are too big and too far away for us to grab, so we must be more flexible in what it means to “shake” one. Fortunately, the universe performs a version of this experiment for us. A time-dependent tidal gravitational perturbation effectively makes an object oscillate in response, producing its own gravitational field. The details of this induced gravitational field depend on the make-up of the object, which is encoded in (dynamical) Love numbers [1, 2]. When other heavy objects move past black holes, or orbit them, their changing gravitational field induces a dynamical response in the black hole. This response is physical, and in principle can be measured, for example, in the spectrum of gravitational waves emitted by a binary. Here we compute these dynamical Love numbers of Schwarzschild black holes.
The tidal responses of black holes are well-studied. The leading response of any object to an external tidal gravitational field is to mechanically deform. This response is captured by the object’s static Love number. A remarkable—and somewhat counterintuitive—fact is that the static Love numbers of black holes vanish exactly in (four-dimensional) general relativity [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. From the perspective of effective field theory (EFT), this result implies that, as far as static perturbations are concerned, black holes in general relativity behave like elementary particles, indistinguishable from point-like objects with no internal structure. This emergent simplicity, peculiar to four-dimensional general relativity [6, 7, 15, 16, 17, 18, 19, 20, 21, 22], has been the focus of much interest in recent years, not least because it suggests the presence of new symmetries of general relativity [8, 9, 23, 24, 25, 26, 10, 27, 28, 29, 30, 31]. Given the vanishing of the static linear Love numbers, it is both theoretically interesting and observationally motivated to ask to what extent this hidden simplicity persists at subleading order. Broadly speaking, two classes of subdominant finite-size effects, beyond the linear static tidal response, can be identified: dynamical (i.e., time-dependent) effects and nonlinearities. Nonlinear corrections to the Love numbers have recently been investigated for both black holes and neutron stars. Interestingly, explicit calculations show that the nonlinear Love numbers of black holes continue to vanish in four-spacetime dimensions [14, 32, 33, 34, 35, 27, 36, 37, 38], hinting at a fully nonlinear symmetry structure underlying the static sector of general relativity [27, 38, 29].
In this paper we study subleading in frequency tidal effects, namely the dynamical response. Specifically, we aim to compute the induced time-dependent tidal deformation of Schwarzschild black holes by matching results obtained in general relativity to a point-particle EFT description of black holes [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 6, 7, 13]. This provides an unambiguous definition of the dynamical Love numbers. The general topic of black hole dynamical response is well-studied [49, 50, 51, 32, 13, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69]. In particular, it is well known that frequency-dependent effects are nonzero. For example, a black hole absorbs radiation through its horizon, which, at linear order in frequency, corresponds to a positive imaginary part of the response coefficients [70, 13, 71]. Solutions to the perturbation equations at second order in frequency have also been obtained by various methods. These include the Mano–Suzuki–Takasugi (MST) formalism [72, 73]—which expresses the solution as a series of special functions truncated at a chosen order in the small-frequency limit—as well as more standard perturbative approaches [60, 67]. It is by now well understood that the dynamical Love numbers, unlike their static counterparts, exhibit a logarithmic dependence on distance, which is interpreted as a classical example of renormalization-group running in the point-particle EFT description of the object [49, 13, 56, 59, 60, 67, 74]. However, although the coefficient of the logarithm has been previously computed using the above methods, as well as other symmetry-based arguments [55], existing results in the literature remain incomplete for gravitational perturbations (see, however, [59, 74] for a scalar field tidal matching). In particular, a full computation of the (renormalized) dynamical Love numbers cannot be performed at tree level, but requires evaluating certain (classical) higher-loop diagrams in the point-particle EFT after subtracting ultraviolet divergences.
In this work, we extend previous analyses in several directions. Foremost, we perform (for the first time) the complete matching between general relativity and EFT, so that in addition to the logarithmic running of the gravitational dynamical Love numbers, we compute their scheme-dependent finite terms. Unlike the logarithmic coefficient, which is universal, these terms depend on the chosen renormalization scheme (we will work here in dimensional regularization). This work also contains a number of technical features of interest. The computation of dynamical Love numbers requires solutions to the equations of black hole perturbation theory at subleading orders in the frequency expansion. We develop a systematic expansion in frequency that allows us to relatively easily extract the terms of interest. In order to account for gravitational effects in the EFT, we follow the elegant approach of [75, 74], though we choose to match the graviton one-point function, as opposed to a scattering amplitude. A benefit of matching this off-shell quantity is that the matching can be performed in a suitably defined near zone. This simplifies both the general-relativistic computation and the EFT matching, obviating the need to resolve the far-zone dynamics. For completeness, we also demonstrate in a toy scalar-field example that, when extended to the far zone, our result reproduces the MST solution and agrees with [74].
The most salient feature of our results is that the dynamical Love numbers of Schwarzschild black holes are indeed nonzero. Aside from this, there are some features worth noting. The responses in the parity even and odd sectors are the same (or more properly, there is a renormalization scheme in which they are equal). This can be viewed as a consequence of Chandrasekhar’s symmetry that maps these two sectors into each other [76, 77, 78]. In addition, the responses display intriguing patterns that suggest that there are further insights to be mined from the study of black hole responses.
Outline: In Section 2, as a preliminary example, we analyze dynamical scalar response. While our main focus is on gravitational perturbations, the scalar field case allows us to present the underlying logic in a simpler setup and to highlight the technical differences with respect to previous works. In Section 3, we consider gravitational dynamical tidal effects. First, we solve the Regge–Wheeler and Zerilli equations in a full general-relativistic setup perturbatively in the frequency. We then perform the matching with the point-particle EFT. In the EFT, we employ dimensional regularization and introduce a renormalization scheme to remove ultraviolet divergences. Through this matching, we obtain the renormalized dynamical Love number couplings up to second order in frequency. Several appendices collect complementary and more technical results that are relevant but lie somewhat outside the main line of the text. In Appendix A we derive general expressions for the running of the scalar dynamical response for generic multipole number. In Appendix B, we further analyze the far-zone of the scalar field example, which we use as a crosscheck to show that our approach reproduces the MST solution and previous results. Appendix C contains explicit expressions for the gravitational field solutions and their renormalized counterparts. Finally, Appendix D discusses some technical aspects of the gravitational tidal field.
Conventions: We use the mostly-plus metric convention (,+,,+)(-$\cdots,+$), and denote the spacetime dimension by $D=d+1$, with $d$ the spatial dimension. Spacetime indices are denoted by Greek letters $\mu,\nu,\cdots$, while spatial indices are denoted by Roman letters $i,j,k,\cdots a,b,c,\cdots$. We will often use multi-index notation, where the multi-index $A_{\ell}=i_{1}\cdots i_{\ell}$. The notation $({\cdots})_{T}$ indicates the trace-subtracted symmetrization of the enclosed indices. We (anti)symmetrize indices with weight one, so that for example $A_{[ij]}=\frac{1}{2}(A_{ij}-A_{ji})$. In many cases we decompose fields using spherical harmonics, where we denote the angular momentum by $\ell$, and the magnetic quantum number by $m$.
2 Warm-up: Dynamical scalar response
In order to orient ourselves, we first consider a preliminary example: the dynamical response to an external massless scalar field profile. From the general relativistic perspective, this problem involves solving the Klein–Gordon equation for a massless scalar field propagating in a Schwarzschild spacetime. We begin by explaining how to solve this equation perturbatively in the frequency of the field, obtaining a solution that is valid at all spatial distances. Then, following [74], we use these solutions to match to worldline EFT. The final results of this computation are known. The new contribution of the approach taken here is that: (1) We show that, to perform the EFT matching and obtain the response coefficients, it is not necessary to compute far-field observables; it is enough to know the solution in a suitably defined intermediate zone. (2) By introducing appropriate far regions, we show that our perturbative solution reproduces the MST [72, 73] result, up to linear order in the small-frequency limit.
While this section is useful for setting up notation, its results are not strictly essential for the remainder of the work. Readers already familiar with these ideas may choose to skip directly to Section 3.
2.1 Small-frequency expansion of relativistic solution
We begin by considering the general relativity side. We are interested in the dynamics of a massless scalar field, $\Phi$, on a four-dimensional Schwarzschild spacetime, defined by the line element
${\rm d}s^{2}=-f(r){\rm d}t^{2}+f(r)^{-1}{\rm d}r^{2}+r^{2}({\rm d}\theta^{2}+\sin^{2}\theta\,{\rm d}\varphi^{2})\,,\qquad f(r)\equiv 1-\frac{r_{s}}{r}\,$, (2.1)
where $r_{s}\equiv 2GM$ is the Schwarzschild radius. After decomposing the field in spherical harmonics and in frequency space
$\Phi(t,r,\theta,\varphi)=\sum_{\ell=0}^{\infty}\sum_{m=-\ell}^{\ell}\int\frac{{\rm d}\omega}{2\pi}\operatorname{e}^{-i\omega t}\phi_{\omega\ell m}(r)Y_{\ell m}(\theta,\varphi)\,$, (2.2)
the Klein–Gordon equation $\nabla_{\mu}\nabla^{\mu}\Phi=0$ takes the form
$\partial_{r}\big[r(r-r_{s})\partial_{r}\phi(r)\big]+\left[\frac{r^{3}\omega^{2}}{r-r_{s}}-\ell(\ell+1)\right]\phi(r)=0\,.$ (2.3)
Here we have dropped the $\omega\ell m$ subscript of $\phi(r)$ for convenience, but we will sometimes restore some of these labels when it is useful.
We would like to solve (2.3) at finite frequency $\omega$. In particular, we want to extract the behavior of the solution at asymptotically far distances ($r=\infty$), after imposing appropriate boundary conditions at the black hole horizon ($r=r_{s}$). From the perspective of the theory of linear differential equations, this type of question is equivalent to deriving the connection formulas for the solutions of (2.3). As is well known, at finite frequency, the Klein–Gordon equation (2.3) is of the confluent Heun type: it has regular singularities at $r=0$ and $r=r_{s}$, and an irregular singularity of rank $1$ at $r=\infty$. [^1] The Klein–Gordon equation (2.3) can be brought to the standard form of the confluent Heun equation [79, 80, 81] $\frac{{\rm d}^{2}}{{\rm d}z^{2}}w(z)+\left(\frac{\gamma}{z}+\frac{\delta}{z-1}+\epsilon\right)\frac{{\rm d}}{{\rm d}z}w(z)+\frac{\alpha z-q}{z(z-1)}w(z)=0\,$, (2.4) via the change of variable $\phi(r)=(r-r_{s})^{i\omega r_{s}}\operatorname{e}^{i\omega r}w(z(r))$, with $z\equiv r/r_{s}$. It has parameters $\gamma=1$, $\delta=1+2i\omega r_{s}$, $\epsilon=\alpha=2i\omega r_{s}$ and $q=\ell(\ell+1)$. Connection formulas for Heun equations—or, more generally, for Fuchsian equations with more than three regular singularities—are not known generally in closed form. Explicit analytic expressions can be obtained in certain limits, or perturbatively in some parameter. [^2] See e.g. [82, 83, 84, 85, 86] for recent results in this context. This makes solving the problem at finite frequency difficult. However, in some cases it is sufficient to understand the behavior in the limit $\omega\to 0$. Indeed, it is in this limit that the connection to an effective field theory description of compact objects is most transparent.
We are therefore motivated to work in the adiabatic regime. That is, we will assume $\omega r_{s}\ll 1$ and seek solutions to (2.3) perturbatively in $\omega r_{s}$. It is important to note that, regardless of how small the frequency is, the term $r^{3}\omega^{2}/(r-r_{s})$ in (2.3) is not always guaranteed to be small. For instance, sufficiently close to the horizon ($r\rightarrow r_{s}$), this term dominates and eventually becomes the leading component of the potential. Likewise, this term grows at large distances, as $r\rightarrow\infty$. As such, the small frequency approximation is necessarily more complicated than simply neglecting the $r^{3}\omega^{2}/(r-r_{s})$ term. A natural way to proceed is to adopt an asymptotic expansion approach, reducing the problem to a simplified set of equations by making approximations which are valid in different regions with some overlap. The complete solution can then be reconstructed by matching the various solutions at the boundaries of their respective regions of validity. [^3] This type of approach is sometimes referred to as “boundary layer theory”. See [87] for a more rigorous introduction to the theory of boundary layer problems. See also [88, 89, 90, 91, 92] for a non-exhaustive list of scalar field applications on various black hole spacetimes.
We will define three different zones (following [88]):
  • Near Zone (NZ): Defined by the condition that $r-r_{s}\ll r_{s}$.
  • Intermediate Zone (IZ): Defined by $r_{s}\lesssim r\ll\omega^{-1}$.
  • Far Zone (FZ): Where $r\gg r_{s}$.
In each of these regions, we will be able to neglect certain terms in the original equation (2.3), allowing us to find an approximate, analytic closed-form solution. A central conceptual point of the analysis is that in order to match to point-particle EFT, it suffices to know the solution in the Intermediate Zone, as this has an overlapping regime of validity with the EFT description, if one matches the one-point function. In [74] they match to the EFT in the Far Zone, by matching scattering amplitudes to scattering computed in GR using the MST formalism. In order to make contact with this description, we show in Appendix B that the solution at small frequency reproduces the results of MST when matched to the Far Zone.
We now turn to solving the Klein–Gordon equation in each of these zones and matching across their interfaces.
2.1.1 Near zone
We begin by considering the Near Zone. In the vicinity of the horizon, the potential in (2.3) is dominated by the term $r^{3}\omega^{2}/(r-r_{s})$, and the equation can be approximated by
$\partial_{r}\big[r(r-r_{s})\partial_{r}\phi_{\rm NZ}\big]+\frac{r^{3}\omega^{2}}{r-r_{s}}\phi_{\rm NZ}=0\,.$ (2.5)
Up to corrections of order $(r-r_{s})/r_{s}\ll 1$, the independent solutions of (2.5) are $\operatorname{e}^{\pm i\omega r_{\star}}/r$, where $r_{\star}$ is the tortoise coordinate defined by
$\frac{{\rm d}r_{\star}}{{\rm d}r}\equiv\left(1-\frac{r_{s}}{r}\right)^{-1},\qquad r_{\star}=r+r_{s}\log\left(\frac{r}{r_{s}}-1\right).$ (2.6)
Imposing ingpoing boundary conditions at the horizon, the physical near-zone solution thus reads
$\phi_{\text{NZ}}(r)=B\frac{r_{s}}{r}\mathrm{e}^{-i\omega\left(r+r_{s}\log\left(\frac{r}{r_{s}}-1\right)\right)},$ (2.7)
where $B$ is some arbitrary integration constant.
2.1.2 Intermediate zone
In the Intermediate Zone, defined such that $r_{s}\lesssim r\ll\omega^{-1}$, the term involving $\omega^{2}$ in (2.3) is genuinely small compared to all the other terms in the equation. We can thus treat it perturbatively, and look for a series solution in $\omega$. To enable this, it is convenient to introduce the quantities
$x\equiv\frac{2r}{r_{s}}-1\,,\qquad\epsilon\equiv\omega r_{s}\,$, (2.8)
where $\epsilon$ is the small parameter we will perturb in. In terms of these variables (2.3) takes the form
$(1-x^{2})\partial_{x}^{2}\phi_{\rm IZ}-2x\partial_{x}\phi_{\rm IZ}+\ell(\ell+1)\phi_{\rm IZ}=\epsilon^{2}\frac{(x+1)^{3}}{4(x-1)}\phi_{\rm IZ}\,$, (2.9)
where we have simply moved the term proportional to $\epsilon^{2}$ to the right-hand side, in order to treat it as a source. The idea is to look for a solution of the form
$\phi_{\text{IZ}}=\phi^{(0)}+\epsilon\phi^{(1)}+\epsilon^{2}\phi^{(2)}+\cdots,$ (2.10)
and solve perturbatively in $\epsilon$. At each order, the source on the right hand side can be written in terms of lower-order solutions, allowing us to find a solution using Green’s function methods.
Since the right-hand side of (2.9) starts at order $\epsilon^{2}$, both $\phi^{(0)}$ and $\phi^{(1)}$ solve the homogeneous equation. This equation is Legendre’s differential equation, so the solutions can be written in terms of the Legendre polynomials $P_{\ell}(x)$ and $Q_{\ell}(x)$. [^4] We are using the definition of $Q_{\ell}(x)$ with argument $x>1$. That is, $Q_{\ell}(x)=\frac{1}{2}P_{\ell}(x)\log\left(\frac{x+1}{x-1}\right)-\sum_{n=1}^{\ell}\frac{1}{n}P_{n-1}(x)P_{\ell-n}(x).$ (2.11) See also (A.1) in Appendix A. We can therefore parameterize them as
$\phi_{\ell}^{(0)}(x)$ $=b_{1}^{(0)}P_{\ell}(x)+b_{2}^{(0)}Q_{\ell}(x),$ (2.12)
$\phi_{\ell}^{(1)}(x)$ $=b_{1}^{(1)}P_{\ell}(x)+b_{2}^{(1)}Q_{\ell}(x),$ (2.13)
for some integration constants $b_{1}^{(k)}$ and $b_{2}^{(k)}$. [^5] The constants $b_{1}^{(k)}$ and $b_{2}^{(k)}$ are $\ell$-dependent, as we will explicitly see in some examples below. These can be determined by matching the perturbative solutions $\phi_{\ell}^{(0)}(x)$ and $\phi_{\ell}^{(1)}(x)$ to the near-zone solution (2.7) across the boundary at $r=r_{s}$. To this end, it is useful to recall the asymptotic behavior of the Legendre polynomials near $x=1$ (see e.g., [93, 94]):
$P_{\ell}(x)\underset{x\to 1}{\sim}1\,,\qquad\quad Q_{\ell}(x)\underset{x\to 1}{\sim}-\frac{1}{2}\log\left(\frac{x-1}{2}\right)-H_{\ell}\,$, (2.14)
where $H_{\ell}$ is the harmonic number, $H_{\ell}\equiv\sum_{k=1}^{\ell}\frac{1}{k}=\gamma_{E}+\psi(\ell+1)$, with $\psi(z)\equiv\Gamma^{\prime}(z)/\Gamma(z)$ the digamma function, and $\gamma_{E}$ the Euler–Mascheroni constant.
Let us first compare the $x\rightarrow 1$ limit of the intermediate-zone solution $\phi_{\ell}^{(0)}(x)$ with the near-zone solution (2.7) expanded at zeroth order in $\omega$. By matching the two solutions, one readily finds
$b_{1}^{(0)}=B,\qquad b_{2}^{(0)}=0.$ (2.15)
Proceeding similarly for the $\phi_{\ell}^{(1)}(x)$ solution, and comparing it with $\phi_{\text{NZ}}$ at linear order in $\omega$, we obtain
$b_{1}^{(1)}=iB\left(2H_{\ell}-1\right),\qquad\qquad b_{2}^{(1)}=2iB.$ (2.16)
We now want to find the intermediate-zone solution at second order in the small-$\epsilon$ expansion. Plugging the expansion (2.10) into (2.9), and truncating at order $\epsilon$, we obtain the following inhomogeneous equation for $\phi_{\ell}^{(2)}$:
$(1-x^{2})\partial_{x}^{2}\phi_{\ell}^{(2)}(x)-2x\partial_{x}\phi_{\ell}^{(2)}(x)+\ell(\ell+1)\phi_{\ell}^{(2)}(x)=\frac{(1+x)^{3}}{4(x-1)}\phi_{\ell}^{(0)}(x)\equiv S_{\ell}(x),$ (2.17)
where the source on the right-hand side is fully determined by the zeroth-order solution $\phi_{\ell}^{(0)}$. A general solution to (2.17) can be written as
$\phi_{\ell}^{(2)}(x)=\phi_{\ell,\text{p}}^{(2)}(x)+\phi_{\ell,\text{h}}^{(2)}(x)\,$, (2.18)
where $\phi_{\ell,\text{h}}^{(2)}(x)$ is the most general homogeneous solution to the equation
$\phi_{\ell,\text{h}}^{(2)}(x)=b_{1}^{(2)}P_{\ell}(x)+b_{2}^{(2)}Q_{\ell}(x),$ (2.19)
and where $\phi_{\ell,\text{p}}^{(2)}(x)$ is a particular solution to the equation, which can be expressed as a convolution between the Green’s function of the homogeneous equation and the source, as follows: [^6] We have used that the combination $W\equiv(1-x^{2})[P_{\ell}(x)\partial_{x}Q_{\ell}(x)-Q_{\ell}(x)\partial_{x}P_{\ell}(x)]$, proportional to the Wronskian of the homogeneous equation, is independent of $x$ for all $\ell$’s, i.e. $\partial_{x}W=0$. In particular, it is straightforward to check that $W\equiv 1$.
$\phi_{\ell,\text{p}}^{(2)}(x)=Q_{\ell}(x)\int^{x}_{x_{0}}{\rm d}y\,P_{\ell}(y)S_{\ell}(y)-P_{\ell}(x)\int^{x}_{x_{0}}{\rm d}y\,Q_{\ell}(y)S_{\ell}(y),$ (2.20)
where $x_{0}$ is an arbitrary constant, which can be fixed to any convenient value when computing the integrals. Different choices of $x_{0}$ can be reabsorbed into the coefficients $b_{1}^{(2)}$ and $b_{2}^{(2)}$ of the homogeneous solution. [^7] For this reason one can alternatively leave the lower bounds unspecified and treat the integrals as indefinite. We specify the lower limit to prevent any ambiguity.
All together, the second-order intermediate-zone solution $\phi_{\ell}^{(2)}$ reads
$\phi_{\ell}^{(2)}(x)=\left[b_{1}^{(2)}-\frac{B}{4}\int^{x}_{x_{0}}{\rm d}y\frac{(1+y)^{3}}{y-1}P_{\ell}(y)Q_{\ell}(y)\right]P_{\ell}(x)+\left[b_{2}^{(2)}+\frac{B}{4}\int^{x}_{x_{0}}{\rm d}y\frac{(1+y)^{3}}{y-1}P_{\ell}(y)^{2}\right]Q_{\ell}(x).$ (2.21)
Once again, the integration constants $b_{1}^{(2)}$ and $b_{2}^{(2)}$ can be fixed by matching the intermediate-zone solution $\phi_{\ell}^{(2)}(x)$ with the near-zone solution (2.7) expanded at second order in the frequency. For instance, in the case $\ell=0$, it is straightforward to find (with $x_{0}=2$)
$b_{1,\ell=0}^{(2)}$ $=\frac{B}{24}\left[24\text{Li}_{2}\left(-\frac{1}{2}\right)-27-90\log 3+4\log 2\,(11+\log 8)\right],$ (2.22)
$b_{2,\ell=0}^{(2)}$ $=B\left(\frac{35}{6}-\log 4\right),$
and similarly for higher values of $\ell$.
For sufficiently small $\omega$, the intermediate zone ($r_{s}\lesssim r\ll\omega^{-1}$) extends arbitrarily far from $r_{s}$. For later convenience, it will useful to write down the large-$r$ expansion of this intermediate-zone solution. The solutions are the following, where we keep subleading terms up to $r^{-\ell-1}$ (here we have solved for the $b_{1}^{(2)}$ and $b_{2}^{(2)}$ coefficients for $\ell=1,2$ by matching with the near zone as in (2.22)):
$\phi_{\ell=0}(r)\underset{\frac{r}{r_{s}}\to\infty}{\sim}B$ $+Bi\omega r_{s}\left(\frac{r_{s}}{r}-1\right)$ (2.23)
$+B\omega^{2}r_{s}^{2}\bigg[-\frac{r^{2}}{6r_{s}^{2}}-\frac{5r}{6r_{s}}+\frac{3-\pi^{2}}{6}-\frac{11}{6}\log\frac{r}{r_{s}}+\frac{r_{s}}{r}\left(2+\log\frac{r}{r_{s}}\right)\bigg]+\cdots\,$,
$\phi_{\ell=1}(r)\underset{\frac{r}{r_{s}}\to\infty}{\sim}$ $B\left(\frac{2r}{r_{s}}-1\right)+Bi\omega r_{s}\left(\frac{2r}{r_{s}}-1+\frac{r_{s}^{2}}{6r^{2}}\right)$ (2.24)
$\penalty 10000\ \penalty 10000\ +B\omega^{2}r_{s}^{2}\bigg[-\frac{r^{3}}{5r_{s}^{3}}-\frac{7r^{2}}{10r_{s}^{2}}-\frac{r}{r_{s}}\left(\frac{101+10\pi^{2}}{30}+\frac{19}{15}\log\frac{r}{r_{s}}\right)$
$\penalty 10000\ \penalty 10000\ \hskip 42.67912pt+\frac{113+5\pi^{2}}{30}+\frac{19}{30}\log\frac{r}{r_{s}}+\frac{r_{s}}{2r}+\frac{r_{s}^{2}}{6r^{2}}\left(\frac{11}{6}+\log\frac{r}{r_{s}}\right)\bigg]+\cdots\,$,
$\phi_{\ell=2}(r)\underset{\frac{r}{r_{s}}\to\infty}{\sim}$ $B\left(\frac{6r^{2}}{r_{s}^{2}}-\frac{6r}{r_{s}}+1\right)+Bi\omega r_{s}\left(\frac{12r^{2}}{r_{s}^{2}}-12\frac{r}{r_{s}}+2+\frac{r_{s}^{3}}{30r^{3}}\right)$ (2.25)
$\penalty 10000\ \penalty 10000\ +B\omega^{2}r_{s}^{2}\bigg[-\frac{3r^{4}}{7r_{s}^{4}}-\frac{8r^{3}}{7r_{s}^{3}}-\frac{r^{2}}{r_{s}^{2}}\left(\frac{4133+210\pi^{2}}{210}+\frac{79}{35}\log\frac{r}{r_{s}}\right)$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \hskip 42.67912pt+\frac{r}{r_{s}}\left(\frac{1003+42\pi^{2}}{42}+\frac{79}{35}\log\frac{r}{r_{s}}\right)-\frac{972+35\pi^{2}}{210}$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \hskip 42.67912pt-\frac{79}{210}\log\frac{r}{r_{s}}+\frac{r_{s}}{6r}+\frac{5r_{s}^{2}}{24r^{2}}+\frac{r_{s}^{3}}{30r^{3}}\left(\frac{287}{60}+\log\frac{r}{r_{s}}\right)\bigg]+\cdots\,.$
Some interesting physical observations can be abstracted from these explicit solutions for particular $\ell$ values. First of all, as expected, one notices the standard flat-space falloffs $r^{\ell}$ and $r^{-\ell-1}$ of the Klein–Gordon equation, with $\omega$-dependent coefficients. Technically, this is because—even though we are parametrically distant from $r_{s}$—we are not yet in the asymptotically far region (what we called the Far Zone), where the falloffs (at finite frequency) differ (see Appendix B). In addition, we see logarithmic scaling in $r$ in the decaying $r^{-\ell-1}$ terms. These are well known and can be interpreted as a classical running of the response in the infrared solution [6, 7, 59, 95, 74]. The coefficient of the logarithm remains the same at all distances, and is scheme-independent, so there is a sense in which it is universal. This can be computed in full generality for arbitrary $\ell$ from (2.21), as we show in Appendix A. In contrast, finite constant corrections to effective tidal response couplings will in general depend on the regularization prescription in the EFT, and require an explicit matching calculation. This will be the subject of the next subsection.
To compute the EFT coefficients, we will match to the large-$r$ expansion of the intermediate-zone solutions above. The intermediate zone overlaps with the region of validity of the EFT ($r\gg r_{s}$), so it is not necessary to compute far-zone observables (like scattering amplitudes) for this purpose. Since it will not play a major role in the following, we postpone the discussion of the far zone to Appendix B, where we will in particular demonstrate that it correctly reproduces previous results obtained via different methods.
2.2 EFT calculation of scalar dynamical response
In order to provide an unambiguous definition of the physical responses of compact objects, we will employ point-particle effective field theory. This effective description takes advantage of the fact that any localized object can be approximated as a point particle from sufficiently far away. Its internal composition is encoded in the effective couplings of the particle’s worldline to external fields. This approach has twin advantages of being systematic and unambiguous. Our goal in this Section is to derive the effective couplings of a Schwarzschild black hole to a scalar probe. To do so, we employ the formalism of [74] to treat the effects of coupling the point particle to gravity. One difference in our detailed approach is that we choose to match the off-shell scalar field profile in the intermediate zone between the general relativistic calculation and the EFT description.
The calculation proceeds in three steps. We first derive the response of a point particle to an external scalar probe in the absence of gravity (or nearby to the point particle). We then couple to gravity using the Born series organization of the interactions [75, 74]. This yields a bare solution which must then be renormalized and matched to the general relativistic solutions obtained in Section 2.1. The end output of the calculation are the scalar response coefficients of the black hole.
In order to regulate some ultraviolet divergences in the EFT calculation, it is useful to conduct the calculation in generic dimension. We therefore need the Schwarzschild metric in $D$ spacetime dimensions:
${\rm d}s^{2}=-f(r){\rm d}t^{2}+f(r)^{-1}{\rm d}r^{2}+r^{2}{\rm d}\Omega_{S^{D-2}}^{2},$ (2.26)
where ${\rm d}\Omega_{S^{D-2}}^{2}$ is the line element on the (D2)(D-2)-sphere. [^8] For completeness, the sphere metric in hyperspherical coordinates can be defined recursively using the relation ${\rm d}\Omega_{S^{n}}^{2}={\rm d}\theta_{n}^{2}+\sin^{2}\theta_{n}{\rm d}\Omega_{S^{n-1}}^{2}$, with ${\rm d}\Omega_{S^{1}}^{2}={\rm d}\theta_{1}^{2}$. In this notation, $\theta_{1}\in(0,2\pi)$ while all other angles $\theta_{i}\in(0,\pi)$. The function $f(r)$ is defined as
$f(r)\equiv 1-\left(\frac{r_{s}}{r}\right)^{D-3}=1-\frac{2GMn_{D}\mu^{2\varepsilon}}{r^{D-3}},\qquad{\rm where}\penalty 10000\ \penalty 10000\ n_{D}\equiv\frac{4\pi^{\frac{3-D}{2}}\Gamma\left(\frac{D-1}{2}\right)}{D-2}.$ (2.27)
Here we have used the relation between the Schwarzschild radius $r_{s}$ and the asymptotically flat black hole mass $M$, and we defined $\varepsilon\equiv\frac{4-D}{2}$, which parameterizes the deviation from four dimensions. In (2.27), we introduced a scale $\mu$ with the dimension of energy to ensure that the units of $G$ are independent of $\varepsilon$, and in particular remain the same as in $D=4$ [74]. This is useful as we will eventually take the limit $\varepsilon\rightarrow 0$ ($D\rightarrow 4$). In this limit, the scale $\mu$ is analogous to the arbitrary scale introduced in dimensional regularization of standard quantum field theory calculations with Feynman diagrams.
2.2.1 EFT setup
Famously, things can fall into black holes and not escape. As such, they are dissipative systems, whose detailed microstate we cannot track. Consequently their effective description is an open EFT [96, 97] (see also [98, 99, 100, 101, 102, 103, 104, 105, 19] for recent developments). The point-particle EFT action in the Schwinger–Keldysh approach, including dissipative effects, is [39, 40, 106, 56, 19, 74]
$S=S_{\text{bulk}}+S_{\text{pp}}+S_{\text{int}}.$ (2.28)
The first term, $S_{\text{bulk}}$, is the $D$-dimensional scalar action,
$S_{\text{bulk}}=-\int{\rm d}^{D}x\sqrt{-g}g^{\mu\nu}\partial_{\mu}\Phi_{+}\partial_{\nu}\Phi_{-},$ (2.29)
which describes the scalar field’s dynamics in the bulk spacetime. The second term is the worldline action of the object in the point-particle approximation,
$S_{\text{pp}}=-M\int{\rm d}\tau=-M\int{\rm d}\lambda\sqrt{-g_{\mu\nu}\frac{{\rm d}x^{\mu}}{{\rm d}\lambda}\frac{{\rm d}x^{\nu}}{{\rm d}\lambda}},$ (2.30)
where $x^{\mu}(\lambda)$ parametrizes the location of the point particle as a function of the affine parameter $\lambda$. Finally, we introduce couplings between the scalar and the particle worldline
$S_{\text{int}}=\int{\rm d}\tau\sum_{\ell=0}^{\infty}Q_{I}^{A_{\ell}}(\tau)\nabla^{(\ell)}_{A_{\ell}}\Phi^{I}(x(\tau))\,.$ (2.31)
Here we have defined the traceless combination of derivatives $\nabla^{(\ell)}_{A_{\ell}}\equiv\nabla_{(i_{1}}\cdots\nabla_{i_{\ell})_{T}}$, where $(\cdots)_{T}$ denotes the symmetrized trace-free part of the enclosed indices. [^9] Note that (2.31) is written in terms of spatial indices, and so does not look covariant. This can be rectified by defining the projected covariant derivative $\nabla^{\bot}_{\mu}\equiv{\mathbb{P}^{\nu}}_{\mu}\nabla_{\nu}$, where ${\mathbb{P}^{\nu}}_{\mu}\equiv\delta_{\mu}^{\nu}+u_{\mu}u^{\nu}$ projects onto the plane orthogonal to the four velocity $u^{\mu}\equiv{\rm d}x^{\mu}/{\rm d}\tau$ (where $\tau$ is the proper time, and normalized such that $u^{\mu}u_{\mu}=-1$). The composite operator multipole operator $Q_{I}^{A_{\ell}}(\tau)$ is built from internal degrees of freedom $X_{\pm}$ localized on the worldline of the particle, which physically we can think of as modeling the degrees of freedom into which energy and other charges are dissipating. [^10] Introducing the auxiliary worldline degrees of freedom $X$ is not strictly necessary. One can instead directly write down the effective couplings (2.34) between the $\pm$ fields, following the Schwinger–Keldysh approach [107, 108, 109, 110, 111]. The two approaches are equivalent and lead to the same description of the open system. Correlation functions of the effective operator $Q$ describe how the point particle reacts to the presence of an external field $\Phi$.
To account for dissipative effects, in (2.29) and (2.31) we have introduced two copies of the scalar field $\{\Phi_{1},\Phi_{2}\}$ and defined $\Phi_{+}\equiv\frac{1}{2}(\Phi_{1}+\Phi_{2})$ and $\Phi_{-}\equiv\Phi_{1}-\Phi_{2}$. The indices $I,J$ run over these two copies +,+,- and are contracted with the off-diagonal unit matrix. Each copy lives on a different branch of a two-sided closed-time contour running from $t=$-$\infty$ up to some time of interest and then back to $t=$-$\infty$.
Since we are interested in the response to external $\Phi$ profiles, we can integrate out the internal degrees of freedom $X$ to obtain an effective action
$\operatorname{e}^{\Gamma_{\text{int}}^{\text{in-in}}[\Phi_{\pm}]}=\int\mathcal{D}X_{+}\mathcal{D}X_{-}\operatorname{e}^{iS\left[\Phi_{\pm},X_{\pm}\right]}.$ (2.32)
At leading order, we can do this by replacing $Q$ by its linear response
$\langle Q_{I}^{A_{\ell}}(\tau)\rangle=\int{\rm d}\tau^{\prime}K_{IJ}^{(\Phi)\penalty 10000\ A_{\ell}|B_{\ell^{\prime}}}(\tau-\tau^{\prime})\nabla^{(\ell)}_{B_{\ell^{\prime}}}\Phi^{J}(\tau^{\prime})\,$, (2.33)
where $K_{IJ}^{(\Phi)\,A_{\ell}|B_{\ell^{\prime}}}$ is a Green’s function of the $Q$ degrees of freedom, corresponding to the two-point function of $Q$ in the Keldysh basis via [46, 112, 56, 19]
$\langle Q_{I}^{A_{\ell}}(\tau)Q_{J}^{B_{\ell^{\prime}}}(\tau^{\prime})\rangle=-iK_{IJ}^{(\Phi)\penalty 10000\ A_{\ell}|B_{\ell^{\prime}}}(\tau-\tau^{\prime})\,.$ (2.34)
Making this replacement, the effective action takes the form
$\Gamma^{\text{in-in}}_{\text{int}}=\int{\rm d}\tau_{1}{\rm d}\tau_{2}\sum_{\ell=0}^{\infty}K_{{IJ}}^{(\Phi)\,\ell}(\tau_{2}-\tau_{1})\nabla^{(\ell)}_{A_{\ell}}\Phi^{I}(\tau_{2})\nabla^{(\ell)}{}^{A_{\ell}}{\Phi}^{J}(\tau_{1})\,$, (2.35)
where we have taken advantage of the fact that in the case of interest the system is rotationally symmetric. [^11] This means that the Green’s function (2.34) is diagonal in $\ell,\ell^{\prime}$ and its index structure can be written in terms of $\delta_{ij}$. The end result is that the indices on the two derivative operators acting on $\Phi$ end up getting contracted. Correspondingly the responses at a given $\ell$ are independent of the magnetic quantum number $m$. The function $K^{(\Phi)}$ encodes the response of the particle to external sources. In frequency space, real terms capture conservative responses and imaginary terms capture dissipative response.
2.2.2 Linear response
To derive the linear response of the classical field $\Phi_{+}$ to a background $\mkern 2.5mu\overline{\mkern-2.5mu\Phi\mkern-2.5mu}\mkern 2.5mu_{+}$ induced by the interaction at the location of the worldline (2.35), one has to compute its one-point function, expanding $\operatorname{e}^{i\Gamma_{\text{int}}^{\text{in-in}}}$ at linear order in the perturbation:
$\langle\Phi_{+}(t,\vec{x})\rangle_{\text{in-in}}$ $=\int\mathcal{D}\Phi_{+}\mathcal{D}\Phi_{-}\,\Phi_{+}(t,\vec{x})\operatorname{e}^{i\Gamma_{\text{int}}^{\text{in-in}}[\mkern 2.5mu\overline{\mkern-2.5mu\Phi\mkern-2.5mu}\mkern 2.5mu+\Phi]}$ (2.36)
$=i\sum_{\ell=0}^{\infty}(-1)^{\ell}\int{\rm d}\tau_{1}{\rm d}\tau_{2}\int\frac{{\rm d}{\omega}}{2\pi}\operatorname{e}^{-i\tilde{\omega}(\tau_{2}-\tau_{1})}K^{(\Phi)}_{\ell}({\omega})\nabla^{(\ell)}_{A_{\ell}}\langle\Phi_{+}(t,\vec{x})\Phi_{-}(\tau_{2})\rangle\nabla^{(\ell)}{}^{A_{\ell}}\bar{\Phi}_{+}(\tau_{1})\,$,
where $K_{+-}^{(\Phi)}\equiv K^{(\Phi)}$ related to the retarded (thus causal) Green’s function via
$K_{+-}^{(\Phi)}(\tau_{2}-\tau_{1})=-G_{R}^{(Q)}(\tau_{2}-\tau_{1})\equiv i\langle\left[Q_{+}(\tau_{2}),Q_{-}(\tau_{1})\right]\rangle\theta(\tau_{2}-\tau_{1}).$ (2.37)
The two-point function of the scalar field appearing in (2.36) is related to the field retarded Green’s function by
$\langle\Phi_{+}(t,\vec{x})\Phi_{-}(\tau_{2},0)\rangle=iG^{(\Phi)}_{\text{R}}(t-\tau_{2},\vec{x}).$ (2.38)
To further simplify (2.36), we first note that the fact that the scalar tidal response coefficients of black holes vanish implies that $K_{\ell}^{(\Phi)}(\omega)$ starts at linear order in $\omega$ (which can be checked a posteriori). We can therefore neglect order $O(\omega^{2})$ terms coming from $\nabla^{(\ell)}_{i_{L}}\langle\Phi^{+}(t,\vec{x})\Phi^{-}(\tau_{2})\rangle\nabla^{(\ell)}{}^{i_{L}}\bar{\Phi}^{+}(\tau_{1})$, meaning that it will be sufficient to treat the propagation of the scalar as instantaneous, so that
$iG^{(\Phi)}_{\text{R}}(t-\tau_{2},\vec{x})=-i\delta(t-\tau_{2})\int\frac{{\rm d}^{D-1}\vec{p}}{(2\pi)^{D-1}}\frac{\operatorname{e}^{i\vec{p}\cdot\vec{x}}}{\vec{p}^{2}}\,.$ (2.39)
We additionally need the explicit form of the tidal field $\bar{\Phi}_{+}$. Solving the Klein–Gordon equation in flat spacetime, $\square\bar{\Phi}_{+}=0$, one can write $\bar{\Phi}_{+}$ in cartesian coordinates as a Taylor expansion around $\vec{x}=0$:
$\mkern 2.5mu\overline{\mkern-2.5mu\Phi\mkern-2.5mu}\mkern 2.5mu_{+}(\tau_{1})=e^{-i\omega\tau_{1}}\sum_{\ell}c_{j_{1}\cdots j_{\ell}}x^{j_{1}}\cdots x^{j_{\ell}}+O(\omega^{2})\,$, (2.40)
with $c_{j_{1}\cdots j_{\ell}}$ traceless and fully symmetric. Using this and (2.39) in (2.36), we find
$\langle\Phi_{+}(t,\vec{x})\rangle_{\text{in-in}}=\operatorname{e}^{-i\omega t}\sum_{\ell=0}^{\infty}(-i)^{\ell}\ell!\,K^{(\Phi)}_{\ell}(\omega)c^{i_{1}\cdots i_{\ell}}\int\frac{{\rm d}^{D-1}\vec{p}}{(2\pi)^{D-1}}\operatorname{e}^{i\vec{p}\cdot\vec{x}}\frac{p_{i_{1}}\cdots p_{i_{\ell}}}{\vec{p}^{2}}\,.$ (2.41)
Fourier transforming back to position space, this is [^12] We use the Fourier integral $i^{L}\int\frac{{\rm d}^{d}\vec{p}}{(2\pi)^{d}}\operatorname{e}^{i\vec{p}\cdot\vec{x}}\,\frac{p^{(i_{1}}\cdots p^{i_{L})_{T}}}{\vec{p}^{2}}=\frac{\Gamma(\tfrac{d}{2}-1)\Gamma(2-\tfrac{d}{2})}{2^{L}(4\pi)^{d/2}\Gamma(2-\tfrac{d}{2}-L)}x^{(i_{1}}\cdots x^{i_{L})_{T}}\left(\frac{\vec{x}^{2}}{4}\right)^{1-\tfrac{d}{2}-L}\,$, (2.42) to convert between momentum and position space.
$\langle\Phi_{+}(t,\vec{x})\rangle_{\text{in-in}}=\operatorname{e}^{-i\omega t}\sum_{\ell=0}^{\infty}(-1)^{\ell}K^{(\Phi)}_{\ell}(\omega)\frac{2^{\ell-2}\ell!\,\Gamma\left(\frac{D-3}{2}\right)\Gamma\left(\frac{5-D}{2}\right)}{\pi^{\frac{D-1}{2}}\Gamma\left(\frac{5-D}{2}-\ell\right)}c_{i_{1}\cdots i_{\ell}}\frac{x^{i_{1}}\cdots x^{i_{\ell}}}{|\vec{x}|^{2\ell+D-3}}\,.$ (2.43)
Simplifying the ratio of gamma functions and writing the tidal field in spherical coordinates as $c_{i_{1}\cdots i_{\ell}}x^{i_{1}}\cdots x^{i_{\ell}}=c_{\text{ext}}r^{\ell}Y_{\ell}^{m}$, we obtain
$\langle\Phi_{+}\rangle_{\text{in-in}}=\operatorname{e}^{-i\omega t}\sum_{\ell=0}^{\infty}K^{(\Phi)}_{\ell}(\omega)\frac{2^{\ell-2}\ell!\,\Gamma\left(\ell+\frac{D-3}{2}\right)}{\pi^{\frac{D-1}{2}}}\frac{c_{\text{ext}}Y_{\ell}^{m}}{r^{\ell+D-3}}\,.$ (2.44)
Combining this with the external field and replacing $K_{\ell}^{(\Phi)}$ by $\mu^{2\varepsilon}K_{\ell}^{(\Phi)}$ (where $\mu$ is a dimensionful constant in order to keep the response function dimension independent) yields
$\mkern 2.5mu\overline{\mkern-2.5mu\Phi\mkern-2.5mu}\mkern 2.5mu_{+}+\langle\Phi_{+}\rangle_{\text{in-in}}=c_{\text{ext}}\operatorname{e}^{-i\omega t}\sum_{\ell=0}^{\infty}Y_{\ell}^{m}\left(r^{\ell}+\mu^{2\varepsilon}K^{(\Phi)}_{\ell}(\omega)\frac{2^{\ell-2}\ell!\Gamma\left(\ell+\frac{D-3}{2}\right)}{\pi^{\frac{D-1}{2}}}r^{-\ell-D+3}\right).$ (2.45)
This result gives the scalar field response, generated by the interaction term of the action $S_{\text{int}}$, in the vicinity of the point particle. However, in order to match to the UV result, we have to account for the effects of gravity, which we have so far neglected in the EFT.
2.2.3 Coupling to gravity via the Born series
The result (2.45) represents the leading response of the point particle in the $G_{\rm N}\to 0$ limit. However, in order to make contact with the UV solution at subleading order in $G_{\rm N}$ we need to include gravitational effects. To do this, we follow the nice approach of [75, 74] and utilize the Born series to capture nonlinear $G_{\rm N}$ corrections to the Minkowski spacetime solution.
The philosophy is effectively the following: we consider the full bulk equation of motion following from (2.29) (including gravity), which can be written as
$\vec{\nabla}^{2}\Phi_{+}=\left(V_{G}-\omega^{2}\right)\Phi_{+},$ (2.46)
where in spherical coordinates the potential reads
$V_{G}=\frac{2G_{\text{N}}Mn_{D}\mu^{2\varepsilon}}{r^{1-2\varepsilon}}\left[\frac{{\rm d}^{2}}{{\rm d}r^{2}}+\frac{1}{r}\frac{{\rm d}}{{\rm d}r}-\omega^{2}\sum_{n=0}^{\infty}\left(\frac{2G_{\text{N}}Mn_{D}\mu^{2\varepsilon}}{r^{1-2\varepsilon}}\right)^{n}\right].$ (2.47)
We can then imagine solving (2.46) perturbatively in $G_{\rm N}$ and $\omega$. Since we are truncating the $G_{\rm N}$ expansion, the solutions will generically be divergent as $r\to 0$. We then view the source of this divergence as the point particle itself. The properties of the long-distance $\Phi$ solution will then be determined by $K(\omega)$ in (2.45). This will relate the worldline couplings to the parameters of the $\Phi$ solution at large distances. We can then match these solutions (which include gravity) to the GR calculations from Section 2.1 to determine $K(\omega)$, and hence the worldline couplings.
As a practical matter, at each order, (2.46) takes the form of an inhomogeneous second-order differential equation. The most general solution is given by a sum of homogeneous and particular solutions. The homogeneous solution has the same form at all orders in $G_{\text{N}}$ and $\omega$. We will therefore solve it once and allow the integration constants to be generic functions of $G_{\text{N}}$ and $\omega$. We will then impose boundary conditions at the point particle’s location and plug this solution in the source on the right-hand side to obtain the particular solution. In the following, we briefly review the steps of the procedure in the scalar field case, previously discussed in [74], which we will later extend to gravitational perturbations in Section 3.
Homogeneous solution and boundary conditions
The homogeneous part of (2.46) is simply Laplace’s equation for a static scalar field in $D$ dimensions
$\vec{\nabla}^{2}\Phi_{+}=0.$ (2.48)
If we decompose the field in higher-dimensional spherical harmonics as [^13] We collectively denote by $\vec{\theta}$ the angular coordinates of the hyperspherical harmonics $Y_{\ell}^{m}(\vec{\theta})$ on the $S^{D-2}$ sphere, and, with a slight abuse of notation, use $m$ as a multi-index to label all the magnetic quantum numbers. See, e.g., Appendix A of [7] for details. Note that the transformed field $\varphi$ in (2.49) carries implicit $\ell,m$ labels that we suppress for notational simplicity.
$\Phi_{+}=\sum_{\ell,m}\varphi(r)\,r^{\frac{2-D}{2}}Y_{\ell}^{m}(\vec{\theta})\,$, (2.49)
then (2.49) takes the form
$\left(\frac{{\rm d}^{2}}{{\rm d}r^{2}}-\frac{(\ell-\varepsilon)(\ell-\varepsilon+1)}{r^{2}}\right)\varphi^{(h)}(r)=0.$ (2.50)
The general solution to this equation is
$\varphi^{(h)}(r)=B_{\text{reg}}r^{\ell+1-\varepsilon}+\frac{\mu^{2\varepsilon}B_{\text{irr}}}{2\ell+1-2\varepsilon}r^{-\ell+\varepsilon}\,$, (2.51)
with $B_{\text{reg}}$ and $B_{\text{irr}}$ the two arbitrary integration constants. [^14] For convenience, we extracted the constant prefactor $\frac{\mu^{2\epsilon}}{2\ell+1-2\epsilon}$ from the decaying branch. Implicitly, the $r$-independent coefficients $B_{\text{reg}}$ and $B_{\text{irr}}$ should be thought of as expanded in powers of $G$ and $\omega$, as we discussed, and will see explicitly later on. The former branch of solution is regular everywhere in space, including the origin, with $B_{\text{reg}}$ corresponding to the amplitude of the asymptotic external tidal field that we use to probe the point object. On the other hand, the decaying $r^{-\ell+\varepsilon}$ solution is divergent at the location of the particle. We fix its free coefficient $B_{\text{irr}}$ by demanding that the divergence is sourced precisely by the localized action term $S_{\text{int}}$ (2.31). In practice, we determine the Love numbers $K_{\ell}^{(\Phi)}(\omega)$ by matching the homogeneous solution for $\Phi$
$\Phi_{+}^{(h)}=\operatorname{e}^{-i\omega t}\sum_{\ell=0}^{\infty}Y_{\ell}^{m}\,B_{\rm reg}\left(r^{\ell}+\frac{B_{\rm irr}}{B_{\rm reg}}\frac{\mu^{2\epsilon}}{2\ell+1-2\varepsilon}r^{-\ell-1+2\varepsilon}\right),$ (2.52)
to the one-point function (2.45). This yields
$\frac{2^{\ell-1}\ell!\Gamma\left(\ell+\frac{D-1}{2}\right)}{\pi^{\frac{D-1}{2}}}K^{(\Phi)}_{\ell}(\omega)=\frac{B_{\text{irr}}}{B_{\text{reg}}}\,.$ (2.53)
As per [74], this implies that the ratio of irregular to regular solution is the worldline tidal response. We next include the particular solution to the equation including the gravitational potential.
Particular solution
The solution (2.52) with (2.53) is the homogeneous solution to (2.46) with the correct boundary conditions at the location of the point particle. We now turn to deriving the particular solution in the presence of the $V_{G}$ potential. To this end, we consider points away from the origin and solve
$\vec{\nabla}^{2}\Phi_{+}=\left(V_{G}-\omega^{2}\right)\Phi_{+}.$ (2.54)
In terms of the variable $\varphi(r)$ defined in (2.49), this equation becomes
$\left(\frac{{\rm d}^{2}}{{\rm d}r^{2}}-\frac{(\ell-\varepsilon)(\ell-\varepsilon+1)}{r^{2}}\right)\varphi(r)=V_{\varphi}(r)\varphi(r),$ (2.55)
with potential
$V_{\varphi}\equiv\sum_{n=1}^{\infty}\left(\frac{2G_{\text{N}}Mn_{D}\mu^{2\varepsilon}}{r^{1-2\varepsilon}}\right)^{n}\left[\frac{2\varepsilon-1}{r}\frac{{\rm d}}{{\rm d}r}+\frac{\ell^{2}+\ell+1-\varepsilon(3+2\ell)+2\varepsilon^{2}}{r^{2}}-(n+1)\omega^{2}\right]-\omega^{2}.$ (2.56)
We can solve (2.55) formally perturbatively in $G_{\text{N}}$ and $\omega$ via the Born series [74]:
$\varphi(r)=\varphi^{(h)}(r)$ $+\int^{r}{\rm d}r^{\prime}\,G(r,r^{\prime})V_{\varphi}(r^{\prime})\,\varphi^{(h)}(r^{\prime})$ (2.57)
$+\int^{r}{\rm d}r^{\prime}\,G(r,r^{\prime})V_{\varphi}(r^{\prime})\int^{r^{\prime}}{\rm d}r^{\prime\prime}\,G(r^{\prime},r^{\prime\prime})V_{\varphi}(r^{\prime\prime})\,\varphi^{(h)}(r^{\prime\prime})+\cdots.$
In (2.57), $\varphi^{(h)}$ is the homogeneous solution (2.51), with $B_{\text{reg}}$ and $B_{\text{irr}}$ satisfying (2.53), while $G(r,r^{\prime})$ is a Green’s function solving
$\left(\frac{{\rm d}^{2}}{{\rm d}r^{2}}-\frac{(\ell-\varepsilon)(\ell-\varepsilon+1)}{r^{2}}\right)G(r,r^{\prime})=\delta(r-r^{\prime})\,.$ (2.58)
In order to ensure that the particular solution does not alter the boundary condition (2.53) of the homogeneous solution, we will require the Green’s function to be proportional to the theta function $\theta(r-r^{\prime})$, so that it vanishes as $r\rightarrow 0$. From the continuity of $G(r,r^{\prime})$ and the jump condition of $\partial_{r}G(r,r^{\prime})$ at $r=r^{\prime}$, it follows that [74]
$G(r,r^{\prime})=\frac{r^{\varepsilon-\ell}(r^{\prime})^{1+\ell-\varepsilon}-(r^{\prime})^{\varepsilon-\ell}r^{1+\ell-\varepsilon}}{2\varepsilon-2\ell-1}\theta(r-r^{\prime})\,.$ (2.59)
For the same reason, all integrals in the perturbative series (2.57) can be treated as indefinite integrals; in other words, we omit the lower integration bound, which would otherwise modify the homogeneous solution $\varphi^{(h)}(r)$.
To illustrate, here we record the solutions for the $\ell=0,2$ multipoles [74]. Keeping only the terms scaling as $r^{\ell}$ and $r^{-\ell-1}$, and taking the $\varepsilon\rightarrow 0$ limit, one has
$\Phi_{\ell=0}(r)\equiv\varphi_{\ell=0}(r)r^{\frac{2-D}{2}}=B_{\rm reg}\bigg(1$ $-\frac{11\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{6\varepsilon}+\frac{16\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{9}-\frac{22}{3}\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}\log(\mu r)\bigg)$
$+B_{\text{irr}}\left(-\frac{\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu\omega^{2}}{\varepsilon}-\frac{\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu\omega^{2}}{2}-4\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu\omega^{2}\log(\mu r)\right)$ (2.60)
$+\frac{B_{\rm reg}}{r}\left(\frac{2\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{3}\omega^{2}}{\varepsilon}+\frac{80\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{3}\omega^{2}}{3}+12\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{3}\omega^{2}\log(\mu r)\right)$
$+\frac{B_{\text{irr}}}{r}\left(1+\frac{11\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{6\varepsilon}+\frac{68}{3}\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}+11\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}\log(\mu r)\right)+\cdots,$
where we have adopted the convention of [74], by defining [^15] This is just a matter of choice of renormalization scheme. Choosing to work with $G_{\text{N}}$, instead of $\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu$}, would result in different finite $O(\varepsilon^{0})$ terms in the expanded solutions.
$\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu\equiv G_{\text{N}}Mn_{D}.$ (2.61)
As expected, there are UV divergences in the $\varepsilon\rightarrow 0$ limit, as well as logarithmic terms in addition to the standard polynomials in rr. [^16] The divergent $\frac{1}{\varepsilon}$ terms and the $\log(r)$ result from the integrals $\int^{r}{\rm d}r^{\prime}(r^{\prime})^{-1+a\varepsilon}=\frac{r^{a\varepsilon}}{a\varepsilon}\underset{\varepsilon\to 0}{\sim}\frac{1}{a\varepsilon}+\log(r)+O(\varepsilon)$. Similarly, for the quadrupole we have
$\Phi_{\ell=2}(r)=\,B_{\rm reg}r^{2}\bigg(1$ $-\frac{79\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{210\varepsilon}+\frac{11429\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{11025}-\frac{158}{105}\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}\log(\mu r)\bigg)$ (2.62)
$+\frac{B_{\rm reg}}{r^{3}}\left(\frac{8\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{7}\omega^{2}}{135\varepsilon}+\frac{9328\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{7}\omega^{2}}{2025}+\frac{112}{135}\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{7}\omega^{2}\log(\mu r)\right)$
$+\frac{B_{\text{irr}}}{r^{3}}\left(\frac{1}{5}+\frac{79\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{1050\varepsilon}+\frac{39416\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{55125}+\frac{79}{175}\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}\log(\mu r)\right)+\dots\,$,
which displays the same characteristic features.
2.2.4 Renormalization and matching
The infinities appearing in (2.60) and (2.62) as $\varepsilon\rightarrow 0$ are reminiscent of UV divergences in perturbative calculations in quantum field theory (QFT)—with the difference that here we are dealing with classical worldline loops, rather than quantum loops [39, 6, 113, 114, 59, 95]. In QFT, one introduces a regularization scheme and a set of renormalization conditions that allow the identification of the physical parameters—those measured in experiments—and their relation to the bare parameters appearing in the Lagrangian. Counterterms are then introduced into the Lagrangian to remove the divergences and enforce the renormalization conditions. A byproduct of this perturbative renormalization procedure is the appearance of scale dependence in the coefficients and logarithmic running governed by precise renormalization group equations. In the present context, the running parameters are $B_{\text{reg}}$ (the scalar field amplitude) and $B_{\text{irr}}$ (or, equivalently, the tidal response couplings $K_{\ell}^{(\Phi)}(\omega)$ via (2.53)). We can relate the renormalized coefficients $\bar{B}_{\text{reg}}$ and $\bar{B}_{\text{irr}}$ to the bare ones via the relations
$B_{\text{reg}}$ $=\bar{B}_{\text{reg}}(1+\omega^{2}\delta_{11})+\bar{B}_{\text{irr}}\omega^{2}\delta_{12}\,$, (2.63)
$B_{\text{irr}}$ $=\bar{B}_{\text{irr}}(1+\omega^{2}\delta_{22})+\bar{B}_{\text{reg}}\omega^{2}\delta_{21}\,.$
Defining the matrix
$\delta=\begin{pmatrix}\delta_{11}&\delta_{12}\\ \delta_{21}&\delta_{22}\end{pmatrix}$, (2.64)
one finds the following values are needed to cancel the $1/\varepsilon$ divergences [74]:
$\delta^{\ell=0}=\begin{pmatrix}\frac{11\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}}{6\varepsilon}&\frac{\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu}{\varepsilon}\\ -\frac{2\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{3}}{\varepsilon}&-\frac{11\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}}{6\varepsilon}\end{pmatrix}\qquad\qquad\delta^{\ell=2}=\begin{pmatrix}\frac{79\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}}{210\varepsilon}&0\\ -\frac{\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{7}}{\varepsilon}&-\frac{79\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}}{210\varepsilon}\end{pmatrix}$. (2.65)
Plugging these back into the scalar field solution (2.57) now yields a finite limit as $\varepsilon\rightarrow 0$: [^17] We stress the substitution should be done not in (2.60) and (2.62), but in $\Phi(r)$ before taking the $\varepsilon\to 0$ limit.
$\Phi^{\text{R}}_{\ell=0}=$ $\,\bar{B}_{\text{reg}}\left[1+\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}\left(\frac{16}{9}-\frac{22}{3}\log(\mu r)\right)+\frac{\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{3}\omega^{2}}{r}\left(\frac{68}{3}+8\log(\mu r)\right)\right]$ (2.66)
$+\bar{B}_{\text{irr}}\left[\frac{1}{r}-\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu\omega^{2}\left(\frac{1}{2}+4\log(\mu r)\right)+\frac{\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{r}\left(19+\frac{22}{3}\log(\mu r)\right)\right]+\cdots\,$,
and for $\ell=2$:
$\Phi^{\text{R}}_{\ell=2}=$ $\,\bar{B}_{\text{reg}}\left[r^{2}+\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}\left(\frac{11429}{11025}-\frac{158}{105}\log(\mu r)\right)+\frac{\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{7}\omega^{2}}{r^{3}}\left(\frac{63064}{14175}+\frac{32}{45}\log(\mu r)\right)\right]$ (2.67)
$+\bar{B}_{\text{irr}}\left[\frac{1}{5r^{3}}+\frac{\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{r^{3}}\left(\frac{37757}{55125}+\frac{158}{525}\log(\mu r)\right)\right]+\cdots\,$,
where $\Phi^{\text{R}}(r)$ is the renormalized scalar field solution. As a consistency check, it is worth noting that the conditions (2.65) are enough to remove all divergences from the renormalized solution $\Phi^{\text{R}}(r)$, not only the ones in the coefficients of the $r^{\ell}$ and $r^{-\ell-1}$ fall-offs, displayed in (2.66) and (2.67).
These renormalized solutions are directly the EFT analogues of the scalar field solutions obtained in Section 2.1. Comparing (2.66) and (2.67) with the full general relativistic solutions (2.23)–(2.25), we can match and obtain the EFT parameters:
$\bar{B}_{\text{reg}}^{\ell=0}=$ $B\bigg[1-i\omega r_{s}+\omega^{2}r_{s}^{2}\left(\frac{1}{18}-\frac{\pi^{2}}{6}+\frac{11}{6}\log(\mu r_{s})\right)\bigg]$ (2.68)
$\bar{B}_{\text{irr}}^{\ell=0}=$ $B\bigg[i\omega r_{s}^{2}-\omega^{2}r_{s}^{3}\left(\frac{5}{6}+\log(\mu r_{s})\right)\bigg]$
$\bar{B}_{\text{reg}}^{\ell=1}=$ $B\bigg[2+2i\omega r_{s}+\omega^{2}r_{s}^{2}\left(-\frac{1733+150\pi^{2}}{450}+\frac{19}{15}\log(\mu r_{s})\right)\bigg]$
$\bar{B}_{\text{irr}}^{\ell=1}=$ $B\bigg[\frac{i\omega r_{s}^{4}}{2}-\omega^{2}r_{s}^{5}\left(\frac{39}{80}+\frac{1}{2}\log(\mu r_{s})\right)\bigg]$
[…]
$\bar{B}_{\text{irr}}^{\ell=2}=$ $B\bigg[\frac{i\omega r_{s}^{6}}{6}-\omega^{2}r_{s}^{7}\left(\frac{232}{945}+\frac{1}{6}\log(\mu r_{s})\right)\bigg]\,$,
where we have also included the results for $\ell=1$, though we did not display the intermediate steps. Note that for $\ell=2$, redefined $B\to Br_{s}^{2}$, so that $B_{\text{reg}}$ is a polynomial in $r_{s}$. Plugging (2.68) into (2.53), we can finally deduce the $\ell=0,1,2$ scalar response coefficients:
$\frac{1}{4\pi}K^{(\Phi)}_{\ell=0}(\omega)$ $=i\omega r_{s}^{2}+\omega^{2}r_{s}^{3}\left(-\frac{11}{6}-\log(\mu r_{s})\right),$ (2.69) $\frac{3}{\pi}K^{(\Phi)}_{\ell=1}(\omega)$ $=i\omega r_{s}^{4}+\omega^{2}r_{s}^{5}\left(\frac{1}{40}-\log(\mu r_{s})\right),$ $\frac{270}{\pi}K^{(\Phi)}_{\ell=2}(\omega)$ $=i\omega r_{s}^{6}+\omega^{2}r_{s}^{7}\left(\frac{166}{315}-\log(\mu r_{s})\right)\,$,
which agree with [59, 74]. Notice that the responses include an imaginary part (corresponding to dissipation) at leading order in the frequency, and a conservative (real) part at order $\omega^{2}$, which both has a logarithmic running component, and a particular finite contribution. As a matter of principle there is no obstruction to carrying out this procedure to subleading orders in either $G_{\rm N}$ or $\omega$, or for higher multipoles. However, we now turn to the analogous computation of gravitational tidal responses (Love numbers).
3 Dynamical Love numbers
The responses of objects to tidal gravitational perturbations are encoded in the Love numbers of the object. Here we compute the dynamical Love numbers of a Schwarzschild black hole.
The dynamics of perturbations of Schwarzschild black holes in general relativity is described by the Regge–Wheeler [115] and Zerilli [116] equations. Similar to the Schwarzschild Klein–Gordon equation for a massless scalar, both the Regge–Wheeler and Zerilli equations feature an irregular singular point at $r=\infty$, and can be recast into the standard form of the confluent Heun equation. [^18] In the way it is commonly written down [116] (see eq. (3.3) with $D=4$), the Zerilli equation exhibits an additional regular singular point at a negative value of $r$. This singularity can, however, be eliminated through a suitable field redefinition. Indeed, it is well known that the Zerilli equation can be mapped to the Regge–Wheeler equation via the Chandrasekhar transformation [76, 77, 78]. For the boundary value problem of interest here—where the field asymptotically approaches a tidal-field profile at large distances—small-frequency corrections to the static solution have been computed using a range of semi-analytical techniques, approximation schemes, and symmetry arguments (see, e.g. [49, 32, 13, 117, 83, 56, 55, 54, 53, 60, 63, 118, 65, 61, 67]).
We first construct perturbative-in-frequency solutions to the Regge–Wheeler and Zerilli equations, paralleling the scalar analysis of Section 2 (see also [60, 67] for related treatments). We then match these solutions to the worldline EFT in dimensional regularization, and explicitly compute the dynamical Love number coefficients, including, for the first time, the finite scheme-dependent terms.
3.1 General relativistic solution via small-frequency expansion
We begin by considering the general relativistic computation of tidal responses. We first set up notation and briefly recall the derivation of the Regge–Wheeler and Zerilli equations. For later convenience, we present these expressions in general $D$ dimensions. In the general relativistic part of our calculation we will exclusively work in $D=4$, but keeping the equations in generic $D$ will be important for implementing dimensional regularization within the EFT framework.
We are interested in the dynamics of perturbations around a background Schwarzschild geometry (2.26). Denoting it with $\mkern 2.5mu\overline{\mkern-2.5mug\mkern-2.5mu}\mkern 2.5mu_{\mu\nu}$, we perturb the metric as ${g}_{\mu\nu}=\mkern 2.5mu\overline{\mkern-2.5mug\mkern-2.5mu}\mkern 2.5mu_{\mu\nu}+h_{\mu\nu}$. To make maximal use of the ${\rm SO}(D-1)$ background symmetry, it is convenient to decompose $h_{\mu\nu}$ into scalar, vector, and tensor spherical harmonics. Concretely, we can write $h_{\mu\nu}$ as follows [119, 120, 121, 7, 122]:
[…]
$+\sum_{\ell,m}\begin{pmatrix}0&0&h_{0}(t,r)Y_{i}^{(T)}{}_{\ell}^{m}\\ *&0&h_{1}(t,r)Y_{i}^{(T)}{}_{\ell}^{m}\\ *&*&r^{2}h_{2}(t,r)\nabla_{(i}Y_{j)_{T}}^{(T)}{}_{\ell}^{m}\end{pmatrix}+\sum_{\ell,m}\begin{pmatrix}0&0&0\\ *&0&0\\ *&*&r^{2}h_{T}(t,r)\end{pmatrix}Y^{(TT)}_{ij}{}_{\ell}^{m}\,$,
where the entries denoted by * are the same as the entries across the diagonal because $h_{\mu\nu}$ is symmetric. In (3.1), $Y_{\ell}^{m}$ are scalar spherical harmonics, $Y_{i}^{(T)}{}_{\ell}^{m}$ are (transverse) vector harmonics, and $Y_{ij}^{(TT)}{}_{\ell}^{m}$ are (transverse and traceless) tensor harmonics. They are orthogonal to each other and satisfy standard eigenvalue equations of the Laplace operator on the $S^{D-2}$ sphere (see [7] for details). The term proportional to the scalar harmonics $Y_{\ell}^{m}$ corresponds to the even (or, polar) sector in $D=4$, while term proportional to $Y_{i}^{(T)}{}_{\ell}^{m}$ describes to the odd (or, axial) sector of perturbations in $D=4$. Since we are ultimately interested in the $D=4$ limit, we will henceforth ignore the last term in (3.1), proportional to $Y_{ij}^{(TT)}$, which corresponds to the tensor sector that is present only in higher-dimensional spacetimes. Owing to the background symmetry, the component $h_{T}(t,r)$ decouples from the other metric fluctuations and is non-dynamical in $D=4$.
After fixing the gauge $h_{2}={\cal H}_{0}={\cal K}=G=0$, [^19] Note that this choice slightly differs from what is commonly referred to as the Regge–Wheeler gauge [115]. and solving for the constraint variables, one finds the following equations for the physical degrees of freedom $\Psi_{\text{RW}}$ and $\Psi_{\text{Z}}$ [7]:
$\frac{{\rm d}^{2}\Psi_{\rm RW}}{{\rm d}r_{\star}^{2}}+\Big(\omega^{2}-V_{\rm RW}(r)\Big)\Psi_{\rm RW}$ $=0$ (3.2)
$\frac{{\rm d}^{2}\Psi_{\rm Z}}{{\rm d}r_{\star}^{2}}+\Big(\omega^{2}-V_{\rm Z}(r)\Big)\Psi_{\rm Z}$ $=0,$ (3.3)
where ${\rm d}r_{\star}/{\rm d}r\equiv f^{-1}$, and where the potentials $V_{\rm RW}$ and $V_{\rm Z}$ are given by
$V_{\rm RW}(r)$ $=f\frac{(\ell+1)(\ell+D-4)}{r^{2}}+f^{2}\frac{(D-4)(D-6)}{4r^{2}}-ff^{\prime}\frac{(D+2)}{2r}\,$, (3.4)
$V_{\text{Z}}(r)$ $=\bigg[4(D-4)(D-2)^{4}f^{3}-8(D-2)^{2}(D-2)(D-6)\ell(\ell+D-3)f^{2}$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +4(D-2)(D-2)(D-12)\ell^{2}(\ell+D-3)^{2}f$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +2(D-2)^{3}(D+2)r^{3}f^{\prime}{}^{3}-4(D-2)^{2}(D-6)\ell(\ell+D-3)r^{2}f^{\prime}{}^{2}$ (3.5)
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ -8(D-2)^{2}\ell^{2}(\ell+D-3)^{2}rf^{\prime}+12(D-2)^{5}rf^{2}f^{\prime}+(D-2)^{3}(D(D+10)-32)r^{2}ff^{\prime}{}^{2}$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ -4(D-2)^{2}(D-2)(3D-8)\ell(\ell+D-3)rff^{\prime}$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +16\ell^{2}(\ell+D-3)^{2}(D-2)\ell(\ell+D-3)\bigg]\frac{f[2\ell(\ell+D-3)+(D-2)(rf^{\prime}-2f)]^{-2}}{4(D-2)r^{2}}\,.$
In $D=4$, the equations (3.2) and (3.3) reduce to the Regge–Wheeler and Zerilli equations, respectively [115, 116]. The exact relations between the Regge–Wheeler and Zerilli fields $\Psi_{\text{RW}}$ and $\Psi_{\text{Z}}$, and the metric perturbations in (3.1) can be found in [7].
We next move on to solving (3.2) and (3.3) perturbatively in $\omega$. As in the scalar case, we identify three distinct regions: a near zone ($r-r_{s}\ll r_{s}$), an intermediate zone ($r_{s}\lesssim r\ll\omega^{-1}$), and a far zone ($r\gg r_{s}$). In each region, we employ an approximation scheme that yields an analytic solution, valid within that regime. From the scalar example, we learned—and explicitly showed—that the far zone solution is not strictly required for matching to the EFT and determining the Love number couplings. Consequently, in the present general relativistic setup described by (3.2) and (3.3), we restrict our attention to computing the near and intermediate-zone solutions only. For simplicity in the remainder of this subsection we set $D=4$ in (3.2) and (3.3). We will return to the case of arbitrary $D$ in the EFT discussion of Section 3.2.
3.1.1 Odd sector
We first focus on the odd-parity sector of gravitational perturbations, which is described by the Regge–Wheeler equation (3.2) in $D=4$:
$\partial_{r}\big(f(r)\partial_{r}\Psi_{\text{RW}}(r)\big)+\left(\frac{\omega^{2}}{f(r)}-\frac{\ell(\ell+1)}{r^{2}}+\frac{3r_{s}}{r^{3}}\right)\Psi_{\text{RW}}(r)=0.$ (3.6)
We define two overlapping zones (near and intermediate) where we can solve this equation and use boundary conditions at the black hole horizon to fix the solution.
Near zone:
In the near-zone limit $r\rightarrow r_{s}$ (where $f\rightarrow 0$), the potential is dominated by the $\omega^{2}$ term. By employing the tortoise coordinate, the resulting equation can be written in the form of the usual wave equation
$\left(\frac{{\rm d}^{2}}{{\rm d}r_{\star}^{2}}+\omega^{2}\right)\Psi_{\text{RW}}=0\,.$ (3.7)
The independent solutions are $\mathrm{e}^{\pm i\omega r_{\star}}$. As in the scalar field case, imposing standard infalling boundary conditions at the horizon selects the following near-zone solution:
$\Psi_{\text{RW}}^{{\rm NZ}}(r)=B\operatorname{e}^{-i\omega\left(r+r_{s}\log\left(\frac{r}{r_{s}}-1\right)\right)},$ (3.8)
with $B$ an arbitrary integration constant. The solution (3.8) is needed to set the correct boundary conditions for the intermediate zone solution by matching across their region of overlap, which we now discuss.
Intermediate zone:
For values of $r$ satisfying $r_{s}\lesssim r\ll\omega^{2}$, the $\omega^{2}$ term is always small if $\omega$ is, and it can be treated perturbatively. Defining $\epsilon\equiv\omega r_{s}$, we can expand the Regge–Wheeler field as
$\Psi_{\text{RW}}=\Psi^{(0)}+\epsilon\Psi^{(1)}(r)+\epsilon^{2}\Psi^{(2)}(r)+\cdots,$ (3.9)
and then solve order-by-order in $\epsilon$. It is convenient to define a new coordinate $z\equiv r_{s}/r$ and the redefine the field $\Psi$
$u_{\ell}(z(r))$ $\equiv\sqrt{\frac{(\ell+2)(\ell-1)}{2}}\frac{r^{\ell}}{r_{s}^{\ell+1}}\Psi_{\text{RW}}(r)\,$, (3.10)
in order to recast (3.6) into
$z(1-z)u_{\ell}^{\prime\prime}+\big[2\ell+2-(2\ell+3)z\big]u_{\ell}^{\prime}-(\ell+3)(\ell-1)u_{\ell}=-\frac{\epsilon^{2}}{z^{3}(1-z)}u_{\ell}\,.$ (3.11)
The differential operator on the left hand side is hypergeometric, so that the homogeneous equation is in the standard form of a hypergeometric equation.
At zeroth and first order in $\epsilon$, we can set the right-hand side of (3.11) to zero and solve the un-sourced hypergeometric equation for $u_{\ell}$. Denoting with $u_{\ell}^{(n)}$ the $O(\epsilon^{n})$ intermediate-zone solution for $u_{\ell}$, we have
$u_{\ell}^{(n)}(z)=b_{1}^{(n)}u_{\text{reg}}(z)+b_{2}^{(n)}u_{\text{irr}}(z),\qquad\text{for }n=0,1,$ (3.12)
with $b_{1}^{(n)}$ and $b_{2}^{(n)}$ arbitrary integration constants, to be determined by matching to the near-zone solution (3.8). The solutions $u_{\text{reg}}$ and $u_{\text{irr}}$ are the two independent hypergeometric functions solving the homogeneous version of (3.11): [^20] Note that the regular solution (3.20) can be equivalently written as $u_{\text{reg}}(z)=-\frac{24(2\ell)!}{\left((\ell+2)!\right)^{2}}z^{-2\ell-1}\,{}_{2}F_{1}\left[\begin{array}[]{c}-\ell-2\,,\,-\ell+2\\[-3.0pt] -2\ell\end{array}\Big\rvert\,z\,\right]\,$, (3.13) by using hypergeometric identities.
$u_{\text{irr}}(z)$ $=\,{}_{2}F_{1}\left[\begin{array}[]{c}\ell-1\,,\,\ell+3\\[-3.0pt] 2\ell+2\end{array}\Big\rvert\,z\,\right],$ (3.16)
$u_{\text{reg}}(z)$ $=(-z)^{-\ell-3}\,{}_{2}F_{1}\left[\begin{array}[]{c}2-\ell\,,\,\ell+3\\[-3.0pt] 5\end{array}\Big\rvert\,\frac{1}{z}\,\right]$ (3.19)
$=24(-z)^{-\ell-3}\sum_{k=0}^{\ell-2}\frac{(-1)^{k}(\ell-2)!}{(\ell-2-k)!k!}\frac{\Gamma(\ell+k+3)}{\Gamma(\ell+3)\Gamma(k+5)}z^{-k},$ (3.20)
where $u_{\text{reg}}$ is regular at $z=\infty$ ($r=0$), while $u_{\text{irr}}$ is singular there. In particular, $u_{\text{reg}}$ is a finite polynomial in $1/z$, while $u_{\text{irr}}$ contains a $\log(1-z)$.
Comparing the $z\rightarrow 1$ ($r\to r_{s}$) limit of the intermediate zone solution (3.12) with the small-$\epsilon$ expansion of the near-zone solution (3.8), which in these variables reads
$u_{\text{NZ}}(z)=\frac{A_{\ell}}{z^{\ell}}\mathrm{e}^{-i\epsilon\left(\frac{1}{z}+\log\left(\frac{1-z}{z}\right)\right)},\qquad A_{\ell}\equiv\frac{B}{r_{s}}\sqrt{\frac{(\ell+2)(\ell-1)}{2}},$ (3.21)
we find the following matching conditions for the constant coefficients appearing in (3.12): [^21] In doing this, we used the limit of the hypergeometric function $\lim_{z\rightarrow 1^{-}}\,{}_{2}F_{1}\left[\begin{array}[]{c}a\,,\,b\\[-3.0pt] a+b\end{array}\Big\rvert\,z\,\right]=-\frac{\Gamma(a+b)}{\Gamma(a)\Gamma(b)}\left[\log(1-z)-2\gamma_{E}+\psi(a)+\psi(b)\right]+O(1-z)\,$, (3.22) with $a=\ell-1$ and $b=\ell+3$, and the Chu–Vandermonde identity $_{2}F_{1}\left[\begin{array}[]{c}-m\,,\,b\\[-3.0pt] c\end{array}\Big\rvert\,1\,\right]=\frac{(-1)^{m}(b-c)!}{(b-c-m)!}\frac{\Gamma(c)}{\Gamma(c+m)}\qquad\qquad\text{for }m=0,1,2,\cdots,$ (3.23) with $m=\ell-2$, $b=\ell+3$ and $c=5$.
$b_{1}^{(0)}$ $=-\frac{A_{\ell}}{24}\frac{(\ell+2)!}{(\ell-2)!},$ $\qquad b_{2}^{(0)}$ $=0,$ (3.24)
$b_{1}^{(1)}$ $=-\frac{iA_{\ell}}{24}\frac{(\ell+2)!}{(\ell-2)!}\left(H_{\ell+2}+H_{\ell-2}-1\right),$ $\qquad b_{2}^{(1)}$ $=iA_{\ell}\frac{(\ell+2)!(\ell-2)!}{(2\ell+1)!}\,.$ (3.25)
At second order in $\epsilon$, the intermediate-zone solution $u^{(2)}$ satisfies
$z(1-z)\partial_{z}^{2}u_{\ell}^{(2)}+\left[2\ell+2-(2\ell+3)z\right]\partial_{z}u_{\ell}^{(2)}-(\ell+3)(\ell-1)u_{\ell}^{(2)}=-\frac{u_{\ell}^{(0)}(z)}{z^{3}(1-z)}\equiv S^{(u)}_{\ell}(z).$ (3.26)
which is an inhomogeneous equation with source $S^{(u)}_{\ell}(z)$, which can be solved using a Green’s function. The general solution reads: [^22] We used the Green’s function $G(x,y)=\big[u_{\text{irr}}(x)u_{\text{reg}}(y)-u_{\text{reg}}(x)u_{\text{irr}}(y)\big]\frac{y^{2\ell+1}}{W_{\ell}}\theta(x-y)\,.$ (3.27) It is also useful to recall the standard formula for the Wronskian: $W\left[f_{1}(y),f_{2}(y)\right]=(1-c)y^{-c}(1-y)^{c-a-b-1}$, where $f_{1}(y)=\,{}_{2}F_{1}\left[\begin{array}[]{c}a\,,\,b\\[-3.0pt] c\end{array}\Big\rvert\,y\,\right]\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ {\rm and}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ f_{2}(y)=y^{1-c}\,{}_{2}F_{1}\left[\begin{array}[]{c}a-c+1\,,\,b-c+1\\[-3.0pt] 2-c\end{array}\Big\rvert\,y\,\right]\,.$ (3.28) Here $f_{1}(z)$ and $f_{2}(z)$ correspond to the homogeneous solutions $u_{\text{irr}}(z)$ and $u_{\text{reg}}(z)$ respectively (with $u_{\text{reg}}(z)$ given in (3.13)), for $a=\ell-1$, $b=\ell+3$ and $c=2\ell+2$.
$u_{\ell}^{(2)}(z)$ $=\left[b_{1}^{(2)}+b_{1}^{(0)}\int^{z}{\rm d}y\frac{y^{2\ell-2}}{(1-y)W_{\ell}}u_{\text{reg}}(y)u_{\text{irr}}(y)\right]u_{\text{reg}}(z)$ (3.29)
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \,\penalty 10000\ +\left[b_{2}^{(2)}-b_{1}^{(0)}\int^{z}{\rm d}y\frac{y^{2\ell-2}}{(1-y)W_{\ell}}u_{\text{reg}}(y)^{2}\right]u_{\text{irr}}(z)\,$,
where $W_{\ell}=-24(2\ell+1)!/((\ell+2)!)^{2}$. In (3.29) we are omitting the lower integration bound—see the discussion around equation (2.21).
Evaluating the integrals in (3.29) and matching the result to the near zone solution at second order in $\epsilon$, one can fix the coefficients $b_{1}^{(2)}$ and $b_{2}^{(2)}$ (see (2.22) for an example in the scalar field case). Plugging the expressions back into (3.10), one finds the following $\ell=2$ intermediate-zone solution for the Regge–Wheeler field:
$\Psi^{\rm IZ}_{\text{RW},\ell=2}(r)\underset{\frac{r}{r_{s}}\to\infty}{=}$ $\penalty 10000\ B\frac{r^{3}}{r_{s}^{3}}+iB\omega r_{s}\left(\frac{13r^{3}}{12r_{s}^{3}}+\frac{r_{s}^{2}}{5r^{2}}\right)$ (3.30)
$+B\omega^{2}r_{s}^{2}\bigg[-\frac{r^{5}}{14r_{s}^{5}}-\frac{13r^{4}}{42r_{s}^{4}}+\frac{r^{3}}{r_{s}^{3}}\left(-\frac{95+8\pi^{2}}{48}+\frac{107}{210}\log\frac{r_{s}}{r}\right)+\frac{319r^{2}}{420r_{s}^{2}}$
$\hskip 48.36958pt+\frac{153r}{280r_{s}}+\frac{223}{420}+\frac{363r_{s}}{560r}+\frac{r_{s}^{2}}{25r^{2}}\left(6-5\log\frac{r_{s}}{r}\right)\bigg]+O\left(\frac{r_{s}^{3}}{r^{3}}\right)\,.$
The analogous solutions for $\ell=3,4$ are provided in Appendix C. We will match these general relativistic solutions to point-particle EFT in Section 3.2. Before proceeding, we present the analogous analysis for the Zerilli equation in the even sector.
3.1.2 Even sector
The even-parity sector is governed by the Zerilli equation, which in $D=4$ is [116]
$\partial_{r}\left(f(r)\partial_{r}\Psi_{\text{Z}}(r)\right)+\left(\frac{\omega^{2}}{f(r)}-\frac{r_{s}}{r^{3}}-\frac{2\lambda}{3r^{2}}-\frac{8\lambda^{2}(2\lambda+3)}{3(2\lambda r+3r_{s})^{2}}\right)\Psi_{\text{Z}}(r)=0\,$, (3.31)
where $\lambda\equiv(\ell-1)(\ell+2)/2$. As before, we split space up into a a near zone ($r-r_{s}\ll r_{s}$) and an intermediate zone ($r_{s}\lesssim r\ll\omega^{-1}$). (The far zone is again not needed for the matching calculation we are undertaking.) We consider each of these in turn.
Near zone:
In the near zone regime, ($r-r_{s}\ll r_{s}$), the $\omega$-independent piece of the potential is subdominant, and can be neglected. The Zerilli equation then takes the same form as the Regge–Wheeler equation in this region:
$\left(\frac{{\rm d}^{2}}{{\rm d}r_{\star}^{2}}+\omega^{2}\right)\Psi_{\text{Z}}=0.$ (3.32)
Imposing ingoing boundary conditions at the horizon determines the following solution:
$\Psi_{\text{Z}}^{\rm NZ}(r)=C\operatorname{e}^{-i\omega\left(r+r_{s}\log\left(\frac{r}{r_{s}}-1\right)\right)},$ (3.33)
with $C$ a free integration constant.
Intermediate zone:
In the intermediate zone, one can derive a solution perturbatively in $\omega$, as in the scalar and the gravitational odd cases. However, a more direct and efficient approach is to take advantage of Chandrasekhar’s symmetry [76, 77, 78]. The Chandrasekhar duality is a symmetry—present only in D=4D=4—that relates the Regge–Wheeler and Zerilli equations. Mathematically, it belongs to the class of Darboux transformations of second-order ordinary differential equations, and can be used as a way of generating solutions [123, 124, 122]. In particular, if we know a solution $\Psi_{\text{RW}}(r)$ of the Regge–Wheeler equation (3.6), the Chandrasekhar duality guarantees that $\Psi_{\text{Z}}(r)$, defined by
$\Psi_{\text{Z}}(r)=\Big(f(r)\partial_{r}-{\cal W}(r)\Big)\Psi_{\text{RW}}(r),$ (3.34)
where $\cal W$} is given by
${\cal W}(r)\equiv\frac{3r_{s}(r_{s}-r)}{r^{2}(3r_{s}+2\lambda r)}-\frac{2\lambda(\lambda+1)}{3r_{s}}\,$, (3.35)
solves the $D=4$ Zerilli equation (3.31). We can therefore easily generate solutions to the Zerilli equation by mapping our known solutions to the Regge–Wheeler equation.
To carry this out in practice, let us thus redefine $\Psi_{\text{RW}}(r)$ in (3.34) using the the field redefinition (3.10), with $u_{\ell}(z)$ given by
$u_{\ell}(x)=u_{\ell}^{(0)}(z)+\epsilon u_{\ell}^{(1)}(z)+\epsilon^{2}u_{\ell}^{(2)}(z),$ (3.36)
where $u_{\ell}^{(0)}$ and $u_{\ell}^{(1)}$ are (cf. (3.12))
$u_{\ell}^{(n)}(z)=c_{1}^{(n)}u_{\text{reg}}(z)+c_{2}^{(n)}u_{\text{irr}}(z),\qquad\text{for }n=0,1,$ (3.37)
while $u_{\ell}^{(2)}$ is (cf. (3.29))
$u_{\ell}^{(2)}(z)=$ $\left[c_{1}^{(2)}+\int^{z}{\rm d}y\frac{y^{2\ell-2}}{(1-y)W_{\ell}}u_{\ell}^{(0)}(y)u_{\text{irr}}(y)\right]u_{\text{reg}}(z)$ (3.38)
$+\left[c_{2}^{(2)}-\int^{z}{\rm d}y\frac{y^{2\ell-2}}{(1-y)W_{\ell}}u_{\ell}^{(0)}(y)u_{\text{reg}}(y)\right]u_{\text{irr}}(z).$
Here $u_{\text{irr}}$ and $u_{\text{reg}}$ can be read off from (3.16)–(3.20). Then, it is straightforward to check that the field $\Psi_{\text{Z}}$ constructed in (3.34) satisfies (3.31) up to second order in $\omega$. We stress that, up to this point, we are not yet making any statements about the physical implications of the symmetry. We are merely using the duality as a solution-generating technique. In particular, the integration constants $c_{j}^{(n)}$ appearing in (3.37) and (3.38) are unrelated to the $b_{j}^{(n)}$ in (3.12) and (3.12). We will discuss the significance of the duality, thought of as a symmetry, for the Love numbers later on.
By matching to the near zone solution (3.33), we can then easily determine the $c_{j}^{(n)}$ constants in terms of the amplitude $C$. For $\ell=2$, this yields the following result for the intermediate-zone solution to the Zerilli equation:
$\Psi_{\text{Z},\ell=2}^{\rm IZ}(r)\underset{\frac{r}{r_{s}}\to\infty}{=}$ $C\left(\frac{r^{3}}{r_{s}^{3}}+\frac{3r^{2}}{4r_{s}^{2}}-\frac{9r}{16r_{s}}-\frac{21}{64}+\frac{63r_{s}}{256r}-\frac{189r_{s}^{2}}{1\,024r^{2}}\right)$ (3.39)
$+iC\omega r_{s}\left(\frac{4r^{3}}{3r_{s}^{3}}+\frac{r^{2}}{r_{s}^{2}}-\frac{3r}{4r_{s}}-\frac{7}{16}+\frac{21r_{s}}{64r}-\frac{59r_{s}^{2}}{1\,280r^{2}}\right)$
$+C\omega^{2}r_{s}^{2}\bigg[-\frac{r^{5}}{14r_{s}^{5}}-\frac{67r^{4}}{168r_{s}^{4}}-\frac{r^{3}}{r_{s}^{3}}\left(\frac{\pi^{2}}{6}+\frac{1\,711}{672}-\frac{107}{210}\log\frac{r_{s}}{r}\right)$
$\hskip 49.79231pt-\frac{r^{2}}{r_{s}^{2}}\left(\frac{\pi^{2}}{8}+\frac{11\,119}{13\,440}-\frac{107}{280}\log\frac{r_{s}}{r}\right)+\frac{r}{r_{s}}\left(\frac{3\pi^{2}}{32}+\frac{8\,739}{3\,584}-\frac{321}{1\,120}\log\frac{r_{s}}{r}\right)$
$\hskip 49.79231pt+\frac{7\pi^{2}}{128}+\frac{34\,591}{30\,720}-\frac{107}{640}\log\frac{r_{s}}{r}-\frac{r_{s}}{r}\left(\frac{21\pi^{2}}{512}+\frac{6\,431}{40\,960}-\frac{321}{2\,560}\log\frac{r_{s}}{r}\right)$
$\hskip 49.79231pt+\frac{r_{s}^{2}}{r^{2}}\left(\frac{63\pi^{2}}{2\,048}+\frac{431\,313}{819\,200}-\frac{3\,011}{10\,240}\log\frac{r_{s}}{r}\right)\bigg]+O\left(\frac{r_{s}^{3}}{r^{3}}\right).$
One can carry out the same exercise for higher multipoles. See Appendix C for the explicit $\ell=3,4$ solutions.
3.2 EFT calculation of dynamical Love numbers
We now match the general relativistic solutions we have just obtained to the worldline EFT of a point particle coupled to gravity, order by order in frequency. As argued in the toy example of the scalar field, the matching can be carried out at large $r$, but still within the intermediate zone of the full solution, where the EFT is well defined.
The point-particle EFT in $D$ spacetime dimensions, to quadratic order in the bulk graviton fluctuation $g_{\mu\nu}=\eta_{\mu\nu}+2h_{\mu\nu}/M_{\rm Pl}^{(D-2)/2}$, can be written as [7, 125]
$S=S_{\rm pp}$ $+\int{\rm d}^{D}x\sqrt{-g}\left[-\frac{1}{2}\nabla_{\lambda}h_{\mu\nu}\nabla^{\lambda}h^{\mu\nu}+\nabla_{\lambda}h_{\mu\nu}\nabla^{\nu}h^{\mu\lambda}-\nabla_{\mu}h\nabla_{\nu}h^{\mu\nu}+\frac{1}{2}\nabla_{\mu}h\nabla^{\mu}h\right]$ (3.40)
$+\int{\rm d}\tau\sum_{\ell=2}^{\infty}\left[Q_{B}^{A_{\ell}\,j}(\tau)B^{(\ell)}_{A_{\ell}\,j}+Q_{E}^{A_{\ell}}(\tau)E^{(\ell)}_{A_{\ell}}\right$],
where $S_{\rm pp}$ is the worldline action (2.30) of the point particle, and where $Q_{B,E}$ are composite operators describing the induced response of the object to external gravitational $B$ and $E$ fields, built from some internal degrees of freedom, $X$. In (3.40), we introduced the multi-index notation $A_{\ell}\equiv i_{1}\cdots i_{\ell}$ and the operators [46, 112, 56, 19]
[…]
which are symmetrized derivatives of the gravito-magnetic and gravito-electric fields, which are themselves defined in terms of the Weyl tensor as [^23] In $D>4$, $B_{i_{1}i_{2}j}$ and $E_{i_{1}i_{2}}$ do not exhaust all independent components of the Weyl tensor. While in $D=4$ $C_{ijkl}$ can be re-expressed in terms of the electric and magnetic components, this is not the case in higher spacetime dimensions. This implies that in (3.40) one must add an additional action term describing the response of a purely tensorial degree of freedom [7]. Since this extra sector is decoupled from the $B$ and $E$ operators at quadratic order, and will play no role in the following discussion, we have omitted it from (3.40).
$B_{i_{1}i_{2}j}$ $\equiv C_{0i_{1}i_{2}j},$ (3.43)
$E_{i_{1}i_{2}}$ $\equiv C_{0i_{1}0i_{2}}.$ (3.44)
In (3.40) we implicitly put ourselves in the rest frame of the point particle. If desired, the expressions can be covariantized using the projector ${\mathbb{P}^{\nu}}_{\mu}\equiv\delta_{\mu}^{\nu}+u_{\mu}u^{\nu}$, as discussed in Section 2.2.
We will follow the same strategy as in the scalar field case. First, we will obtain the induced gravitational field sourced by the point-particle in the presence of a background field, neglecting the nonlinearities of gravity. This will fix the homogeneous solution, up to an overall amplitude. Then, we will include the coupling to gravity using the Born series [75, 74] to determine the particular solution induced by the Schwarzschild corrections to the bulk potential. Finally we will match this EFT solution to the general relativity calculations performed in Section 3.1. As before, we treat each parity sector separately, beginning with the parity odd case.
3.2.1 Odd sector
First consider the odd sector of the point-particle EFT action (3.40). We double the number of fields on a closed-time contour and denote the graviton fields on the two-sided path by $h_{1}^{\mu\nu}$ (inserted on the forward part of the contour) and $h_{2}^{\mu\nu}$ (inserted on the backward part of the contour). As before we define the Keldysh basis
$h_{+}\equiv\frac{1}{2}\left(h_{1}+h_{2}\right),\qquad h_{-}\equiv h_{1}-h_{2}.$ (3.45)
We obtain an effective action for $h_{\pm}$ by integrating out the $X$ degrees of freedom on which the $Q$ depend
$\operatorname{e}^{i\Gamma^{\text{in-in}}_{\text{int}}[h_{\pm}]}=\int\mathcal{D}X_{+}\mathcal{D}X_{-}\,\operatorname{e}^{iS\left[h_{\pm},X_{\pm}\right]}.$ (3.46)
Perturbatively we can replace $Q_{B}$, at leading order, by its linear response
$\langle Q_{B,\,I}^{A_{\ell}\,j}(\tau)\rangle=\int{\rm d}\tau^{\prime}\,K_{IJ}^{(B)}{}^{A_{\ell}\,j|A_{\ell^{\prime}}\,j^{\prime}}(\tau-\tau^{\prime}){B}^{J}_{A_{\ell^{\prime}}\,j^{\prime}}(\tau^{\prime})\,$, (3.47)
where $I,J=\{+,-\}$, and $K_{IJ}^{(B)}{}^{A_{\ell}\,j|A_{\ell^{\prime}}\,j^{\prime}}$ is a response kernel, related to the two-point function of $Q_{B}$ in the Keldysh basis [46, 112, 56, 19] by
[…]
As we discuss below, we remain agnostic about its detailed form, which we will parametrize in the most general way, compatible with the symmetries of the problem. [^24] We are assuming that the kernel admits an expansion in small frequency, which relies on the characteristic time scale of the black hole’s internal dynamics being much faster than the timescales on which we are probing the system. In this case, this is natural because we are considering $\omega\ll r_{s}^{-1}$. For non-spinning objects, the tensorial structure of the kernel in the spatial indices reduces to a product of Kronecker deltas, leading to the following in-in effective action:
[…]
which we can use to compute the one-point function of the gravito-magnetic field.
Magnetic one-point function
The effective action (3.49) depends on the gravitational field $h$ through the magnetic component of the Weyl tensor. Thus, rather than computing the one-point function of the graviton field $h$, it is more convenient to evaluate the one-point function of $B$ itself. We therefore wish to compute:
$\langle B_{+\,abc}(t,\vec{x})\rangle_{\text{in-in}}=\int\mathcal{D}h_{+}\mathcal{D}h_{-}B_{+\,abc}(t,\vec{x})\operatorname{e}^{i\Gamma^{\text{in-in}}_{\text{int}}[h_{\pm},X_{\pm}]},$ (3.50)
in the presence of a background $\mkern 2.5mu\overline{\mkern-2.5muB\mkern-2.5mu}\mkern 2.5mu$ for the magnetic field. We are computing the expectation value of $B^{+}$ because this is the field combination that has a classical interpretation. For the external classical source we fix $h^{+}_{1}=h^{+}_{2}\equiv h^{+}$ (equivalently, $B^{+}_{1}=B^{+}_{2}\equiv B^{+}$ for the linearized Weyl tensor). Using the explicit form of the in-in effective action (3.49), we obtain the following expression for the one-point function
$\langle B_{+\,abc}(t,\vec{x})\rangle_{\text{in-in}}=i\sum_{\ell=2}^{\infty}\int{\rm d}\tau_{1}{\rm d}\tau_{2}\,K_{\ell}^{(B)}(\tau_{2}-\tau_{1})\langle B_{+\,abc}(t,\vec{x}){B}_{-\,A_{\ell}\,j}\left(\tau_{2}\right)\rangle\mkern 2.5mu\overline{\mkern-2.5mu{B}\mkern-2.5mu}\mkern 2.5mu_{+}^{A_{\ell}\,j}\left(\tau_{1}\right).$ (3.51)
where $K_{\ell}^{(B)}\equiv K_{+-,\ell}^{(B)}$ is the response function
$K_{+-}^{(B)}(\tau_{2}-\tau_{1})=-G_{R}^{(Q_{B})}(\tau_{2}-\tau_{1})\equiv i\langle\left[Q_{B,+}(\tau_{2}),Q_{B,-}(\tau_{1})\right]\rangle\theta(\tau_{2}-\tau_{1}).$ (3.52)
The UV information is stored in the response function $K_{\ell}^{(B)}$, which depends on the microscopic details of the object (the boundary conditions at the surface, the object’s internal dynamics, etc.). Assuming that the timescale of the object’s dynamics is parametrically faster than the timescale on which we probe the system (so that we can think of the internal dynamics as instantaneous), we can expand the Fourier transform of the response function
$K_{\ell}^{(B)}(\tau_{2}-\tau_{1})$ $=\int\frac{{\rm d}{\omega}}{2\pi}\operatorname{e}^{-i{\omega}(\tau_{2}-\tau_{1})}K_{\ell}^{(B)}(\omega)$ (3.53)
$=\frac{1}{\ell!}\int\frac{{\rm d}{\omega}}{2\pi}\operatorname{e}^{-i{\omega}(\tau_{2}-\tau_{1})}\left[\frac{\lambda_{0,\ell}^{B}}{2}+i{\omega}r_{s}\frac{\lambda_{1,\ell}^{B}}{2}+({\omega}r_{s})^{2}\frac{\lambda_{2,\ell}^{B}}{2}+\cdots\right]\,$,
where the parameters $\lambda_{i}$ are the response coefficients. Terms with odd powers of $\omega$ (which are time-reversal odd) capture the dissipative response of the object, while those with even powers of $\omega$ parametrize the conservative tidal deformability. In particular, $\lambda_{0,\ell}^{B}$ is related to the magnetic static Love numbers, while $\lambda_{2,\ell}^{B}$ corresponds to the quadratic-in-frequency dynamical Love numbers that we wish to compute.
To determine the one-point function (3.51), we need to compute the two-point function of $B$. We first express the magnetic component of the Weyl tensor in terms of the graviton field. Since the linearized Weyl tensor on Minkowski space is a gauge-invariant quantity, it is useful to fix a gauge. A convenient choice is the de Donder gauge, defined by $\partial^{\mu}(h_{\mu\nu}-\frac{1}{2}\eta_{\mu\nu}h)=0$, with $h\equiv\eta^{\mu\nu}h_{\mu\nu}$. In this gauge, the linearized equations of motion are $\square h_{\mu\nu}=0$, and the on-shell magnetic component of the Weyl tensor reads
$B_{abc}=C_{0abc}=2\left(\partial_{a}\partial_{[b}h_{c]0}-\partial_{0}\partial_{[b}h_{c]a}\right).$ (3.54)
Similarly, we can write the symmetrized derivative of the Weyl tensor as
$B_{A_{\ell}\,j}=\partial_{i_{L}}h_{0j}-\partial_{j}\partial_{(i_{1}}\cdots\partial_{i_{\ell-1}}h_{i_{\ell})_{T}0}+\partial_{0}F_{A_{\ell}\,j}\,$, (3.55)
where we have defined the combination
$F_{A_{\ell}j}\equiv-\partial_{(i_{1}}\cdots\partial_{i_{\ell-1}}h_{i_{\ell})_{T}\,j}+\partial_{j}\partial_{(i_{1}}\cdots\partial_{i_{\ell-2}}h_{i_{\ell-1}i_{\ell})_{T}}\,$, (3.56)
which is symmetric and traceless in the first $i_{1}\cdots i_{\ell}$ indices.
Let us start by computing the product ${B}^{-}_{A_{\ell}\,j}\mkern 2.5mu\overline{\mkern-2.5mu{B}\mkern-2.5mu}\mkern 2.5mu^{+\,A_{\ell}j}$. It is convenient to unpack it using (3.55) and (3.56), and evaluate term by term. First, notice the following identity:
[…]
which holds when the $\ell$ indices denoted by the multi-index $A_{\ell}$ are symmetrized on the left-hand side (we suppress the $\pm$ superscripts on $h$ and $\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu$ in these intermediate steps for simplicity, and restore them at the end). Removing the traces from (3.57)—as prescribed by (3.55)—amounts to correcting the right-hand side of (3.55) by additional terms with the schematic form $\partial_{k}\partial^{k}h$, $\partial_{k}\partial^{k}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu$, or $\partial^{k}h_{k0}\partial^{j}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{j0}$. From the equations of motion in de Donder gauge, it follows that all these terms are $O(\omega^{2})$. We can neglect these terms since the static Love numbers of black holes vanish, $\lambda_{0,\ell}^{B}=0$ so $K_{\ell}^{(B)}(\omega)$ starts linearly in $\omega$. Therefore, such terms would correspond to at least $O(\omega^{3})$ contributions in the one-point function (3.51), which we neglect. Similarly, the combination $\partial_{0}F_{A_{\ell}\,j}\partial_{0}\mkern 2.5mu\overline{\mkern-2.5muF\mkern-2.5mu}\mkern 2.5mu^{A_{\ell}\,j}$, arising from the product ${B}^{-}_{A_{\ell}\,j}\mkern 2.5mu\overline{\mkern-2.5mu{B}\mkern-2.5mu}\mkern 2.5mu^{+\,A_{\ell}\,j}$, is of $O(\omega^{2})$ and gives subdominant corrections to the dynamical Love numbers for black holes. Therefore, we are left with
${B}_{-\,A_{\ell}j}\mkern 2.5mu\overline{\mkern-2.5mu{B}\mkern-2.5mu}\mkern 2.5mu_{+}^{A_{\ell}\,j}=$   $2\frac{\ell+1}{\ell}\partial_{A_{\ell}}h_{0j}\partial^{A_{\ell-1}}\partial^{[i_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{j]}{}_{0}+\partial_{A_{\ell}}h_{0j}\partial_{0}\mkern 2.5mu\overline{\mkern-2.5muF\mkern-2.5mu}\mkern 2.5mu^{A_{\ell}\,j}-(\partial_{j}\partial_{(A_{\ell-1}}h_{i_{\ell})_{T}0})\partial_{0}\mkern 2.5mu\overline{\mkern-2.5muF\mkern-2.5mu}\mkern 2.5mu^{A_{\ell}\,j}$ (3.58)
$+\partial_{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{0j}\partial_{0}F^{A_{\ell}\,j}-(\partial_{j}\partial_{(A_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{j_{\ell})_{T}0})\partial_{0}F^{A_{\ell}\,j}+O(\omega^{2}).$
Let us focus for a moment on the terms involving $F$ in (3.58). For the same reasons as above, removing traces from $\partial_{A_{\ell-1}}{h}_{j_{\ell}0}$ and $\partial_{A_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{j_{\ell}0}$ corresponds to subtracting terms of order $\omega$, as a consequence of the equations of motion and gauge conditions. Moreover, since the $A_{\ell}$ indices in $F^{A_{\ell}\,j}$ and $\mkern 2.5mu\overline{\mkern-2.5muF\mkern-2.5mu}\mkern 2.5mu^{A_{\ell}j}$ are already symmetrized, we can drop the brackets $(\cdots)_{T}$ from $(\partial_{j}\partial_{(A_{\ell-1}}{h}_{j_{\ell})_{T}0})\partial^{0}\mkern 2.5mu\overline{\mkern-2.5muF\mkern-2.5mu}\mkern 2.5mu^{A_{\ell}j}$ and $(\partial_{j}\partial_{(A_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{j_{\ell})_{T}0})\partial^{0}F^{i_{L}j}$, up to $O(\omega^{3})$ corrections, which we neglect. As a result, (3.58) is
${B}_{-\,A_{\ell}\,j}\mkern 2.5mu\overline{\mkern-2.5mu{B}\mkern-2.5mu}\mkern 2.5mu_{+}^{A_{\ell}\,j}=$   $2\frac{\ell+1}{\ell}\partial_{A_{\ell}}h_{0j}\partial^{A_{\ell-1}}\partial^{[i_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{j]}{}_{0}$ (3.59)
$-2\frac{\ell+1}{\ell}\left[\partial_{A_{\ell}}h_{0j}\left(\partial_{0}\partial^{A_{\ell-2}}\partial^{[i_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{j]i_{\ell}}\right)+(h\leftrightarrow\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu)\right]+O(\omega^{2}).$
Combining this expression with (3.54), we thus have
$\langle B_{+\,abc}{B}_{-\,A_{\ell}\,j}\rangle\mkern 2.5mu\overline{\mkern-2.5mu{B}\mkern-2.5mu}\mkern 2.5mu_{+}^{A_{\ell}\,j}=$ $\,\,\frac{4(\ell+1)}{\ell}\langle(\partial_{a}\partial_{[b}h_{+\,c]0})(\partial_{A_{\ell}}h_{-\,j0})\rangle\left[\partial^{A_{\ell-1}}\partial^{[i_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j]}{}_{0}-\partial_{0}\partial^{A_{\ell-2}}\partial^{[i_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j]i_{\ell}}\right]$ (3.60)
$-\frac{4(\ell+1)}{\ell}\langle(\partial_{a}\partial_{[b}h_{+\,c]0})(\partial_{0}\partial_{A_{\ell-2}}\partial_{[i_{\ell-1}}h_{-\,j]i_{\ell}})\rangle\partial^{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j}{}_{0}$
$-\frac{4(\ell+1)}{\ell}\langle(\partial_{0}\partial_{[b}h_{+\,c]a})(\partial_{A_{\ell}}h_{-\,0j})\rangle\partial^{A_{\ell-1}}\partial^{[i_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j]}{}_{0}+O(\omega^{2})$
To compute this up to linear order in $\omega$, we can use the instantaneous propagator for the graviton in de Donder gauge [7]:
$\langle h_{+\,\mu\nu}(t,\vec{x})h_{-\,\rho\sigma}(\tau_{1},\vec{0})\rangle=i\delta(t-\tau_{1})\mathcal{P}_{\mu\nu\rho\sigma}^{\text{dD}}\int\frac{{\rm d}^{D-1}\vec{p}}{(2\pi)^{D-1}}\frac{\operatorname{e}^{i\vec{p}\cdot\vec{x}}}{\vec{p}^{2}},$ (3.61)
where the propagator numerator is
$\mathcal{P}_{\mu\nu\rho\sigma}^{\text{dD}}\equiv-\frac{1}{2}\left(\eta_{\mu\sigma}\eta_{\nu\rho}+\eta_{\mu\rho}\eta_{\nu\sigma}-\frac{2}{D-2}\eta_{\mu\nu}\eta_{\rho\sigma}\right).$ (3.62)
Upon substituting (3.61) into (3.60), the tensor structure of $\mathcal{P}_{\mu\nu\rho\sigma}^{\text{dD}}$ forces the two-point functions on the second and third lines of (3.60) to vanish. We are thus left with
$\langle B_{+\,abc}{B}_{-\,A_{\ell}\,j}\rangle\mkern 2.5mu\overline{\mkern-2.5mu{B}\mkern-2.5mu}\mkern 2.5mu_{+}^{A_{\ell}\,j}=\frac{4(\ell+1)}{\ell}\langle(\partial_{a}\partial_{[b}h_{+\,c]0})(\partial_{A_{\ell}}h_{-\,j0})\rangle\left[\partial^{A_{\ell-1}}\partial^{[i_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j]}{}_{0}-\partial_{0}\partial^{A_{\ell-2}}\partial^{[i_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j]i_{\ell}}\right]+O(\omega^{2}).$ (3.63)
It is convenient at this point to use the explicit form of the external tidal field $\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}$. Recall that we are interested in solving the Minkowski spacetime equation of motion in de Donder gauge, $\square\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{\mu\nu}=(\vec{\nabla}^{2}+\omega^{2})\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{\mu\nu}=0$, up to linear order in $\omega$. At this order, the tidal field solution can be expressed as a Taylor expansion around $\vec{x}=0$, with the time dependence factored out as $\operatorname{e}^{-i\omega t}$. In cartesian coordinates, the component $\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{0j}$ ca be written as
$\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{0j}(t,\vec{x})=\operatorname{e}^{-i\omega t}\sum_{\ell}c_{j|j_{1}\cdots j_{\ell}}x^{j_{1}}\cdots x^{j_{\ell}}+O(\omega^{2}),$ (3.64)
with $c$ symmetric and traceless in its $j_{1}\cdots j_{\ell}$ indices (and vanishing when totally symmetrized).
We can write the other components of $h_{\mu\nu}$ similarly. Notice that the last term in (3.63) is already of order $O(\omega)$. Therefore, it is sufficient to evaluate it using the static solution for $\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{ij}$. From the form of this solution, it is easy to show that this term vanishes identically. We provide the details in Appendix D (see in particular (D.8)). As a result, (3.60) reduces to its first term, and the one-point function for the magnetic component of the Weyl tensor reads
$\langle B_{+\,abc}(t,\vec{x})\rangle_{\text{in-in}}=4i\sum_{\ell=2}^{\infty}\frac{(-1)^{\ell}(\ell+1)}{\ell}\int$ ${\rm d}\tau_{1}{\rm d}\tau_{2}\int\frac{{\rm d}{\omega}}{2\pi}\operatorname{e}^{-i{\omega}(\tau_{2}-\tau_{1})}K^{(B)}_{\ell}({\omega})$ (3.65)
[…]
where the $(-1)^{\ell}$ prefactor arises from moving the derivatives $\partial_{i_{L}}$ out of the correlator. Plugging in the expressions for the propagator (3.61) and the external field (3.64), we obtain
$\langle B_{+\,abc}(t,\vec{x})\rangle_{\text{in-in}}=-\operatorname{e}^{-i\omega t}\sum_{\ell=2}^{\infty}\frac{4(-i)^{\ell}(\ell+1)!}{\ell}K^{(B)}_{\ell}(\omega)c^{[j|i_{1}]\cdots i_{\ell}}\mathcal{P}_{0j0[c}^{\text{dD}}\partial_{b]}\partial_{a}\int\frac{{\rm d}^{D-1}\vec{p}}{(2\pi)^{D-1}}\operatorname{e}^{i\vec{p}\cdot\vec{x}}\frac{p_{i_{1}}\cdots p_{i_{\ell}}}{\vec{p}^{2}}\,.$ (3.66)
We can Fourier transform this back into position space
$\langle B_{+abc}(t,\vec{x})\rangle_{\text{in-in}}\!=\!\operatorname{e}^{-i\omega t}\sum_{\ell=2}^{\infty}\frac{(-1)^{\ell+1}(\ell+1)!}{\ell}K^{(B)}_{\ell}(\omega)\frac{2^{\ell-1}\Gamma\left(\frac{D-3}{2}\right)\Gamma\left(\frac{5-D}{2}\right)}{\pi^{\frac{D-1}{2}}\Gamma\left(\frac{5-D}{2}-\ell\right)}c_{[j|i_{1}]\cdots i_{\ell}}\delta^{j}_{[c}\partial_{b]}\partial_{a}\frac{x^{i_{1}}\cdots x^{i_{\ell}}}{|\vec{x}|^{2\ell+D-3}}\,$, (3.67)
and simplify the tensor structure using
$c_{[j|i_{1}]\cdots i_{\ell}}\delta^{j}_{[c}\partial_{b]}\partial_{a}\left(x^{i_{1}}\cdots x^{i_{\ell}}\right)=\partial_{a}\partial_{[b}\left(x^{i_{1}}\cdots x^{i_{\ell}}\right)c_{c]|A_{\ell}}.$ (3.68)
Considering the $B_{rij}$ component of (3.66), and writing the tidal field in spherical coordinates as
$c_{c|i_{1}\cdots i_{\ell}}x^{i_{1}}\cdots x^{i_{\ell}}=c_{\text{ext}}r^{\ell+1}Y_{c}^{(T)}{}_{\ell}^{m}$, (3.69)
with $Y_{c}^{(T)}{}_{\ell}^{m}$ a vector spherical harmonic [7], we finally obtain
$\langle B_{+\,rij}(x)\rangle_{\text{in-in}}=-\operatorname{e}^{-i\omega t}\sum_{\ell=2}^{\infty}\frac{(\ell+1)!}{\ell}K^{(B)}_{\ell}(\omega)\frac{2^{\ell-1}\Gamma\left(\ell+\frac{D-3}{2}\right)}{\pi^{\frac{D-1}{2}}}\nabla_{[i}Y_{j]}^{(T)}{}_{\ell}^{m}r^{2}\partial_{r}\left(r^{-2}\frac{c_{\text{ext}}}{r^{\ell+D-4}}\right).$ (3.70)
We can also write the background tidal field in spherical coordinates
$\mkern 2.5mu\overline{\mkern-2.5muB\mkern-2.5mu}\mkern 2.5mu_{rij}$ $=2\nabla_{r}\nabla_{[i}h_{j]0}-2\partial_{t}\nabla_{[i}h_{j]r}$ (3.71)
$=2\operatorname{e}^{-i\omega t}\sum_{\ell=2}^{\infty}c_{\text{ext}}r^{2}\partial_{r}\big(r^{\ell-1}\big)\big(1+O(\omega r)\big)\nabla_{[i}Y_{j]}^{(T)}{}_{\ell}^{m}$.
The $O(\omega r)$ term in $\mkern 2.5mu\overline{\mkern-2.5muB\mkern-2.5mu}\mkern 2.5mu_{rij}$ represents a far-zone, frequency-dependent correction to the static tidal field. It exhibits a different scaling in $r$ compared to the standard $r^{\ell}$ and $r^{-\ell-1}$ falloffs. Therefore, as long as we are concerned with matching the EFT in the intermediate zone, we can consistently neglect this contribution. Putting everything together and replacing $K_{\ell}^{(B)}$ by $\mu^{2\varepsilon}K_{\ell}^{(B)}$ (as we did in the scalar section), we obtain
[…]
$\times\left(r^{\ell}+\mu^{2\varepsilon}K^{(B)}_{\ell}(\omega)\frac{(\ell+1)!(D+\ell-2)}{\ell(\ell-1)}\frac{2^{\ell-2}\Gamma\left(\ell+\frac{D-3}{2}\right)}{\pi^{\frac{D-1}{2}}}r^{-\ell+3-D}\right),$
The result (3.72) gives the one-point function of the magnetic component of the Weyl tensor in the presence of a tidal field, up to quadratic order in $\omega$, which we will use to match to the full black hole solution. Before doing this, we must introduce the coupling to gravity and regulate the UV divergences.
Gravitational corrections
The solution (3.72) captures the leading in $r$ behavior of the $B$ field, but in order to match to the relativistic solution, we must include the coupling to gravity which generates subleading terms. To do this, we work in dimensional regularization and a Born series expansion following [75, 74] to capture nonlinear $G_{\rm N}$-corrections to the Minkowski-spacetime solution. This procedure is more easily carried out in Regge–Wheeler gauge. However, the one-point function above was derived using the graviton solution in de Donder gauge. So, we must first relate the function $K_{\ell}^{(B)}(\omega)$ to the integration constants of the homogeneous solution of the Regge–Wheeler equation. This will ensure that the homogeneous solution for $\Psi_{\text{RW}}$ used below satisfies the correct boundary condition at the point particle’s location, with the worldline delta function sourcing its decaying falloff at infinity.
Homogeneous solution and boundary conditions: Away from the origin, and neglecting all $G_{\rm N}$-corrections, the $D$-dimensional Regge–Wheeler equation (3.2) in the bulk reads
$\left(\frac{{\rm d}^{2}}{{\rm d}r^{2}}-\frac{(\ell-\varepsilon)(\ell-\varepsilon+1)}{r^{2}}\right)\Psi_{\text{RW}}(r)=0\,.$ (3.73)
This is analogous to the free scalar equation (2.50) on Minkowski space, and admits the following standard growing and decaying solutions:
$\Psi_{\text{RW}}^{(h)}(r)=\mu^{-\varepsilon}B_{\text{reg}}r^{\ell+1-\varepsilon}+\frac{\mu^{\varepsilon}B_{\text{irr}}}{2\ell+1-2\varepsilon}r^{-\ell+\varepsilon}.$ (3.74)
In order for $\Psi_{\text{RW}}$ to have the same mass dimension in $D>4$, we rescaled the growing and decaying branches by $\mu^{-\varepsilon}$ and $\mu^{\varepsilon}$, respectively (cf. the scalar field case, in particular (2.49) and (2.51)).
Now we recall that the metric perturbations $h_{0}$ and $h_{1}$, defined in (3.1), are related to the Regge–Wheeler variable via (see [7] for details)
$h_{0}$ $=-q_{\ell}\operatorname{e}^{-i\omega t}r\mu^{\varepsilon}\left[(D-2)r^{\frac{2-D}{2}}\Psi_{\text{RW}}+r\partial_{r}\left(r^{\frac{2-D}{2}}\Psi_{\text{RW}}\right)\right],$ (3.75)
$h_{1}$ $=i\omega q_{\ell}\operatorname{e}^{-i\omega t}\mu^{\varepsilon}r^{3-\frac{D}{2}}\Psi_{\text{RW}},$
where $q_{\ell}\equiv 1/\sqrt{2(\ell-1)(D-2+\ell)}$. The components (3.75) can be used to evaluate the Weyl tensor in Regge–Wheeler gauge. In particular, we find $B_{rij}$ to be
$B_{rij}^{\text{RW}}$ $=\sum_{\ell=2}^{\infty}\left[r^{2}\partial_{r}\left(r^{-2}h_{0}(t,r)\right)-\partial_{0}h_{1}(t,r)\right]\nabla_{[i}Y_{j]}^{(T)}{}_{\ell}^{m}$ (3.76)
$=-\sum_{\ell=2}^{\infty}\frac{q_{\ell}\operatorname{e}^{-i\omega t}}{4}\mu^{\varepsilon}r^{1-\frac{D}{2}}\left[\left(2D-D^{2}+4r^{2}\omega^{2}\right)\Psi_{\text{RW}}(r)+4r^{2}\Psi_{\text{RW}}^{\prime\prime}(r)\right]\nabla_{[i}Y_{j]}^{(T)}{}_{\ell}^{m}$.
Plugging in the homogeneous solution (3.74) for $\Psi_{\text{RW}}$, we get
$B_{rij}^{\text{RW}}=-\sum_{\ell=2}^{\infty}q_{\ell}\operatorname{e}^{-i\omega t}\left[(\ell-1)(\ell+2-2\varepsilon)+r^{2}\omega^{2}\right]\left(B_{\text{reg}}r^{\ell}+\frac{\mu^{2\varepsilon}B_{\text{irr}}}{2\ell+1-2\varepsilon}r^{-\ell-1+2\varepsilon}\right)\nabla_{[i}Y_{j]}^{(T)}{}_{\ell}^{m}$. (3.77)
Since the linearized Weyl tensor is gauge invariant in flat space, we can directly compare $B_{rij}^{\text{RW}}$ with (3.72), obtained in de Donder gauge. By matching the $r^{\ell}$ and $r^{-\ell-1+2\varepsilon}$ coefficients in the two expressions, we can identify
$\frac{(\ell+1)!(D+\ell-2)}{\ell(\ell-1)}\frac{2^{\ell-1}\Gamma\left(\ell+\frac{D-1}{2}\right)}{\pi^{\frac{D-1}{2}}}K^{(B)}_{\ell}(\omega)=\frac{B_{\text{irr}}}{B_{\text{reg}}}.$ (3.78)
This formula relates the response $K^{(B)}(\omega)$ to the coefficients of the homogeneous solution to the Regge–Wheeler equation.
Particular solution: Now that we have obtained the homogeneous solution (3.74) for the Regge–Wheeler field, with the constants related to the worldline couplings (3.53) via (3.78), we can proceed to compute the $G_{\rm N}$-corrections to that solution and determine the renormalized coefficients. Expanding in powers of $G_{\rm N}$, the Regge–Wheeler equation (3.2) can be cast in the form
$\left(\frac{{\rm d}^{2}}{{\rm d}r^{2}}-\frac{(\ell-\varepsilon)(\ell-\varepsilon+1)}{r^{2}}\right)\Psi_{\text{RW}}(r)=V_{\Psi_{\text{RW}}}(r)\Psi_{\text{RW}}(r),$ (3.79)
where we placed all $G_{\rm N}$ terms on the right-hand side of the equation:
$V_{\Psi_{\text{RW}}}$ $=\sum_{n=1}^{\infty}\left(\frac{2G_{\text{N}}Mn_{D}\mu^{2\varepsilon}}{r^{1-2\varepsilon}}\right)^{n}\left[\frac{2\varepsilon-1}{r}\frac{{\rm d}}{{\rm d}r}+\frac{\ell^{2}+\ell-3-\varepsilon(2\ell-5)-2\varepsilon^{2}}{r^{2}}-(n+1)\omega^{2}\right]-\omega^{2},$ (3.80)
with $n_{D}$ defined in (2.27). These terms will be treated perturbatively in the Born series, in analogy with the scalar field example [74].
We are interested in the perturbative particular solution to (3.79), sourced by $V_{\Psi_{\text{RW}}}$, up to the order $O(\omega^{2}G_{\text{N}}^{2}G_{\text{N}}^{2\ell+1})$. This solution can be obtained from the Born series with up to $2\ell+4$ insertions of the $V_{\Psi_{\text{RW}}}$ potential: [^25] Recall that the static Love numbers scale as $r_{s}^{2\ell+1}$. In a diagrammatic description, they correspond to Feynman diagrams with $2\ell$ (classical) loops and $2\ell+1$ worldline mass insertions. Studying quadratic $(GM\omega)^{2}$-corrections to the tidal response is two orders further in the coupling $G$ in the perturbative expansion.
$\Psi_{\text{RW}}(r)=\Psi_{\text{RW}}^{(h)}(r)$ $+\int^{r}{\rm d}r^{\prime}G(r,r^{\prime})V_{\Psi_{\text{RW}}}(r^{\prime})\Psi_{\text{RW}}^{(h)}(r^{\prime})$ (3.81)
$+\int^{r}{\rm d}r^{\prime}G(r,r^{\prime})V_{\Psi_{\text{RW}}}(r^{\prime})\int^{r^{\prime}}{\rm d}r^{\prime\prime}G(r^{\prime},r^{\prime\prime})V_{\Psi_{\text{RW}}}(r^{\prime\prime})\Psi_{\text{RW}}^{(h)}(r^{\prime\prime})+\dots$
where $\Psi_{\text{RW}}^{(h)}$ is the homogeneous solution (3.74) and $G(r,r^{\prime})$ is the same Green’s function as in (2.59). Taking the $\varepsilon\to 0$ limit, we obtain (for $\ell=2$):
$\Psi_{\text{RW}}^{\ell=2}(r)=$ $\,\,r^{3}B_{\text{reg}}\left(1+\frac{468\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{1225}-\frac{107\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{210\varepsilon}-\frac{107\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{70}\log(\mu r)\right)$ (3.82)
$+\frac{B_{\text{reg}}}{r^{2}}\left(\frac{17704\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{7}\omega^{2}}{315}+\frac{32\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{7}\omega^{2}}{15\varepsilon}+\frac{416\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{7}\omega^{2}}{15}\log(\mu r)\right)$
$+\frac{B_{\text{irr}}}{r^{2}}\left(\frac{1}{5}+\frac{26998\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{55125}+\frac{107\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{1050\varepsilon}+\frac{107\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{210}\log(\mu r)\right),$
where we have introduced $\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu$, as in (2.61).
Renormalization and matching
We now renormalize and match (3.82) to the full solution of the Regge–Wheeler equation. To subtract the infinities, we introduce the renormalized coefficients $\bar{B}_{\text{reg}}$ and $\bar{B}_{\text{irr}}$ as follows:
$B_{\text{reg}}$ $=\bar{B}_{\text{reg}}(1+\omega^{2}\delta_{11})+\bar{B}_{\text{irr}}\omega^{2}\delta_{12},$ (3.83)
$B_{\text{irr}}$ $=\bar{B}_{\text{irr}}(1+\omega^{2}\delta_{22})+\bar{B}_{\text{reg}}\omega^{2}\delta_{21},$
where we have defined the coefficients
$\delta^{\ell=2}=\begin{pmatrix}\frac{107\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}}{210\varepsilon}&0\\ -\frac{32\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{7}}{3\varepsilon}&-\frac{107\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}}{210\varepsilon}\end{pmatrix}$. (3.84)
Plugging (3.83) back into the expression for $\Psi_{\text{RW}}$ before taking the $\varepsilon\to 0$ limit yields the following renormalized solution:
$\Psi_{\text{RW}}^{\text{R},\ell=2}=$ $\penalty 10000\ \bar{B}_{\text{reg}}\left[r^{3}+r^{3}\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}\left(\frac{468}{1225}-\frac{214}{105}\log(\mu r)\right)+\frac{\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{7}\omega^{2}}{r^{2}}\left(\frac{79472}{1575}+\frac{128}{5}\log(\mu r)\right)\right]$ (3.85)
$+\bar{B}_{\text{irr}}\left[\frac{1}{5r^{2}}+\frac{\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{r^{2}}\left(\frac{24751}{55125}+\frac{214}{525}\log(\mu r)\right)\right].$
Comparing $\Psi_{\text{RW}}^{\text{R},\ell=2}$ with (3.30), we match the renormalized constants $\bar{B}_{\text{reg}}$ and $\bar{B}_{\text{irr}}$ as functions of the frequency $\omega$, the Schwarzschild radius $r_{s}$, the renormalization scale $\mu$, and the amplitude $B$ of the UV solution (3.8):
$\bar{B}_{\text{reg}}^{\ell=2}=$ $\,B\left[1+i\omega r_{s}\frac{13}{12}-\omega^{2}r_{s}^{2}\left(\frac{121991}{58800}+\frac{\pi^{2}}{6}-\frac{107}{210}\log\mu r_{s}\right)\right],$ (3.86)
$\bar{B}_{\text{irr}}^{\ell=2}=$ $\,B\left[i\omega r_{s}^{6}-\omega^{2}r_{s}^{7}\left(\frac{1943}{2520}+\log(\mu r_{s})\right)\right].$
(Note that we rescaled $B\to Br_{s}^{2}$ so that $B_{\text{reg}}$ is a polynomial in $r_{s}$.) Using (3.78), we finally obtain
$\frac{45}{\pi}K^{(B)}_{\ell=2}(\omega)=i\omega r_{s}^{6}+\omega^{2}r_{s}^{7}\left(\frac{787}{2520}-\log(\mu r_{s})\right).$ (3.87)
Repeating the same procedure for $\ell=3$ and $\ell=4$ (see Appendix C for details), we get:
$\frac{18900}{\pi}K^{(B)}_{\ell=3}(\omega)$ $=i\omega r_{s}^{8}+\omega^{2}r_{s}^{9}\left(\frac{1\,727}{2\,520}-\log(\mu r_{s})\right),$ (3.88) $\frac{11113200}{\pi}K^{(B)}_{\ell=4}(\omega)$ $=i\omega r_{s}^{10}+\omega^{2}r_{s}^{11}\left(\frac{237\,529}{277\,200}-\log(\mu r_{s})\right).$
We see that, as expected, the magnetic dynamical love numbers are nonzero. It is also worth noting that the relative coefficient between the dissipative response coefficient and the running of the dynamical Love numbers is 1-1.
3.2.2 Even sector
We now turn to the parity-even sector of perturbations. Following the same procedure as for the odd case, we first solve for the expectation value of composite operator $Q_{E}(\tau)$, which appears in point-particle EFT (3.40), in a gravito-electric background using linear response theory:
$\langle Q_{E,I}^{A_{\ell}}(\tau)\rangle=\int{\rm d}\tau^{\prime}K_{IJ}^{(E)A_{\ell}|A_{\ell^{\prime}}}(\tau-\tau^{\prime}){E}^{J}_{A_{\ell^{\prime}}}(\tau^{\prime}),$ (3.89)
where $I,J=\{+,-\}$ again label (polar) gravitational fields (3.45) in the Keldysh basis, and $K_{IJ}^{(E)\,A_{\ell}|A_{\ell^{\prime}}}$ is related to the two-point function of $Q_{E}$ by
$\langle Q_{E,I}^{A_{\ell}}(\tau)Q_{E,J}^{A_{\ell^{\prime}}}(\tau^{\prime})\rangle=-iK_{IJ}^{(E)\,A_{\ell}|A_{\ell^{\prime}}}(\tau-\tau^{\prime})\,$, (3.90)
as in the previous section [46, 112, 56, 19]. This yields the following in-in effective action: [^26] Here we are again using the spherical symmetry of the black hole.
$\Gamma^{\text{in-in}}_{\text{int}}[h_{\pm}]=\int{\rm d}\tau_{1}{\rm d}\tau_{2}\sum_{\ell=2}^{\infty}K_{{IJ},\ell}^{(E)}(\tau_{2}-\tau_{1})E^{I}_{A_{\ell}}(\tau_{2})E^{{J}\,A_{\ell}}(\tau_{1}).$ (3.91)
Electric one-point function
Using $\Gamma^{\text{in-in}}_{\text{int}}$, we can evaluate the one-point function of the electric component of the Weyl tensor in the presence of the external tidal field $\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu_{+}$:
$\langle E_{+\,ab}(t,\vec{x})\rangle_{\text{in-in}}=i\sum_{\ell=2}^{\infty}\int{\rm d}\tau_{1}{\rm d}\tau_{2}\,K_{\ell}^{(E)}(\tau_{2}-\tau_{1})\langle E_{+\,ab}(t,\vec{x})E_{-\,A_{\ell}}\left(\tau_{2}\right)\rangle\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu_{+}^{A_{\ell}}\left(\tau_{1}\right)\,$, (3.92)
with $K_{\ell}^{(E)}\equiv K_{+-,\ell}^{(E)}$ and
$K_{+-}^{(E)}(\tau_{2}-\tau_{1})=-G_{R}^{(Q_{E})}(\tau_{2}-\tau_{1})\equiv i\langle\left[Q_{E\,+}(\tau_{2}),Q_{E\,-}(\tau_{1})\right]\rangle\theta(\tau_{2}-\tau_{1}).$ (3.93)
As in the magnetic sector, we can expand the response kernel in fourier space similarly to (3.53), where
$K_{{+-},\ell}^{(E)}(\tau_{2}-\tau_{1})$ $=\int\frac{{\rm d}{\omega}}{2\pi}\operatorname{e}^{-i{\omega}(\tau_{2}-\tau_{1})}K_{\ell}^{(E)}(\omega)$ (3.94)
$=\frac{1}{\ell!}\int\frac{{\rm d}{\omega}}{2\pi}\operatorname{e}^{-i\tilde{\omega}(\tau_{2}-\tau_{1})}\left[\lambda_{0,\ell}^{E}+i{\omega}r_{s}\lambda_{1,\ell}^{E}+({\omega}r_{s})^{2}\lambda_{2,\ell}^{E}+\cdots\right],$
$\lambda_{0,\ell}^{E}$ are the static Love numbers, $\lambda_{1,\ell}^{E}$ capture the linear-in-frequency dissipative response of the object, and $\lambda_{2,\ell}^{E}$ are the dynamical Love numbers at order $O(\omega^{2})$. Since we are interested in computing $\lambda_{2,\ell}^{E}$ for Schwarzschild black holes in four-dimensional general relativity, we will set $\lambda_{0,\ell}^{E}=0$ in the following. We will allow us to consistently neglect $O(\omega^{2})$ terms in the expressions of $\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu_{+}$ and $\langle E_{+\,ab}(t,\vec{x})E_{-\,A_{\ell}}\left(\tau_{2}\right)\rangle$ in (3.92).
From the definition of the Weyl tensor, we have on shell:
$E_{ab}$ $=C_{0a0b}=2\partial_{a}\partial_{[0}h_{b]0}-2\partial_{0}\partial_{[0}h_{b]a}=-\partial_{a}\partial_{b}h_{00}+2\partial_{0}\partial_{(a}h_{b)0}+O(\omega^{2}),$ (3.95)
$E_{A_{\ell}}$ $=-\partial_{A_{\ell}}h_{00}+2\partial_{0}\partial_{(A_{\ell-1}}h_{i_{\ell})_{T}0}+O(\omega^{2}).$
As in the odd sector, we fix de Donder gauge and expand $\langle E^{+}_{ab}(t,\vec{x})E^{-}_{A_{\ell}}\left(\tau_{2}\right)\rangle\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu^{+\,A_{\ell}}\left(\tau_{1}\right)$ in (3.92) up to order $O(\omega)$. From the gauge constraint $\partial^{\mu}(h_{\mu\nu}-\frac{1}{2}\eta_{\mu\nu}h)=0$ and equations of motion $\square h_{\mu\nu}=0$, it follows that subtracting traces in the definition of $E^{-}_{A_{\ell}}$ and $\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu^{+\,A_{\ell}}$ corresponds to terms of order $O(\omega^{2})$ in the product (3.92). Therefore, in what follows, we can simply replace $({\cdots})_{T}$ by $({\cdots})$. Then, from the contraction of the Weyl $E$-operators we find
$E_{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu^{A_{\ell}}=\partial_{A_{\ell}}h_{00}\partial^{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{00}-2\partial_{A_{\ell}}h_{00}\partial^{0}\partial^{A_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{i_{\ell}0}-2\partial_{0}\partial_{A_{\ell-1}}h_{i_{\ell}0}\partial^{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{00}+O(\omega^{2}).$ (3.96)
Putting this together, we find the contraction
$\langle E_{+\,ab}(t,\vec{x})E_{-\,A_{\ell}}\left(\tau_{2}\right)\rangle\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu_{+}^{A_{\ell}}\left(\tau_{1}\right)=-\langle\partial_{a}\partial_{b}h_{+\,00}\partial_{A_{\ell}}h_{-\,00}\rangle\left(\partial^{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{00}-2\partial^{0}\partial^{A_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{i_{\ell}0}\right)+O(\omega^{2}),$ (3.97)
where we dropped terms that are either of order $O(\omega^{2})$ or proportional to $\mathcal{P}^{\rm dD}_{0a00}=0$, with $\mathcal{P}_{\mu\nu\rho\sigma}^{\text{dD}}$ defined as in (3.62). We stress that $h_{\mu\nu}$ and $\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{\mu\nu}$ are the response and tidal fields, respectively, in de Donder gauge. It is useful to note that the combination in parenthesis is also [^27] To see this, it suffices to write the transformation that relates the fields in the two gauges: $h_{\mu\nu}^{\text{dD}}=h_{\mu\nu}^{\text{RW}}+\partial_{\mu}\xi_{\nu}+\partial_{\nu}\xi_{\mu}$, with $\xi$ the gauge parameter (for simplicity, we omit here all Keldysh indices as well as the bar symbol denoting the tidal field). Taking the linear combination (3.98), one obtains $\partial_{i_{\ell}}{h}_{00}^{\text{dD}}-2\partial_{0}{h}_{0i_{\ell}}^{\text{dD}}=\partial_{i_{\ell}}{h}_{00}^{\text{RW}}-2\partial_{0}{h}_{0i_{\ell}}^{\text{RW}}-2\partial_{0}^{2}\xi_{i_{\ell}}$. Using that ${h}_{0i_{\ell}}^{\text{RW}}=O(\omega)$ in the even sector of perturbations (see, e.g., [34]), (3.98) then follows.
$\partial_{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{00}-2\partial_{0}\partial_{A_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{i_{\ell}0}=\partial_{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{00}^{\text{RW}}+O(\omega^{2}),$ (3.98)
where $\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{00}^{\text{RW}}$ is the $tt$-component of the tidal field metric perturbation in Regge–Wheeler gauge. We can express this as [2]
$\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{00}^{\text{RW}}(t,\vec{x})=\operatorname{e}^{-i\omega t}\sum_{\ell}c_{\rm ext}Y_{\ell}^{m}r^{\ell}=\operatorname{e}^{-i\omega t}\sum_{\ell}c_{j_{1}\cdots j_{\ell}}x^{j_{1}}\cdots x^{j_{\ell}}+O(\omega^{2}),$ (3.99)
where $c^{i_{1}\cdots i_{\ell}}$ is a traceless symmetric tensor. Plugging this into (3.92) for the one-point function and using the expression (3.61) for the graviton propagator, we obtain
$\langle E_{+\,ab}(t,\vec{x})\rangle_{\text{in-in}}=-\operatorname{e}^{-i\omega t}\frac{D-3}{D-2}\sum_{\ell=2}^{\infty}\ell!(-i)^{\ell}K_{\ell}^{(E)}(\omega)c^{i_{1}\cdots i_{\ell}}\partial_{a}\partial_{b}\int\frac{{\rm d}^{D-1}\vec{p}}{(2\pi)^{D-1}}\operatorname{e}^{i\vec{p}\cdot\vec{x}}\frac{p_{i_{1}}\cdots p_{i_{\ell}}}{\vec{p}^{2}}$ (3.100)
Going back to position space, this is
$\langle E_{+\,ab}(t,\vec{x})\rangle_{\text{in-in}}=-\operatorname{e}^{-i\omega t}\sum_{\ell=2}^{\infty}K_{\ell}^{(E)}(\omega)\frac{2^{\ell-2}\ell!\,\Gamma\left(\ell+\frac{D-3}{2}\right)}{\pi^{\frac{D-1}{2}}}\frac{D-3}{D-2}c_{i_{1}\cdots i_{\ell}}\partial_{a}\partial_{b}\frac{x^{i_{1}}\cdots x^{i_{\ell}}}{|\vec{x}|^{2\ell+D-3}}.$ (3.101)
We can similarly write the background expression of the Weyl tensor. Focusing on the $rr$-component,
$\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu_{rr}$ $=-\partial_{r}^{2}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{00}+2\partial_{0}\partial_{r}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{0r}+O(\omega^{2})$ (3.102)
$=-\partial_{r}^{2}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{00}^{\text{RW}}+O(\omega^{2})=-\operatorname{e}^{-i\omega\tau}\sum_{\ell=2}^{\infty}c_{\rm ext}\,\ell(\ell-1)r^{\ell-2}Y_{\ell}^{m}+O(\omega^{2}).$
Using $c_{j_{1}\cdots j_{\ell}}x^{j_{1}}\cdots x^{j_{\ell}}=c_{\rm ext}r^{\ell}Y_{\ell}^{m}$ with $r\equiv|\vec{x}|$, and $\nabla_{r}\nabla_{r}=\partial_{r}\partial_{r}$ in (3.101), and adding it to (3.102), we get
$\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu_{rr}+\langle E_{+\,rr}\rangle_{\text{in-in}}$ $=-c_{\text{ext}}\operatorname{e}^{-i\omega t}\sum_{\ell=2}^{\infty}\ell(\ell-1)Y_{\ell}^{m}$ (3.103)
$\!\!\!\times\left(r^{\ell-2}+\mu^{2\varepsilon}K^{(E)}_{\ell}(\omega)\frac{D-3}{D-2}\frac{(\ell+D-3)(\ell+D-2)}{\ell(\ell-1)}\frac{2^{\ell-2}\ell!\,\Gamma\left(\ell+\frac{D-3}{2}\right)}{\pi^{\frac{D-1}{2}}}r^{-\ell-D+1}\right),$
where we replaced $K^{(E)}_{\ell}$ by $\mu^{2\varepsilon}K^{(E)}_{\ell}$. As before, we need to include gravity, renormalize, and match.
Including gravity
We employ the same methodology to include the nonlinearities of gravity. As before it is convenient to solve the gravitational equations away from the particle in Zerilli variables, and match the (gauge-invariant) Weyl tensor to relate the parameters of the homogeneous solution to the worldline response coefficients.
Homogeneous solution and boundary conditions: At zeroth order in the frequency and to leading order in flat-space limit, the Zerilli equation (3.3) in the bulk simply reduces to
$\left(\frac{{\rm d}^{2}}{{\rm d}r^{2}}-\frac{(\ell-\varepsilon)(\ell-\varepsilon+1)}{r^{2}}\right)\Psi_{\text{Z}}(r)=0.$ (3.104)
As in the previous cases, this admits the standard growing and decaying solutions
$\Psi_{\text{Z}}^{(h)}(r)=\mu^{-\varepsilon}B_{\text{reg}}r^{\ell+1-\varepsilon}+\frac{\mu^{\varepsilon}B_{\text{irr}}}{2\ell+1-2\varepsilon}r^{-\ell+\varepsilon},$ (3.105)
where the $\mu$ factors ensure that the field maintains the correct dimension away from $D=4$. Let us express the Weyl tensor in terms of the even metric perturbations (3.1). In particular, to leading order in the flat-space limit,
$E_{rr}=-\partial_{r}^{2}H_{0}-2i\omega\partial_{r}H_{1}.$ (3.106)
Using the constraint equations (see eq. (3.54) of [7]),
$H_{0}$ $=-\frac{\mu^{\varepsilon}r^{\frac{2-D}{2}}}{2[(D-2)(D-3)\ell(\ell-1)(\ell+D-3)(\ell+D-2)]^{1/2}}\bigg[2(D-3)(D-2)r\partial_{r}\Psi_{\text{Z}}$
$\qquad\qquad\qquad\quad+(D-3)\left(\left(D^{2}+2D(\ell-3)+2\ell(\ell-3)+8\right)-2(D-2)r^{2}\omega^{2}\right)\Psi_{\text{Z}}\bigg],$ (3.107)
$H_{1}$ $=\frac{i\omega\mu^{\varepsilon}r^{\frac{4-D}{2}}}{4[(D-2)(D-3)\ell(\ell-1)(\ell+D-3)(\ell+D-2)]^{1/2}}\bigg[2(D-2)r\partial_{r}\Psi_{\text{Z}}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\left(D^{2}-2\ell(\ell-3)-2D(\ell+1)\right)\Psi_{\text{Z}}\bigg],$ (3.108)
we can express (3.106) in terms of $\Psi_{\text{Z}}$. Comparing the resulting Weyl tensor of the homogeneous solution (3.105) with (3.103), we can extract $K^{(E)}_{\ell}$ as a function of the ratio $B_{\text{irr}}/B_{\text{reg}}$:
$\frac{D-3}{D-2}\frac{(\ell+D-3)(\ell+D-2)}{\ell(\ell-1)}\frac{2^{\ell-1}\ell!\,\Gamma\left(\ell+\frac{D-1}{2}\right)}{\pi^{\frac{D-1}{2}}}K^{(E)}_{\ell}(\omega)=\frac{B_{\text{irr}}}{B_{\text{reg}}}.$ (3.109)
This relates the worldline response to the parameters of the Zerilli solution.
Particular solution: Given the homogeneous solution (3.105), in which the coefficients $B_{\text{irr}}$ and $B_{\text{reg}}$ are related to the effective coupling $K^{(E)}_{\ell}(\omega)$ via (3.109), we can now solve for the particular solutions of the perturbative Zerilli equation
$\left(\frac{{\rm d}^{2}}{{\rm d}r^{2}}-\frac{(\ell-\varepsilon)(\ell-\varepsilon+1)}{r^{2}}\right)\Psi_{\text{Z}}(r)=V_{\Psi_{\text{Z}}}(r)\Psi_{\text{Z}}(r),$ (3.110)
where the right-hand side contains all corrections in $G_{\rm N}$ and in the frequency:
$V_{\Psi_{\text{Z}}}=\sum_{n=1}^{\infty}\left(\frac{2GMn_{D}\mu^{2\varepsilon}}{r^{1-2\varepsilon}}\right)^{n}\left[\frac{2\varepsilon-1}{r}\frac{{\rm d}}{{\rm d}r}+\frac{V^{\text{Z}}_{\ell,n}}{r^{2}}-(n+1)\omega^{2}\right]-\omega^{2},$ (3.111)
where we introduced the (fairly complicated) potential
$V^{\text{Z}}_{\ell,n}$ $\equiv\frac{1}{(\ell-1)^{2}(\ell-2\varepsilon+2)^{2}}\bigg[\ell(1-2\varepsilon)^{2}\left(\varepsilon^{2}-1\right)(\ell-2\varepsilon+1)\zeta(n-2)(\delta_{1,n}-1)$
$-(\varepsilon-1)^{2}\left(4\varepsilon^{4}-16\varepsilon^{3}+5\varepsilon^{2}+11\varepsilon-6\right)s(n-2)(\delta_{1,n}-1)$
$+(\varepsilon-3)(\varepsilon-1)^{2}(2\varepsilon-1)^{3}\zeta(n-3)(\delta_{1,n}-1)(\delta_{2,n}-1)$
$+\ell^{3}(\ell-2\varepsilon+1)^{3}\zeta(n)-\ell^{2}\left(2\varepsilon^{2}-3\varepsilon+1\right)(\ell-2\varepsilon+1)^{2}\zeta(n-1)$
$-2(\varepsilon-1)^{2}\left(\ell^{2}\left(6\varepsilon^{2}-5\varepsilon+4\right)+\ell\left(-12\varepsilon^{3}+16\varepsilon^{2}-13\varepsilon+4\right)+12\varepsilon^{3}-26\varepsilon^{2}+20\varepsilon-6\right)s(n-1)$
$+(\varepsilon-1)\Big[\ell^{4}(\varepsilon+4)-2\ell^{3}\left(2\varepsilon^{2}+7\varepsilon-4\right)+\ell^{2}\varepsilon\left(4\varepsilon^{2}+16\varepsilon-15\right)$
$+\ell\left(-8\varepsilon^{3}+4\varepsilon^{2}+8\varepsilon-4\right)+4(\varepsilon-1)^{2}\varepsilon\Big]s(n)\bigg],$ (3.112)
which involves the functions
$s(n)\equiv(n+1)\left(\frac{(1-\varepsilon)(2\varepsilon-3)}{(\ell-1)(\ell-2\varepsilon+2)}\right)^{n},\qquad\zeta(n)=\sum_{k=0}^{n}s(k).$ (3.113)
In order to display explicit expressions we now set $\ell=2$. By solving for the particular solution using the Born series and taking the $\varepsilon\to 0$ limit, we obtain the following expression for $\Psi_{\text{Z}}^{\ell=2}$:
$\Psi_{\text{Z}}^{\ell=2}(r)=$ $r^{3}B_{\text{reg}}\left([1+\omega^{2}\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\left(-\frac{107}{210\varepsilon}-\frac{2731}{9800}-\frac{107}{210}\log(\mu r)\right)\right]$ (3.114)
$+\frac{B_{\text{reg}}}{r^{2}}\left[-\frac{189\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}}{32}+\omega^{2}\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{7}\left(\frac{4937}{960\varepsilon}+\frac{2802329}{403200}+\frac{64181}{960}\log(\mu r)\right)\right]$
$+\frac{B_{\text{irr}}}{r^{2}}\left[1+\omega^{2}\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\left(\frac{107}{1050\varepsilon}+\frac{129359}{441000}+\frac{107}{210}\log(\mu r)\right)\right].$
This bare solution has divergences that must be renormalized so that we can match to the relativity calculation.
Renormalization and matching
To subtract off the $1/\varepsilon$ divergences, we introduce the renormalized coefficients $\mkern 2.5mu\overline{\mkern-2.5muB\mkern-2.5mu}\mkern 2.5mu_{\text{reg}},\mkern 2.5mu\overline{\mkern-2.5muB\mkern-2.5mu}\mkern 2.5mu_{\text{irr}}$ as defined in (3.83), but where now
$\delta^{\ell=2}=\begin{pmatrix}\frac{107\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}}{210\varepsilon}&0\\ -\frac{32\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{7}}{3\varepsilon}&-\frac{107\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}}{210\varepsilon}\end{pmatrix}$. (3.115)
Plugging this back into the expression of $\Psi_{\text{Z}}^{\ell=2}$ before taking the $\varepsilon\to 0$ limit, we obtain the renormalized Zerilli solution
$\Psi_{\text{Z}}^{\text{R},\ell=2}=\mkern 2.5mu\overline{\mkern-2.5muB\mkern-2.5mu}\mkern 2.5mu_{\text{reg}}$ $\Bigg[r^{3}-\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}r^{3}\left(\frac{2731}{9800}+\frac{214}{105}\log(\mu r)\right)+\frac{\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{7}\omega^{2}}{r^{2}}\left(\frac{1545209}{57600}+\frac{3011}{80}\log(\mu r)\right)-\frac{189\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{5}}{32r^{2}}\Bigg]$
$+\mkern 2.5mu\overline{\mkern-2.5muB\mkern-2.5mu}\mkern 2.5mu_{\text{irr}}\left[\frac{1}{5r^{2}}+\frac{\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{r^{2}}\left(\frac{111383}{441000}+\frac{214}{525}\log(\mu r)\right)\right].$ (3.116)
By matching this to the full solution (3.39), we find
$\mkern 2.5mu\overline{\mkern-2.5muB\mkern-2.5mu}\mkern 2.5mu_{\text{reg}}$ $=C+\frac{4}{3}iC\omega r_{s}+C\omega^{2}r_{s}^{2}\left(\frac{9101}{3675}-\frac{\pi^{2}}{6}+\frac{107}{210}\log(\mu r_{s})\right)$ (3.117)
$\mkern 2.5mu\overline{\mkern-2.5muB\mkern-2.5mu}\mkern 2.5mu_{\text{irr}}$ $=iC\omega r_{s}^{6}-C\omega^{2}r_{s}^{7}\left(\frac{883}{1260}+\log(\mu r_{s})\right),$
and therefore, from (3.103), this implies the following value for the $\ell=2$ electric type Love number
$\frac{45}{2\pi}K^{(E)}_{\ell=2}(\omega)=i\omega r_{s}^{6}-\omega^{2}r_{s}^{7}\left(-\frac{797}{1260}+\log(\mu r_{s})\right).$ (3.118)
Repeating the same procedure for $\ell=3$ and $\ell=4$ (see Appendix C for details) we get:
$\frac{18900}{2\pi}K^{(E)}_{\ell=3}(\omega)$ $=i\omega r_{s}^{8}-\omega^{2}r_{s}^{9}\left(-\frac{709}{840}+\log(\mu r_{s})\right),$ (3.119) $\frac{11113200}{2\pi}K^{(E)}_{\ell=4}(\omega)$ $=i\omega r_{s}^{10}-\omega^{2}r_{s}^{11}\left(-\frac{5\,501}{5\,775}+\log(\mu r_{s})\right).$
We see that the dynamical Love numbers are nonzero, and their logarithmic running is given by the dissipative response coefficients.
3.3 Symmetries
Aside from the fact that they are nonzero, the dynamical Love numbers that we have computed have some properties worth remarking on. Most notably, if we expand the response kernels $K$ as in (3.53) and (3.94) and then use (3.87) and (3.118) along with (3.88) and (3.119) to determine the effective worldline couplings, we see that $\lambda_{1,\ell}^{E}=\lambda_{1,\ell}^{B}$ and the coefficients of the logs in $\lambda_{2,\ell}^{E}$ are exactly the same as those in $\lambda_{2,\ell}^{B}$. [^28] In making this comparison, there is a relative normalization factor of $1/2$ in the magnetic $\lambda$ coefficients coming from the relation between the Weyl tensor and gravito-magnetic field. That is, the dynamical Love numbers in the even and odd sectors are equal—this can be understood as a consequence of Chandrasekhar’s duality.
As we mentioned in Section 3.1.2, Chandrasekhar discovered a mapping between the Regge–Wheeler and Zerilli equations [76, 77, 78]. There we used this fact as a solution-generating technique to obtain solutions in the Zerilli sector from those in the Regge–Wheeler sector. Here we note that this relation can also be read as a symmetry [7]. Specifically, both the Regge–Wheeler potential and Zerilli potential in $D=4$ can be written as
$V_{\rm RW}={\cal W}^{2}+f(r)\frac{{\rm d}{\cal W}}{{\rm d}r}+\beta,\qquad\qquad V_{\rm Z}={\cal W}^{2}-f(r)\frac{{\rm d}{\cal W}}{{\rm d}r}+\beta,$ (3.120)
with $\cal W$} as in (3.35) and $\beta\equiv-4\lambda^{2}(\lambda+1)^{2}/(9r_{s}^{2})$. This implies that the two potentials are supersymmetric partners [126]. One can further check that the following transformation is a symmetry of the action written solely in terms of the physical Regge–Wheeler and Zerilli fields [7]
$\delta\Psi_{\mathrm{Z}}=\left(\frac{\partial}{\partial r_{\star}}-{\cal W}(r)\right)\Psi_{\mathrm{RW}},\qquad\qquad\delta\Psi_{\mathrm{RW}}=\left(\frac{\partial}{\partial r_{\star}}+{\cal W}(r)\right)\Psi_{\mathrm{Z}}.$ (3.121)
This is an off-shell symmetry with a corresponding Noether current. At large distances it interchanges the electric and magnetic parts of the Weyl tensor, and so is a Schwarzschild generalization of electric-magnetic duality. [^29] Note that it keeps the background fixed, it does not map Schwarzschild into Taub–NUT. It is therefore directly a symmetry of linearized perturbations around Schwarzschild.
In the present context, this symmetry maps the worldline response operators (3.40) in the electric and magnetic sectors into each other, relating their coefficients. The astute reader, however, will note that the finite contributions in e.g., (3.87) and (3.118) differ. This would seem to be in conflict with the symmetry. However, we computed these quantities via dimensional regularization, and the symmetry (3.121) exists only in $D=4$. Since we have chosen a regularization scheme that breaks the symmetry, it is not surprising that the final answers do not respect it. Instead, the existence of the symmetry suggests that a scheme exists that preserves the symmetry. Indeed, an important consistency check is that the scheme-independent parts of the response—the imaginary dissipative response and the coefficient of the logarithms—match exactly between the two sectors. In addition, we can certainly pick the scales $\mu$ in the two sectors differently so that the finite pieces agree, which would be a scheme that preserves the duality.
4 Conclusions
In this paper, we have computed the dynamical tidal Love numbers of Schwarzschild black holes in both the even (gravito-electric) and odd (gravito-magnetic) sectors. In order to do this unambiguously, we have employed an approach that synthesizes computations in full general relativity with those done in an effective description of a black hole as a point particle. This latter description allows us to give the tidal responses a gauge-invariant definition in terms of couplings between the worldline and external gravitational fields in the Schwinger–Keldysh effective action.
In order to enable the systematics of the calculation, we have solved the equations of black hole perturbation theory in a small frequency expansion. This expansion is amenable to matching with the EFT description in an intermediate zone, bypassing many of the complexities of the problem associated to the far zone. We dealt with the nonlinearities of gravitational interactions in the EFT by utilizing the Born series approach of [75, 74], and we find also that the field profiles in the EFT must be regularized and renormalized before they can be matched to the general relativity solution.
What we find after matching to define the dynamical Love numbers of black holes is that they have several interesting properties. They are nonzero, and have a characteristic running, where the coefficient of the logarithm in all cases is $\omega r_{s}$ times the corresponding dissipative response coefficient. One way to read this relation is that the beta function for the dynamical Love numbers is determined by the dissipative response. However, this appears to become more complicated at subleading order [74]. Another way to read this relation is that the discontinuity of the conservative part of $K$ is $\beta\omega/2$ times the dissipative part of $K$, where $\beta=4\pi r_{s}$ is the inverse black hole temperature. [^30] This looks suspiciously like a fluctuation–dissipation relation. However, it is a relation involving just the causal Green’s function. (Interestingly this relation continues to hold for the most transcendental parts of the response kernel at subleading orders in frequency, at least in the examples of [74].) It would be nice to understand the features of these logarithmic terms better. Another interesting feature is that, as expected from the symmetries of black hole perturbations, there exists a scheme in which the electric and magnetic dynamical Love numbers (along with the dissipative responses) are equal.
There are a number of future directions suggested by this study that it would be interesting to pursue. Perhaps the most obvious would be to extend the computation done here to general $\ell$ and (relatedly) to higher dimension. In principle there are only technical complications to doing so, and it would be very interesting to see if the patterns seen in the four-dimensional case survive there. Relatedly, one could carry out the computation of dynamical responses to external electromagnetic fields and consider the generalization to charged black hole responses. Perhaps the most phenomenologically interesting generalization would be to compute the full dynamical Love numbers for a spinning Kerr black hole. To do this, the main advances required are technical. In particular, re-summing and regulating gravitational effects in the EFT is most simply done in dimensional regularization, but gravitational couplings to spinning black holes in higher dimensions are somewhat complicated. At leading order, we can extract a Kerr black hole’s dynamical Love number from those of a Schwarzschild black hole by exploiting the fact that a Kerr black hole is a spinning Schwarzschild black hole. The coordinate transformation between the (rotating) black hole’s rest frame and the laboratory frame captures the leading-in-spin effect (much in the same way that a Kerr black hole’s zero frequency dissipation is related to the finite frequency Schwarzschild dissipation) [46, 19]. Nevertheless the computation of the full response for Kerr is a problem that we hope to return to.
One of the most compelling reasons to study the material properties of black holes is that these quantities are in principle measurable. However, there is a practical consideration of understanding how the dynamical Love numbers actually enter into gravitational waveforms, which it would be nice to study in more detail (see [67] for a recent analysis).
An important frontier is to push the computations that we have done to subleading order. This can be done in several directions. Most straightforwardly, one could imagine computing further subleading-in-frequency corrections to black hole tidal responses. In the scalar context, this was done in [74]. In order to try to further understand the structure of black hole responses, it would be useful to amass further theoretical data. In addition to this, it has recently become clear that the nonlinear static tidal responses of black holes share many of the features of their linear counterparts [32, 34, 35, 27, 38, 29]. As such, it would be very interesting to see whether nonlinear dynamical tides have similar features to the linear dynamical tides considered here.
The true goal of precision computation of gravitational observables is insight into the structure of gravitation itself. In this regard, black holes are rare objects because of their simplicity. By studying their properties we learn how gravity organizes itself, and what features are universally shared by all objects. Most intriguingly, the remarkable simplicity of black holes has suggested the presence of symmetries, at least of black hole perturbation theory, that were not previously known. In practice, these symmetries are seen when one truncates to some near region or to zero frequency, and so one might therefore naturally imagine that these symmetries are at best approximate. However, we might take solace in the fact that the dynamical Love numbers seem to share at least a piece of the magic of their static counterparts. In particular the fact that the running of dynamical tides is related in a simple way at leading order to the dissipative response is intriguing, not least because the dissipative responses themselves can be understood as conserved charges of symmetries of a near zone [9]. One might therefore hope that there is some regularity to the pattern of black hole properties that reflects some further hidden symmetries of gravity.
The black holes in our universe are a unique window into physics in extreme environments, but black holes serve as equally valuable theoretical laboratories. There is some irony that black holes are perhaps the most mysterious objects, but nevertheless provide us with continual new lessons about gravity. We look forward to seeing what else we learn.
Acknowledgements:
Thanks to John Joseph Carrasco, Lam Hui, Riccardo Penco, and Mikhail Solon for helpful conversations AJ and DG are supported in part by DOE (HEP) Award DE-SC0025323. The work of MJR is supported through the NSF grant PHY-2309270. The research of LS has been funded by the French National Research Agency (ANR) under project ANR-24-CE31-1097-01. The work is supported by a PhD Joint Program of the International Research Center for Fundamental Scientific Discovery (IRC Discovery).
Appendix A Scalar dynamical response running
Here we compute the coefficient of the logarithmic running of the conservative dynamical response at $O(\omega^{2})$ for a scalar field at generic $\ell$, and relate it to the linear-in-$\omega$ dissipative scalar response. To this end, it is convenient to recall the relation between the Legendre functions $Q_{\ell}$ and $P_{\ell}$:
$Q_{\ell}(x)=\frac{1}{2}\log\left(\frac{x+1}{x-1}\right)P_{\ell}(x)-W_{\ell-1}(x),\qquad\quad W_{\ell-1}(x)\equiv\sum_{n=0}^{\ell-1}c_{n}(x-1)^{n},$ (A.1)
and write $P_{\ell}$ in polynomial canonical form:
$P_{\ell}(x)=\,{}_{2}F_{1}\left[\begin{array}[]{c}\ell+1\,,\,-\ell\\[-3.0pt] 1\end{array}\Big\rvert\,\frac{1-x}{2}\,\right]=\sum_{n=0}^{\ell}a_{n}(x-1)^{n},$ (A.2)
where $c_{n}$ and $a_{n}$ are the $\ell$-dependent constants
$a_{n}=\frac{(\ell+1)_{n}(\ell+1-n)_{n}}{2^{n}(n!)^{2}}\,,\qquad\qquad c_{n}=\frac{(\ell+n)!\left(\psi(\ell+1)-\psi(n+1)\right)}{2^{n}(\ell-n)!(n!)^{2}}\,$, (A.3)
and where $(z)_{n}$ is the Pochhammer symbol, defined by $(z)_{n}\equiv\Gamma(z+n)/\Gamma(z)$.
Let us start by focusing on the first integral in (2.21). Using (A.1), we find
$\int^{x}{\rm d}y\frac{(1+y)^{3}}{y-1}P_{\ell}(y)Q_{\ell}(y)=\int^{x}\frac{{\rm d}y}{2}\frac{(1+y)^{3}}{y-1}P_{\ell}(y)^{2}\log\left(\frac{y+1}{y-1}\right)-\int^{x}{\rm d}y\frac{(1+y)^{3}}{y-1}P_{\ell}(y)W_{\ell-1}(y).$ (A.4)
Consider the first integral in this expression: from the definition of $P_{\ell}$, (A.2), the first integral in (A.4) boils down to a sum of integrals, each of which has the following form:
$I_{n}=\int^{x}{\rm d}y\frac{(1+y)^{3}}{y-1}(y-1)^{n}\log\left(\frac{y+1}{y-1}\right)$ (A.5)
for some integer $n$. Recall that our modest goal here is to simply extract the $\ell$-dependent coefficient of the term that in (2.21) scales as $\log(x)/x^{\ell+1}$ at large $x$; all terms with different scaling in xx—or with no logarithm—will be ignored in what follows. For instance, explicitly evaluating the integral (A.5) for $n=0$ shows that no such $\log(x)/x^{\ell+1}$ is present, so it can be disregarded. Let us thus assume $n>0$. From a straightforward integration by parts, we can write:
$I_{n}$ $=\int^{x}{\rm d}y\left[\frac{{\rm d}}{{\rm d}y}\int^{y}{\rm d}z\frac{(1+z)^{3}}{z-1}(z-1)^{n}\right]\log\left(\frac{y+1}{y-1}\right)$ (A.6)
$=p_{n}(x)\log\left(\frac{x+1}{x-1}\right)+2\int^{x}{\rm d}y\frac{p_{n}(y)}{(y+1)(y-1)},$
where we have introduced the polynomials
$p_{n}(x)\equiv\left(\frac{(x-1)^{3}}{n+3}+\frac{6(x-1)^{2}}{n+2}+\frac{12(x-1)}{n+1}+\frac{8}{n}\right)(x-1)^{n}.$ (A.7)
Using the integral representation of the hypergeometric function, we can write each piece contributing to the last term of (A.6) as
$\int^{x}_{1}{\rm d}y\frac{(y-1)^{m}}{y+1}=\frac{(x-1)^{m+1}}{2(m+1)}\,{}_{2}F_{1}\left[\begin{array}[]{c}1\,,\,m+1\\[-3.0pt] m+2\end{array}\Big\rvert\,\frac{1-x}{2}\,\right]\,$, (A.8)
with $m$ a non-negative integer. [^31] In (A.8), the lower bound in the integral (which we chose to be $1$ for convenience) is completely immaterial, as this simply amounts to a redefinition of the integration constants in (2.21). Note that no $\log(x)/x^{\ell+1}$ scaling is contained in the homogeneous solutions of (2.17). From the properties of the hypergeometric function, each of these terms is of the form $\text{pol}_{1}(x)+\text{pol}_{2}(x)\times\log(x)$, at large $x$, where $\text{pol}_{1}(x)$ and $\text{pol}_{2}(x)$ are generic ordinary polynomials with positive powers of $x$. In conclusion, no $\log(x)/x^{\ell+1}$ term results from the $x\rightarrow\infty$ expansion of the $I_{n}$ integrals (A.6), for all $n$.
The same conclusion is trivially true for the second integral in (A.4), since both $P_{\ell}(y)$ and $W_{\ell-1}(y)$ are polynomials in (y1)(y-1). Therefore, there is no $\log(x)/x^{\ell+1}$ type of term resulting from the first integral in (2.21). Let us then focus on the second integral in (2.21):
$\frac{B}{4}\int^{x}{\rm d}y\frac{(1+y)^{3}}{y-1}P_{\ell}(y)^{2}.$ (A.9)
Using the series representation (A.2) of the Legendre polynomials, the only term of the sum (A.2) that can generate a logarithm in (A.9) is the one with the $a_{0}$ coefficient. All the other terms in the series produce just polynomials in $x$ after integration. Explicitly,
$\frac{B}{4}\int^{x}{\rm d}y\frac{(1+y)^{3}}{y-1}P_{\ell}(y)^{2}=2B\log(x-1)+\text{pol}(x).$ (A.10)
From the large-$x$ expansion of the $Q_{\ell}(x)$ functions,
$Q_{\ell}(x)\underset{x\to\infty}{\sim}\frac{2^{\ell}(\ell!)^{2}}{(2\ell+1)!}\frac{1}{x^{\ell+1}},$ (A.11)
combined with (A.10), we find that the second integral in (2.21) produces at infinity
$\lim_{x\to\infty}\frac{B}{4}\int^{x}{\rm d}y\frac{(1+y)^{3}}{y-1}P_{\ell}(y)^{2}Q_{\ell}(x)\supset 2B\frac{2^{\ell}(\ell!)^{2}}{(2\ell+1)!}\frac{\log(x)}{x^{\ell+1}}+\cdots\,$, (A.12)
where we omitted in $\cdots$ all contributions that scale differently from the one that we wrote explicitly. [^32] The omitted terms are not necessarily subleading—they are not necessarily suppressed by more powers of $x$. Putting everything together in (2.10), and once again focusing only on the asymptotic falloffs $x^{\ell}$ and $x^{-\ell-1}$, the expanded intermediate-zone solution, up to order $\epsilon^{2}$, reads
$\lim_{x\to\infty}\phi_{\text{IZ}}\supset B\frac{(2\ell)!}{2^{\ell}(\ell!)^{2}}x^{\ell}\left(1+O(\epsilon)\right)+B\frac{2^{\ell}(\ell!)^{2}}{(2\ell+1)!}\frac{1}{x^{\ell+1}}\left(2i\epsilon+2\epsilon^{2}\log(x)+O(\epsilon^{3})\right)+\dots$ (A.13)
where we used that $P_{\ell}(x)\underset{x\to\infty}{\sim}\frac{2^{-\ell}(2\ell)!}{(\ell!)^{2}}x^{\ell}$. From the definitions of $x$ and $\epsilon$ (see (2.8)), the ratio of the two falloffs $r^{-\ell-1}$ and $r^{\ell}$ (involving only the logarithmic term at order $\epsilon^{2}$) is
$\kappa_{\ell}=\bigg(i\omega r_{s}+\omega^{2}r_{s}^{2}\log\left(\frac{r}{r_{s}}\right)\bigg)\frac{(\ell!)^{4}}{(2\ell)!(2\ell+1)!}+\cdots$ (A.14)
in agreement with previous results [13, 55]. Note that when expressed as $\log r_{s}$, the relative ratio is precisely 1-1, so that the dynamical Love number running is (minus) the dissipative response.
Appendix B Far zone
Here we study the far zone approximation of the Klein–Gordon equation (2.3) (see also [88]). Before taking the far-zone limit, it is convenient to redefine the field in a way that removes the first-derivative term from the equation. We introduce
$\xi_{\ell}(r)\equiv\sqrt{\Delta(r)}\phi_{\ell}(r),$ (B.1)
where ${\Delta}(r)\equiv{r(r-r_{s})}$. In terms of $\xi$, (2.3) then becomes
$\partial_{r}^{2}\xi_{\ell}(r)+\left(\frac{\omega^{2}r^{4}}{\Delta^{2}}-\frac{\ell(\ell+1)}{\Delta}-\frac{2\Delta\Delta^{\prime\prime}-(\Delta^{\prime})^{2}}{4\Delta^{2}}\right)\xi_{\ell}(r)=0.$ (B.2)
A possible way to obtain a far-zone solution would be to expand the potential in (B.2) at large $r$, keeping terms up to the desired order. [^33] For instance, up to the order $O(r^{-2})$, one gets $\partial_{r}^{2}\xi_{\ell}(r)+\left(\omega^{2}+\frac{2r_{s}\omega^{2}}{r}-\frac{\ell(\ell+1)-3\omega^{2}r_{s}^{2}}{r^{2}}\right)\xi_{\ell}(r)=0.$ (B.3) Neglecting the $O(\omega^{2}/r^{2})$ term in the potential, as in [88], yields precisely the Coulomb wave equation. Although the solutions to the equation expanded this way reproduce asymptotically the standard Coulomb functions, care must be taken to ensure that the range of validity of the approximation extends sufficiently into the intermediate zone, so that matching with the perturbative solution in Section 2.1.2 can be carried out. In particular, we need a solution that remains valid from $\omega r\sim O(1)$ all the way up to $\omega r\to\infty$. To this end, we will expand assuming $r_{s}\ll r$ but without imposing any hierarchy between $\omega$ and $r$.
Concretely, it is convenient to define the quantities
$z\equiv\omega r,\qquad\qquad\epsilon\equiv\omega r_{s}\,.$ (B.4)
We then expand equation (B.2) for small $\frac{\epsilon}{z}=\frac{r_{s}}{r}$, up to order $O(\frac{\epsilon^{2}}{z^{2}})$, without making any assumption about the variable $z$, which can range from $O(1)$ to $\infty$. This yields
$\partial_{z}^{2}\xi_{\ell}+\left(1-\frac{\ell(\ell+1)}{z^{2}}\right)\xi_{\ell}=-\epsilon\left(\frac{2}{z}-\frac{\ell(\ell+1)}{z^{3}}\right)\xi_{\ell}-\epsilon^{2}\left(\frac{3}{z^{2}}+\frac{\frac{1}{4}-\ell(\ell+1)}{z^{4}}\right)\xi_{\ell}+O(\epsilon^{3})\,.$ (B.5)
In the following, we solve this equation to linear order in $\epsilon$. We then use the result to compute the scattering amplitude of a scalar field off the black hole, and finally compare it with the solution obtained from the Mano–Suzuki–Takasugi (MST) formalism [72, 73].
B.1 Far-zone scalar solution
We look for a perturbative solution in $\epsilon$, i.e., we expand the field as
$\xi_{\ell}(r(z))=\xi_{\ell}^{(0)}(z)+\epsilon\xi_{\ell}^{(1)}(z)+\epsilon^{2}\xi_{\ell}^{(2)}(z)+\cdots.$ (B.6)
At zeroth order, $\xi_{\ell}^{(0)}$ solves (B.5) with $\epsilon=0$. The two independent solutions can be expressed in terms of the standard Coulomb functions as
$\xi_{\ell}^{(0)}(z)=c_{1}^{(0)}F_{\ell}(0,z)+c_{2}^{(0)}G_{\ell}(0,z),$ (B.7)
with one of the parameters set to zero. Using the well-known Coulomb wave functions’ asymptotics,
$F_{\ell}(\eta,z)$ $\underset{z\to 0}{\sim}C_{\ell}(\eta)z^{\ell+1}\underset{\eta=0}{=}\frac{1}{(2\ell+1)!!}z^{\ell+1},$ (B.8)
$G_{\ell}(\eta,z)$ $\underset{z\to 0}{\sim}\frac{z^{-\ell}}{(2\ell+1)C_{\ell}(\eta)}\underset{\eta=0}{=}\frac{(2\ell+1)!!}{2\ell+1}z^{-\ell}=(2\ell-1)!\,z^{-\ell},$
with the constant parameters
$C_{\ell}(\eta)\equiv\frac{2^{\ell}\operatorname{e}^{-\pi\eta/2}|\Gamma(\ell+1+i\eta)|}{(2\ell+1)!},\qquad(2\ell+1)!!=\frac{(2\ell+1)!}{2^{\ell}\ell!},$ (B.9)
we can compute the $z\to 0$ limit of the FZ solution:
$\phi_{\ell,\text{FZ}}^{(0)}\underset{z\to 0}{\sim}\frac{c_{1}^{(0)}}{(2\ell+1)!!}\omega z^{\ell}+c_{2}^{(0)}\frac{(2\ell+1)!!}{2\ell+1}\omega z^{-\ell-1},$ (B.10)
where we used the definition (B.1) and have expanded $1/\sqrt{\Delta}$ in powers of $\epsilon$ before taking the small-$z$ limit, [^34] To match with the intermediate-zone solution, we require $\epsilon\ll z$.
$\frac{1}{\sqrt{\Delta}}$ $=\frac{\omega}{\left(z(z-\epsilon)\right)^{1/2}}$ (B.11)
$=\frac{\omega}{z}\left(1+\frac{\epsilon}{2z}+\frac{3\epsilon^{2}}{8z^{2}}\right)+O\left(\frac{\epsilon^{3}}{z^{3}}\right).$
Note that we are doing an expansion in $\epsilon$, and not literally in $\omega$. Hence, the integration constants $c_{1}^{(0)}$ and $c_{2}^{(0)}$ can (and in general will) have a nontrivial dependence on $\omega$. In particular, the FZ solution at a specific order in $\epsilon$ may involve different orders in $\omega$.
Matching at leading order: Comparing with the intermediate-zone solution at order $O(\epsilon^{0})$, which takes the form
$\phi_{\ell,\text{IZ}}^{(0)}=BP_{\ell}\left(2r/r_{s}-1\right)\underset{r\to\infty}{\sim}B\frac{(2\ell)!}{(\ell!)^{2}}\left(\frac{r}{r_{s}}\right)^{\ell}+\cdots,$ (B.12)
and matching at leading order in $r/r_{s}$ yields
$c_{2}^{(0)}=0,\qquad\qquad c_{1}^{(0)}=\frac{B}{\omega^{\ell+1}r_{s}^{\ell}}\frac{(2\ell)!(2\ell+1)!!}{(\ell!)^{2}}.$ (B.13)
Matching at sub-leading order: Let us now consider the matching with the intermediate zone at linear order in $\epsilon$. The $O(\epsilon)$ term on the right-hand side of (B.5) plays the role of a source, which we evaluate on the $O(\epsilon^{0})$ solution $\xi_{\ell}^{(0)}$. Using standard Green’s function methods, we find the following general solution [^35] Recall the Wronskian $W[F_{\ell}(0,z),G_{\ell}(0,z)]\equiv F_{\ell}(0,z)\partial_{z}G_{\ell}(0,z)-G_{\ell}(0,z)\partial_{z}F_{\ell}(0,z)=-1$.
$\xi_{\ell}^{(1)}(z)=\left(c_{1}^{(1)}-I_{F}(z)\right)F_{\ell}(0,z)+\left(c_{2}^{(1)}+I_{G}(z)\right)G_{\ell}(0,z),$ (B.14)
where we defined the indefinite integrals
$I_{F}(z)$ $\equiv\int^{z}{\rm d}t\,G_{\ell}(0,t)\xi_{\ell}^{(0)}(t)\left(\frac{2}{t}-\frac{\ell(\ell+1)}{t^{3}}\right),$ (B.15)
$I_{G}(z)$ $\equiv\int^{z}{\rm d}t\,F_{\ell}(0,t)\xi_{\ell}^{(0)}(t)\left(\frac{2}{t}-\frac{\ell(\ell+1)}{t^{3}}\right).$
These integrals can be computed in terms of (generalized) hypergeometric functions ${{}_{p}}F_{q}$ as
$I_{F}(z)$ $=\frac{c_{1}^{(0)}}{2\ell+1}\bigg(\frac{\ell(\ell+1)}{z}\,{}_{1}F_{2}\bigg[\begin{array}[]{c}-\tfrac{1}{2}\\ \tfrac{1}{2}-\ell\,,\,\tfrac{3}{2}+\ell\end{array}\bigg\rvert-z^{2}\,\bigg]+2z\,{}_{2}F_{3}\bigg[\begin{array}[]{c}\tfrac{1}{2}\,,\,\tfrac{1}{2}\\ \tfrac{3}{2}\,,\,\tfrac{1}{2}-\ell\,,\,\tfrac{3}{2}+\ell\end{array}\bigg\rvert-z^{2}\,\bigg]\bigg)\,$ (B.16)
$I_{G}(z)$ $=\frac{-\sqrt{\pi}c_{1}^{(0)}z^{2\ell}}{4}\bigg((\ell+1)!\,{}_{1}\tilde{F}_{2}\bigg[\!\begin{array}[]{c}\ell\\[-2.0pt] \tfrac{3}{2}+\ell\,,\,2+2\ell\end{array}\bigg\rvert-z^{2}\,\bigg]$
$\hskip 85.35826pt-2(\ell!)^{2}z^{2}{}_{2}\tilde{F}_{3}\bigg[\!\begin{array}[]{c}\ell+1\,,\,\ell+1\\[-1.0pt] \tfrac{3}{2}+\ell\,,\,2+\ell\,,\,2+2\ell\end{array}\bigg\rvert-z^{2}\,\bigg]\bigg)\ ,$
where $\tilde{F}$ denotes the regularized function, defined by
$_{p}\tilde{F}_{q}\left[\begin{array}[]{c}a_{1}\,,\,\cdots\,,\,a_{p}\\[-2.0pt] b_{1}\,,\,\cdots\,,\,b_{q}\end{array}\bigg\rvert\,x\,\right]\equiv\frac{\,{}_{p}F_{q}\left[\begin{array}[]{c}a_{1}\,,\,\cdots\,,\,a_{p}\\[-2.0pt] b_{1}\,,\,\cdots\,,\,b_{q}\end{array}\bigg\rvert\,x\,\right]}{\Gamma(b_{1})\cdots\Gamma(b_{q})}.$ (B.17)
To match with the intermediate-zone solution, we need the small-$z$ expansion of the solution (B.14). First, let us use the series representation of the generalized hypergeometric functions,
$_{p}F_{q}\left[\begin{array}[]{c}a_{1}\,,\,\cdots\,,\,a_{p}\\[-2.0pt] b_{1}\,,\,\cdots\,,\,b_{q}\end{array}\bigg\rvert\,x\,\right]=\sum_{n=0}^{\infty}\frac{(a_{1})_{n}\cdots(a_{p})_{n}}{(b_{1})_{n}\cdots(b_{q})_{n}}\frac{x^{n}}{n!},$ (B.18)
which we use to write $I_{F}(z)$ and $I_{G}(z)$ as
$I_{F}(z)$ $=\frac{c_{1}^{(0)}}{2\ell+1}\sum_{k=0}^{\infty}\left[\ell(\ell+1)f_{1,k}z^{2k-1}+2f_{2,k}z^{2k+1}\right]$ (B.19)
$=\frac{c_{1}^{(0)}}{2\ell+1}\left(\sum_{k=1}^{\infty}\underbrace{\left[\ell(\ell+1)f_{1,k}+2f_{2,k-1}\right]}_{\zeta_{k}^{\ell}}z^{2k-1}+\frac{\ell(\ell+1)f_{1,0}}{z}\right),$
and
$I_{G}(z)=$ $-\frac{\sqrt{\pi}c_{1}^{(0)}}{4}z^{2\ell}\sum_{k=0}^{\infty}\left[(\ell+1)!g_{1,k}z^{2k}-2(\ell!)^{2}g_{2,k}z^{2k+2}\right]$ (B.20)
$=$ $-\frac{\sqrt{\pi}c_{1}^{(0)}}{4}z^{2\ell}\left(\sum_{k=1}^{\infty}\underbrace{\left[(\ell+1)!g_{1,k}-2(\ell!)^{2}g_{2,k-1}\right]}_{\gamma_{k}^{\ell}}z^{2k}+(\ell+1)!g_{1,0}\right),$
where we have introduced the combinations
$f_{1,k}=\frac{(-1)^{k}(-\frac{1}{2})_{k}}{(\frac{1}{2}-\ell)_{k}(\frac{3}{2}+\ell)_{k}k!},\qquad f_{2,k}=\frac{(-1)^{k}((\frac{1}{2})_{k})^{2}}{(\frac{3}{2})_{k}(\frac{1}{2}-\ell)_{k}(\frac{3}{2}+\ell)_{k}k!},$ (B.21)
$g_{1,k}=\frac{(-1)^{k}(\ell)_{k}}{\Gamma(\ell+k+\frac{3}{2})\Gamma(2\ell+k+2)k!},\qquad g_{2,k}=\frac{(-1)^{k}((\ell+1)_{k})^{2}}{\Gamma(\ell+k+\frac{3}{2})\Gamma(2\ell+k+2)\Gamma(\ell+k+2)k!}.$ (B.22)
To expand the Coulomb functions, we can then use their relations to spherical Bessel functions, i.e.,
$F_{\ell}(0,z)=zj_{\ell}(z),\qquad\qquad G_{\ell}(0,z)=-zy_{\ell}(z),$ (B.23)
where $j_{\ell}$ and $y_{\ell}$ admit the following series representation:
$j_{\ell}(z)=z^{\ell}\sum_{n=0}^{\infty}j_{n}^{\ell}z^{2n},\qquad y_{\ell}(z)=-z^{-\ell-1}\sum_{n=0}^{\infty}y_{n}^{\ell}z^{2n},$ (B.24)
with the parameters
$j_{n}^{\ell}$ $=\frac{(-1)^{n}}{2^{n}n!(2\ell+2n+1)!!},$ (B.25)
$y_{n}^{\ell}$ $=\Theta(\ell-n)\frac{(2\ell-2n-1)!!}{2^{n}n!}+\Theta(n-(\ell+1))\frac{(-1)^{\ell+n}}{2^{n}n!(2n-2\ell-1)!!},$
where $\Theta(k)=1$ for $\mathbb{N}\ni k\geq 0$, and $\Theta(k)=0$ otherwise. Finally, the expansion of (B.14) for small $z$ reads
$\xi_{\ell}^{(1)}(z)=$ $\left[c_{1}^{(1)}z^{\ell+1}-\frac{c_{1}^{(0)}}{2\ell+1}\left(\sum_{k=1}^{\infty}\zeta_{k}^{\ell}z^{\ell+2k}+\ell(\ell+1)z^{\ell}\right)\right]\sum_{n=0}^{\infty}j_{n}^{\ell}z^{2n}$ (B.26)
$+\left[c_{2}^{(1)}z^{-\ell}-\frac{c_{1}^{(0)}\sqrt{\pi}}{4}\left(\sum_{k=1}^{\infty}\gamma_{k}^{\ell}z^{\ell+2k}+(\ell+1)!g_{1,0}z^{\ell}\right)\right]\sum_{n=0}^{\infty}y_{n}^{\ell}z^{2n}.$
Plugging this into the definition of $\phi^{(1)}_{\ell,\text{FZ}}=\frac{\omega}{z}(\xi^{(1)}_{\ell}+\frac{1}{2z}\xi^{(0)}_{\ell})$, we obtain the following $z\rightarrow 0$ behavior for $\phi_{\ell,\text{FZ}}^{(1)}$ [^36] We use $y_{0}^{\ell}=(2\ell-1)!!=\frac{(2\ell+1)!!}{2\ell+1}$ to simplify the expressions.
$\phi_{\ell,\text{FZ}}^{(1)}(z)\underset{z\to 0}{\sim}$ $\,\omega\Bigg[c_{2}^{(1)}z^{-\ell-1}\left(y_{0}^{\ell}+\sum_{n=1}^{\ell}y_{n}^{\ell}z^{2n}\right)-c_{1}^{(0)}z^{\ell-1}\left(\frac{\ell(\ell+1)}{2\ell+1}j_{0}^{\ell}+\frac{\sqrt{\pi}(\ell+1)!}{4}g_{1,0}y_{0}^{\ell}-\frac{1}{2(2\ell+1)!!}\right)$
$\quad+c_{1}^{(1)}j_{0}^{\ell}z^{\ell}+O(z^{\ell+1})\Bigg]$
$=$ $\,\,\omega c_{2}^{(1)}\frac{(2\ell+1)!!}{2\ell+1}z^{-\ell-1}\left(1+\sum_{n=1}^{\ell-1}\frac{y_{n}^{\ell}}{y_{0}^{\ell}}z^{2n}\right)-\omega c_{1}^{(0)}\frac{\ell}{2(2\ell+1)!!}z^{\ell-1}$
$+\frac{\omega c_{1}^{(1)}}{(2\ell+1)!!}z^{\ell}+O(z^{\ell+1}).$ (B.27)
The integration constants $c_{1}^{(1)}$ and $c_{2}^{(1)}$ can be determined by comparing $\phi_{\ell,\text{FZ}}^{(1)}$ with the intermediate-zone solution:
$\phi_{\ell,\text{IZ}}^{(1)}$ $=b_{1}^{(1)}P_{\ell}\left({2r}/{r_{s}}-1\right)+b_{2}^{(1)}Q_{\ell}({2r}/{r_{s}}-1)$ (B.28)
$\!\!\!\!\underset{\frac{r}{r_{s}}\to\infty}{\sim}b_{1}^{(1)}\frac{(2\ell)!}{(\ell!)^{2}}\left(\frac{r}{r_{s}}\right)^{\ell}+b_{2}^{(1)}\frac{(\ell!)^{2}}{2(2\ell+1)!}\left(\frac{r}{r_{s}}\right)^{-\ell-1}+O\left(\left({r}/{r_{s}}\right)^{\ell-1}\right),$
where $b_{1}^{(1)}$ and $b_{2}^{(1)}$ are the intermediate-zone integration constants, which are fixed in terms of the amplitude $B$ via (2.16). From the comparison of the $r^{-\ell-1}$ and $r^{\ell}$ falloffs, we find
$c_{1}^{(1)}=\frac{b_{1}^{(1)}}{\omega^{\ell+1}r_{s}^{\ell}}\frac{(2\ell)!(2\ell+1)!!}{(\ell!)^{2}},\qquad\quad c_{2}^{(1)}=\omega^{\ell}r_{s}^{\ell+1}b_{2}^{(1)}\frac{(\ell!)^{2}}{2(2\ell)!(2\ell+1)!!}.$ (B.29)
Once these coefficients are fixed, it is straightforward to verify that all the remaining terms in (B.27) that contribute at the given order match automatically. First, note that upon substituting (B.29) into (B.27), the subleading terms proportional to $c_{2}^{(1)}$ are at least of order $O(\omega^{2})$. Since we are only interested in terms up to first order in $\omega$, we can discard them and focus on comparing the term proportional to $c_{1}^{(0)}$. Substituting the expression for $c_{1}^{(0)}$ from (B.13) into $\epsilon\phi_{\ell,\text{FZ}}^{(1)}$, one finds
$\epsilon\omega c_{1}^{(0)}\frac{\ell}{2(2\ell+1)!!}z^{\ell-1}=-B\frac{\ell(2\ell)!}{2(\ell!)^{2}}\left(\frac{r}{r_{s}}\right)^{\ell-1}.$ (B.30)
This contribution is frequency independent. It must therefore match the subleading contribution in $\phi_{\ell,\text{IZ}}^{(0)}$ as $r\to\infty$. To check this, we can use the series representation of the Legendre polynomials,
$P_{\ell}\left(\frac{2r}{r_{s}}-1\right)=\sum_{n=0}^{\ell}\sum_{k=0}^{n}\alpha_{nk}\left(\frac{r}{r_{s}}\right)^{n-k},\qquad\alpha_{nk}=(-1)^{k}\binom{n}{k}\frac{(\ell+n)!}{(\ell-n)!(n!)^{2}},$ (B.31)
to express the zeroth-order intermediate-zone solution as
$\phi_{\ell,\text{IZ}}^{(0)}=BP_{\ell}\left(\frac{2r}{r_{s}}-1\right)\underset{\frac{r}{r_{s}}\to\infty}{\sim}B\alpha_{\ell 0}\left(\frac{r}{r_{s}}\right)^{\ell}+B\left(\alpha_{(\ell-1)0}+\alpha_{\ell 1}\right)\left(\frac{r}{r_{s}}\right)^{\ell-1}+O\left(\left(r/r_{s}\right)^{\ell-2}\right),$ (B.32)
where
$\alpha_{(\ell-1)0}+\alpha_{\ell 1}=-\frac{\ell}{2}\alpha_{\ell 0}=-\frac{\ell(2\ell)!}{2(\ell!)^{2}}.$ (B.33)
As expected, this reproduces the $O(\omega^{0}r^{\ell-1}$) part of $\phi_{\ell,\text{FZ}}^{(1)}$.
B.2 Comparison with the MST solution
The MST formalism is a systematic approach to compute black hole scattering amplitudes order by order in frequency, constructed by expanding the field solution as an infinite series of hypergeometric functions (see [72, 73] for details, and also [13, 59, 95, 19]). The goal of this section is to show that the far-zone solution obtained above correctly reproduces the MST result for the scalar scattering amplitude at the considered order of approximation.
For this purpose, it is convenient to first solve a slightly different, more general problem, which we will later connect to the far zone above and to MST. Unlike before, let us not make any assumption on $\epsilon$, which can now take any value, and solve (B.5) for large $z$. One could call this the Very Far Zone. Truncated at order $O(z^{-2})$, (B.5) reduces to a standard Coulomb wave equation
$\partial_{z}^{2}\xi_{\ell}(z)+\left(1+\frac{2\epsilon}{z}-\frac{\ell(\ell+1)-3\epsilon^{2}}{z^{2}}\right)\xi_{\ell}(z)=0,$ (B.34)
whose general solutions can be written as
$\xi_{\ell,\epsilon}(z)=d_{1}F_{L}(-\epsilon,z)+d_{2}G_{L}(-\epsilon,z),$ (B.35)
where $d_{1}$ and $d_{2}$ are generic integration constants and $L$ a parameter defined by
$L\equiv$ $-\frac{1}{2}+\frac{1}{2}\sqrt{(2\ell+1)^{2}-4\epsilon^{2}\delta\ell}.$ (B.36)
Note that the FZ solution $\xi_{\ell}$ derived in the previous section can be recovered by taking the $\epsilon\to 0$ limit of (B.35).
The scalar scattering amplitude $\mathcal{A}$ is given by the ratio of the coefficients of the reflected and ingoing waves [127, 56],
$1-i\mathcal{A}=(-1)^{\ell+1}\frac{\phi_{\ell}^{\text{ref}}}{\phi_{\ell}^{\text{in}}},$ (B.37)
where $\phi_{\ell}^{\text{ref}}$ and $\phi_{\ell}^{\text{in}}$ are defined by
$\underset{\omega r\to\infty}{\lim}\phi_{\ell,\text{FZ}}=\underset{\omega r\to\infty}{\lim}\frac{\xi_{\ell,\epsilon}(\omega r)}{r}=\phi_{\ell}^{\text{ref}}\frac{\operatorname{e}^{+i\omega r_{\star}}}{r}+\phi_{\ell}^{\text{in}}\frac{\operatorname{e}^{-i\omega r_{\star}}}{r},$ (B.38)
where we introduced the tortoise coordinate $r_{\star}$, defined as
$\omega r_{\star}=z+\epsilon\log\left(\frac{z}{\epsilon}-1\right)\underset{z\to\infty}{\sim}z+\epsilon\log(z)-\epsilon\log(\epsilon).$ (B.39)
Using, the large-$z$ limit of the Coulomb wave functions
$F_{L}(\eta,z)$ $\underset{z\to\infty}{\sim}\sin\left(z-\eta\log(2z)-\frac{L\pi}{2}+\arg(\Gamma(L+1+i\eta)\right),$ (B.40)
$G_{L}(\eta,z)$ $\underset{z\to\infty}{\sim}\cos\left(z-\eta\log(2z)-\frac{L\pi}{2}+\arg(\Gamma(L+1+i\eta)\right),$
we find the following expressions for the fall-off coefficients, [^37] We used the following relations for the $\Gamma$-functions with complex argument: $\Gamma(a^{*})=\Gamma(a)^{*},\quad\quad\arg\Gamma(a)=-\frac{i}{2}\log\frac{\Gamma(a)}{\Gamma(a^{*})}.$ (B.41)
$\phi_{\ell}^{\text{ref}}$ $=\frac{\operatorname{e}^{-i\frac{\pi L}{2}}}{2}\operatorname{e}^{i\epsilon\log(2\epsilon)}\left(d_{2}-id_{1}\right)\sqrt{\frac{\Gamma(L-i\epsilon+1)}{\Gamma(L+i\epsilon+1)}},$ (B.42)
$\phi_{\ell}^{\text{in}}$ $=\frac{\operatorname{e}^{i\frac{\pi L}{2}}}{2}\operatorname{e}^{-i\epsilon\log(2\epsilon)}\left(d_{2}+id_{1}\right)\sqrt{\frac{\Gamma(L+i\epsilon+1)}{\Gamma(L-i\epsilon+1)}},$
which finally yields
$(-1)^{L+1}\frac{\phi^{\text{ref}}}{\phi^{\text{in}}}=-\frac{\Gamma(L-i\epsilon+1)}{\Gamma(L+i\epsilon+1)}\operatorname{e}^{2i\epsilon\log(2\epsilon)}\frac{d_{2}-id_{1}}{d_{2}+id_{1}}.$ (B.43)
In order to compare (B.43) with the MST result, we take the expression for the same quantity obtained via the MST approach (see Appendix E of [19] for details). In the notation of [19], and for small $\epsilon$,
$(-1)^{L+1}\frac{R^{(\text{ref.})}}{R^{(\text{in.})}}=$ $\frac{A_{-}^{\nu}}{A_{+}^{\nu}}\operatorname{e}^{i\epsilon\left(2\log\epsilon-(1-\kappa)\right)}\frac{K_{\nu}-i\operatorname{e}^{i\pi\nu}K_{-\nu-1}}{K_{\nu}+i\operatorname{e}^{-i\pi\nu}\frac{\sin(\nu+i\epsilon)}{\sin(\nu-i\epsilon)}K_{-\nu-1}}$ (B.44)
$=$ $\frac{A_{-}^{\nu}}{A_{+}^{\nu}}\operatorname{e}^{i\epsilon\left(2\log\epsilon-(1-\kappa)\right)}\bigg[1-i\operatorname{e}^{i\pi L}\left(1+\operatorname{e}^{-2i\pi L}\frac{\sin(\pi(L+i\epsilon)}{\sin(\pi(L-i\epsilon)}\right)\frac{K_{-\nu-1}}{K_{\nu}}\bigg]+O(\epsilon^{2L+3}),$
which we adapted here to the case of a Schwarzschild black hole. In particular, $\nu=L+\omega^{2}\nu_{2}$ and $L$ is the analytic continuation of $\ell$ to non-integer values. Note also that $\frac{K_{-\nu-1}}{K_{\nu}}=O(\epsilon^{2L+1})$. The expression (B.44) has been shown to reflect a near-far factorization of the scattering amplitude [127, 56]. In particular, the term in square brackets arises from the physics of the near zone and its matching to the far region, while the prefactor captures the far-zone contribution. This factorization is not yet apparent from (B.43). To make it manifest, it is useful to re-express the ratio $\frac{d_{2}-id_{1}}{d_{2}+id_{1}}$ by extracting the $\epsilon$-dependence from $d_{1}$ and $d_{2}$. To do so, we first match the asymptotic solution $\xi_{\ell,\epsilon}$ in (B.35) to the FZ solution $\xi_{\ell}$ derived in the previous section for small $\epsilon$. The $d_{1,2}$ constants become then linear combinations of the FZ integration constants $c_{1,2}^{(n)}$, where $n$ is the order in $\epsilon$ we are considering. Given the form of the constants at linear order in $\epsilon$ (see (B.13) and (B.29)), it is convenient to parametrize $d_{1}$ and $d_{2}$ as
$d_{1}=\epsilon^{\ell}d_{1}^{+}+\frac{d_{1}^{-}}{\epsilon^{\ell+1}},\qquad\qquad d_{2}=\epsilon^{\ell}d_{2}^{+}+\frac{d_{2}^{-}}{\epsilon^{\ell+1}},$ (B.45)
where $d_{1,2}^{\pm}$ have a polynomial dependence on $\epsilon$. Plugging these into (B.43), one finds
$(-1)^{L+1}\frac{\phi^{\text{ref}}}{\phi^{\text{in}}}=-\frac{\Gamma(L-i\epsilon+1)}{\Gamma(L+i\epsilon+1)}e^{2i\epsilon\log(2\epsilon)}\frac{d_{2}^{-}-id_{1}^{-}}{d_{2}^{-}+id_{1}^{-}}\left[1+2i\frac{d_{2}^{+}d_{1}^{-}-d_{2}^{-}d_{1}^{+}}{(d_{2}^{-})^{2}+(d_{1}^{-})^{2}}\epsilon^{2\ell+1}+O(\epsilon^{4\ell+2})\right],$ (B.46)
which we expressed so as to recover the form of the near-far factorization of (B.44). We can now perform the comparison by separately expanding the near-zone and far-zone terms at linear order in $\epsilon$. First, we fix the constants $d_{1,2}^{(1)}$ at this order by matching the asymptotic solution with the one from the previous section. Imposing
$\underset{z\to\infty}{\lim}\xi_{\ell}(z)-\underset{\epsilon\to 0}{\lim}\left(\underset{z\to\infty}{\lim}\xi_{\ell,\epsilon}(z)\right)=O(\epsilon^{2}),$ (B.47)
where
$\underset{z\to\infty}{\lim}\xi_{\ell}(z)=$ $\,\epsilon\cos\left(\frac{\pi\ell}{2}-z\right)\left(c_{2}^{(1)}+c_{1}^{(0)}\left(\log(2z)-\psi(\ell+1)-\frac{1}{2}\right)\right)$ (B.48)
$-\left(c_{1}^{(0)}+c_{1}^{(1)}\epsilon+c_{1}^{(0)}\frac{\pi}{2}\epsilon\right)\sin\left(\frac{\pi\ell}{2}-z\right)+O(\epsilon^{2}),$
and
$\underset{\epsilon\to 0}{\lim}\left(\underset{z\to\infty}{\lim}\xi_{\ell,\epsilon}(z)\right)=$ $\,\cos\left(\frac{\pi L}{2}-z\right)(-d_{1}\epsilon\psi(L+1)+d_{1}\epsilon\log(2z)+d_{2})$ (B.49)
$-\sin\left(\frac{\pi L}{2}-z\right)(d_{1}+d_{2}\epsilon\psi(L+1)-d_{2}\epsilon\log(2z))+O(\epsilon^{2}),$
and using $L=\ell+O(\epsilon^{2})$, we obtain
$d_{1}^{-}$ $=\epsilon^{\ell+1}\left(c_{1}^{(0)}+\left(c_{1}^{(1)}+c_{1}^{(0)}\frac{\pi}{2}\right)\epsilon+O(\epsilon^{2})\right),$ $d_{1}^{+}$ $=O(\epsilon^{2}),$ (B.50)
$d_{2}^{-}$ $=-\epsilon^{\ell+1}\frac{c_{1}^{(0)}\epsilon+O(\epsilon^{2})}{2},$ $d_{2}^{+}$ $=\frac{c_{2}^{(1)}\epsilon+O(\epsilon^{2})}{\epsilon^{\ell}},$
where the constants $c$’s are given in (B.13) and (B.29).
Let us now compare the two “far-zone” contributions in (B.44) and (B.46). In the MST result, we can take the integer-$L$ limit, which is smooth, and simply replace $L\rightarrow\ell\in\mathbb{N}$. Expanding the definitions, we find that
$\frac{\Gamma(L-i\epsilon+1)}{\Gamma(L+i\epsilon+1)}\operatorname{e}^{2i\epsilon\log(2\epsilon)}\frac{d_{2}^{-}-id_{1}^{-}}{d_{2}^{-}+id_{1}^{-}}-\frac{A_{-}^{\nu}}{A_{+}^{\nu}}\operatorname{e}^{i\epsilon\left(2\log\epsilon-(1-\kappa)\right)}=O(\epsilon^{2}).$ (B.51)
The “near-zone” term—corresponding to the square bracket in (B.44)—appears instead to be singular in the integer-$\ell$ limit. However, by expressing everything in terms of trigonometric functions, one can easily verify that the divergences cancel at first order in $\epsilon$, yielding
$1-i\frac{K_{-\nu-1}}{K_{\nu}}\left(i\operatorname{e}^{i\pi L}+\operatorname{e}^{-i\pi L}\frac{\sin(\pi(L+i\epsilon)}{\sin(\pi(L-i\epsilon)}\right)\underset{\underset{\epsilon\to 0}{L\to\ell\in\mathbb{N}}}{\sim}1-\epsilon^{2\ell+1}\left(\frac{2^{2\ell+1}(\ell!)^{6}}{((2\ell)!)^{2}((2\ell+1)!)^{2}}\epsilon+O(\epsilon^{2})\right),$ (B.52)
where the term in parenthesis matches the prefactor of the $\epsilon^{2\ell+1}$ term in the square brackets of (B.46), up to subleading orders in $\epsilon$, which we ignored here.
Appendix C Gravitational $\ell=3,4$ solutions
Here we collect the final expressions for the $\ell=3$ and $\ell=4$ solutions of the Regge–Wheeler and Zerilli fields up to second order in frequency. We first present the results from a full general relativistic calculation, in which the Regge–Wheeler and Zerilli equations are solved order by order in frequency. We then provide the corresponding expressions for the Born-series solutions obtained from the point-particle EFT for the black hole.
C.1 General relativistic solutions
We first write the solutions to the results of a general relativity computation of the tidal field.
Odd sector: Solving the Regge–Wheeler equation (3.2) perturbatively up to second-order in $\omega$, as described in Section 3.1, we find for $\ell=3$
$\Psi^{\ell=3}_{\text{RW,IZ}}(r)$ $\underset{\frac{r}{r_{s}}\to\infty}{=}B\bigg(6\frac{r^{4}}{r_{s}^{4}}-5\frac{r^{3}}{r_{s}^{3}}\bigg)+iB\omega r_{s}\bigg(\frac{137}{10}\frac{r^{4}}{r_{s}^{4}}-\frac{137}{12}\frac{r^{3}}{r_{s}^{3}}+\frac{1}{42}\frac{r_{s}^{3}}{r^{3}}\bigg)$ (C.1)
$\penalty 10000\ \hskip 28.45274pt+B\omega^{2}r_{s}^{2}\bigg[-\frac{1}{3}\frac{r^{6}}{r_{s}^{6}}-\frac{r^{5}}{r_{s}^{5}}-\frac{r^{4}}{r_{s}^{4}}\!\left(\pi^{2}+\frac{100\,243}{4\,200}-\frac{13}{7}\log\frac{r_{s}}{r}\right)$
$\penalty 10000\ \hskip 78.24507pt+\frac{5}{6}\frac{r^{3}}{r_{s}^{3}}\!\left(\pi^{2}+\frac{38\,981}{1\,400}-\frac{13}{7}\log\frac{r_{s}}{r}\right)-\frac{109}{210}\frac{r^{2}}{r_{s}^{2}}-\frac{13}{280}\frac{r}{r_{s}}$
$\penalty 10000\ \hskip 78.24507pt+\frac{19}{280}+\frac{337}{2\,800}\frac{r_{s}}{r}+\frac{1}{6}\frac{r_{s}^{2}}{r^{2}}+\frac{1}{42}\frac{r_{s}^{3}}{r^{3}}\!\left(\frac{223}{42}-\log\frac{r_{s}}{r}\right)\bigg]+O\!\left(\frac{r_{s}^{4}}{r^{4}}\right),$
and for $\ell=4$
[…]
$\penalty 10000\ \hskip 28.45274pt+B\omega^{2}r_{s}^{2}\bigg[-\frac{14}{11}\frac{r^{7}}{r_{s}^{7}}-\frac{147}{55}\frac{r^{6}}{r_{s}^{6}}-\frac{r^{5}}{r_{s}^{5}}\!\left(\frac{14\pi^{2}}{3}+\frac{3\,696\,439}{23\,100}-\frac{3\,142}{495}\log\frac{r_{s}}{r}\right)$
$\penalty 10000\ \hskip 78.24507pt+\frac{r^{4}}{r_{s}^{4}}\!\left(7\pi^{2}+\frac{35\,319\,721}{138\,600}-\frac{1\,571}{165}\log\frac{r_{s}}{r}\right)$
$\penalty 10000\ \hskip 78.24507pt-\frac{r^{3}}{r_{s}^{3}}\!\left(\frac{5\pi^{2}}{2}+\frac{762\,181}{7\,920}-\frac{1\,571}{462}\log\frac{r_{s}}{r}\right)+\frac{3\,373}{4\,158}\frac{r^{2}}{r_{s}^{2}}+\frac{4\,759}{27\,720}\frac{r}{r_{s}}$
$\penalty 10000\ \hskip 78.24507pt+\frac{11\,689}{138\,600}+\frac{57\,889}{831\,600}\frac{r_{s}}{r}+\frac{1}{14}\frac{r_{s}^{2}}{r^{2}}+\frac{9}{112}\frac{r_{s}^{3}}{r^{3}}$
$\penalty 10000\ \hskip 78.24507pt+\frac{1}{252}\frac{r_{s}^{4}}{r^{4}}\!\left(\frac{9\,767}{504}-\log\frac{r_{s}}{r}\right)\bigg]+O\!\left(\frac{r_{s}^{5}}{r^{5}}\right).$
Even sector: Similarly we tabulate the solutions to the Zerilli equation in the even sector. The solution for $\ell=3$ is
[…]
$\penalty 10000\ \hskip 78.24507pt+\frac{13\,923}{500\,000}\frac{r_{s}^{2}}{r^{2}}+\frac{1\,622\,851}{105\,000\,000}\frac{r_{s}^{3}}{r^{3}}\bigg)$
$\penalty 10000\ \hskip 28.45274pt+C\omega^{2}r_{s}^{2}\bigg[-\frac{1}{3}\frac{r^{6}}{r_{s}^{6}}-\frac{11}{10}\frac{r^{5}}{r_{s}^{5}}-\frac{r^{4}}{r_{s}^{4}}\!\left(\pi^{2}+\frac{51\,917}{2\,100}-\frac{13}{7}\log\frac{r_{s}}{r}\right)$
$\penalty 10000\ \hskip 78.24507pt+\frac{r^{3}}{r_{s}^{3}}\!\left(\frac{19\pi^{2}}{30}+\frac{399\,451}{21\,000}-\frac{247}{210}\log\frac{r_{s}}{r}\right)$
$\penalty 10000\ \hskip 78.24507pt+\frac{r^{2}}{r_{s}^{2}}\!\left(\frac{31\pi^{2}}{100}+\frac{1\,634\,147}{210\,000}-\frac{403}{700}\log\frac{r_{s}}{r}\right)$
$\penalty 10000\ \hskip 78.24507pt-\frac{r}{10^{3}r_{s}}\!\left(93\pi^{2}+\frac{2\,039\,147}{700}-\frac{1\,209}{7}\log\frac{r_{s}}{r}\right)$
$\penalty 10000\ \hskip 78.24507pt-\frac{221\pi^{2}}{10\,000}-\frac{10\,422\,677}{21\,000\,000}+\frac{2\,873}{70\,000}\log\frac{r_{s}}{r}$
$\penalty 10000\ \hskip 78.24507pt+\frac{r_{s}}{10^{5}r}\!\left(663\pi^{2}+\frac{21\,972\,677}{700}-\frac{8\,619}{7}\log\!\left(\frac{r_{s}}{r}\right)\right)$
$\penalty 10000\ \hskip 78.24507pt+\frac{r_{s}^{2}}{10^{6}r^{2}}\!\left(-1\,989\pi^{2}+\frac{220\,495\,907}{2\,100}+\frac{25\,857}{7}\log\frac{r_{s}}{r}\right)$
$\penalty 10000\ \hskip 78.24507pt+\frac{r_{s}^{3}}{10^{7}r^{3}}\!\left(5\,967\pi^{2}+\frac{58\,238\,757\,859}{44\,100}-\frac{5\,232\,713}{21}\log\frac{r_{s}}{r}\right)\bigg]+O\!\left(\frac{r_{s}^{4}}{r^{4}}\right),$
and the solution for $\ell=4$ is
[…]
$\penalty 10000\ \hskip 44.10185pt-\frac{287}{3\,888}+\frac{287}{23\,328}\frac{r_{s}}{r}-\frac{287}{139\,968}\frac{r_{s}^{2}}{r^{2}}+\frac{287}{839\,808}\frac{r_{s}^{3}}{r^{3}}-\frac{287}{5\,038\,848}\frac{r_{s}^{4}}{r^{4}}\bigg)$
$\penalty 10000\ \hskip 28.45274pt+Ci\omega r_{s}\bigg[\frac{1\,246}{15}\frac{r^{5}}{r_{s}^{5}}-\frac{10\,591}{90}\frac{r^{4}}{r_{s}^{4}}+\frac{15\,931}{540}\frac{r^{3}}{r_{s}^{3}}+\frac{32\,129}{3\,240}\frac{r^{2}}{r_{s}^{2}}-\frac{32\,129}{19\,440}\frac{r}{r_{s}}$
[…]
$\penalty 10000\ \hskip 28.45274pt+C\omega^{2}r_{s}^{2}\bigg[-\frac{14}{11}\frac{r^{7}}{r_{s}^{7}}-\frac{931}{330}\frac{r^{6}}{r_{s}^{6}}-\frac{r^{5}}{r_{s}^{5}}\!\left(\frac{14\pi^{2}}{3}+\frac{2\,798\,434}{17\,325}-\frac{3\,142}{495}\log\frac{r_{s}}{r}\right)$
$\penalty 10000\ \hskip 78.24507pt+\frac{r^{4}}{r_{s}^{4}}\!\left(\frac{119\pi^{2}}{18}+\frac{202\,591\,567}{831\,600}-\frac{26\,707}{2\,970}\log\frac{r_{s}}{r}\right)$
$\penalty 10000\ \hskip 78.24507pt-\frac{r^{3}}{r_{s}^{3}}\!\left(\frac{179\pi^{2}}{108}+\frac{332\,182\,607}{4\,989\,600}-\frac{281\,209}{124\,740}\log\frac{r_{s}}{r})\right)$
$\penalty 10000\ \hskip 78.24507pt-\frac{r^{2}}{r_{s}^{2}}\!\left(\frac{361\pi^{2}}{648}+\frac{603\,088\,753}{29\,937\,600}-\frac{567\,131}{748\,440}\log\frac{r_{s}}{r}\right)$
[…]
where the subleading terms that we have not displayed are $O(r_{s}^{5}/r^{5})$.
C.2 EFT solutions
Here we present the particular solutions obtained from computing the Born series for $\ell=3,4$.
Odd sector: We begin with the odd sector. The $\ell=3$ solution is
$\Psi_{\text{RW}}^{\ell=3}(r)=$ $\,\,r^{4}B_{\text{reg}}\left(1+\frac{361\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{441}-\frac{13\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{42\varepsilon}-\frac{13\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{14}\log(\mu r)\right)$ (C.5)
$+\frac{B_{\text{reg}}}{r^{3}}\left(\frac{287\,552\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{9}\omega^{2}}{19\,845}+\frac{8\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{9}\omega^{2}}{63\varepsilon}+\frac{136\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{9}\omega^{2}}{63}\log(\mu r)\right)$
$+\frac{B_{\text{irr}}}{r^{3}}\left(\frac{1}{7}+\frac{1\,108\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{3\,087}+\frac{13\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{294\varepsilon}+\frac{65\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{294}\log(\mu r)\right),$
while the $\ell=4$ solution is
[…]
$+\frac{B_{\text{reg}}}{r^{4}}\left(\frac{2\,291\,068\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{1}1\omega^{2}}{363\,825}+\frac{32\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{1}1\omega^{2}}{2\,205\varepsilon}+\frac{32\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{7}\omega^{2}}{105}\log(\mu r)\right)$
[…]
To subtract the infinities, we introduce the renormalized coefficients $\bar{B}_{\text{reg}}$ and $\bar{B}_{\text{irr}}$ as
$B_{\text{reg}}=\bar{B}_{\text{reg}}(1+\omega^{2}\delta_{11})+\bar{B}_{\text{irr}}\omega^{2}\delta_{12},\qquad\quad B_{\text{irr}}=\bar{B}_{\text{irr}}(1+\omega^{2}\delta_{22})+\bar{B}_{\text{reg}}\omega^{2}\delta_{21},$ (C.7)
where the $\delta$ parameters are (arranged in a matrix as in (2.64))
$\delta^{\ell=3}=\begin{pmatrix}\frac{13\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}}{42\varepsilon}&0\\ -\frac{8\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{9}}{9\varepsilon}&-\frac{13\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}}{42\varepsilon}\end{pmatrix},\qquad\delta^{\ell=4}=\begin{pmatrix}\frac{1571\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}}{6\,930\varepsilon}&0\\ -\frac{32\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{1}1}{245\varepsilon}&-\frac{1\,571\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}}{6\,930\varepsilon}\end{pmatrix}$. (C.8)
Plugging the renormalized coefficients (C.7) back into the expression for $\Psi_{\text{RW}}$ before taking the $\varepsilon\to 0$ limit yields the following renormalized solutions. For $\ell=3$ we have
$\Psi_{\text{RW}}^{\text{R},\ell=3}$ $=\bar{B}_{\text{reg}}\left[r^{4}+r^{4}\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}\left(\frac{361}{441}-\frac{26}{21}\log(\mu r)\right)+\frac{\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{9}\omega^{2}}{r^{3}}\left(\frac{278\,512}{19\,845}+\frac{128}{63}\log(\mu r)\right)\right]$ (C.9)
$\penalty 10000\ \penalty 10000\ \penalty 10000\ +\bar{B}_{\text{irr}}\left[\frac{1}{7r^{3}}+\frac{\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{r^{3}}\left(\frac{1\,069}{3\,087}+\frac{26}{147}\log(\mu r)\right)\right],$
while the renormalized solution for $\ell=4$ is
[…]
$\penalty 10000\ \penalty 10000\ \penalty 10000\ +\bar{B}_{\text{irr}}\left[\frac{1}{9r^{4}}+\frac{\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{r^{4}}\left(\frac{20\,550\,337}{72\,037\,350}+\frac{3\,142}{31\,185}\log(\mu r)\right)\right].$ (C.10)
Comparing with the solutions $\Psi_{\text{RW,IZ}}^{\ell}$ from Section C.1, we can fix the renormalized constants $\bar{B}_{\text{reg}}$ and $\bar{B}_{\text{irr}}$ as
$\bar{B}_{\text{reg}}^{\ell=3}=$ $\,B\left[6+i\omega r_{s}\frac{137}{10}-\omega^{2}r_{s}^{2}\left(\frac{737\,801}{29\,400}+\pi^{2}-\frac{13}{7}\log(\mu r_{s})\right)\right],$ (C.11)
$\bar{B}_{\text{irr}}^{\ell=3}=$ $\,B\left[\frac{i\omega r_{s}^{8}}{6}-\omega^{2}r_{s}^{9}\left(\frac{4\,027}{15\,120}+\frac{1}{6}\log(\mu r_{s})\right)\right],$
$\bar{B}_{\text{reg}}^{\ell=4}=$ $\,B\left[28+i\omega r_{s}\frac{413}{5}-\omega^{2}r_{s}^{2}\left(\frac{1\,142\,563\,973}{6\,860\,700}+\frac{14\pi^{2}}{3}-\frac{3\,142}{495}\log(\mu r_{s})\right)\right],$
$\bar{B}_{\text{irr}}^{\ell=4}=$ $\,B\left[\frac{i\omega r_{s}^{10}}{28}-\omega^{2}r_{s}^{11}\left(\frac{580\,211}{7\,761\,600}+\frac{1}{28}\log(\mu r_{s})\right)\right].$
From the ratio of these coefficients, we then obtain the response functions (3.88).
Even sector: In the even sector, the Born series computations are similar, but the expressions are considerably more lengthy. To simplify things, we exploit the absence of decaying fall-off in the static limit of the full black hole solution. In the following expressions, we will therefore take $B_{\text{irr}}\equiv\omega\tilde{B}_{\text{irr}}$, where $\tilde{B}_{\text{irr}}$ starts at order $O(\omega^{0})$. Then the solutions are
$\Psi_{\text{Z}}^{\ell=3}(r)=$ $\,\,r^{4}B_{\text{reg}}\left(1+\frac{8\,563\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{11\,025}-\frac{13\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{42\varepsilon}-\frac{13\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}}{14}\log(\mu r)\right)$ (C.12)
[…]
and for $\ell=4$
[…]
$\hskip 42.67912pt+\frac{28\,164\,541\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{11}\omega^{2}}{90\,935\,460}\log(\mu r)\bigg)+\frac{\omega\tilde{B}_{\text{irr}}}{9r^{4}}.$
In this notation, the renormalized coefficients read
$B_{\text{reg}}=\bar{B}_{\text{reg}}(1+\omega^{2}\delta_{11})+\bar{B}_{\text{irr}}\omega^{2}\delta_{12},\qquad\quad\omega\tilde{B}_{\text{irr}}=\bar{B}_{\text{irr}}(1+\omega^{2}\delta_{22})+\bar{B}_{\text{reg}}\omega^{2}\delta_{21},$ (C.14)
with
$\delta^{\ell=3}=\begin{pmatrix}\frac{13\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}}{42\varepsilon}&0\\ -\frac{8\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{9}}{9\varepsilon}&0\end{pmatrix},\qquad\qquad\delta^{\ell=4}=\begin{pmatrix}\frac{1\,571\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}}{6\,930\varepsilon}&0\\ -\frac{32\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{11}}{245\varepsilon}&0\end{pmatrix}$. (C.15)
Then, the renormalized EFT solutions are
$\Psi_{\text{Z}}^{\text{R},\ell=3}=$ $\,\bar{B}_{\text{reg}}\bigg[r^{4}+r^{4}\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}\left(\frac{8\,563}{11\,025}-\frac{26}{21}\log(\mu r)\right)\bigg]$ (C.16)
[…]
$\Psi_{\text{Z}}^{\text{R},\ell=4}=$ $\,\bar{B}_{\text{reg}}\bigg[r^{5}+r^{5}\mkern 2.5mu\overline{\mkern-2.5muG\mkern-2.5mu}\mkern 2.5mu^{2}\omega^{2}\left(\frac{8\,999\,989}{9\,604\,980}-\frac{3\,142}{3\,465}\log(\mu r)\right)\bigg]$ (C.17)
[…]
Matching with $\Psi_{\text{Z,IZ}}^{\ell}$, we find
$\bar{B}_{\text{reg}}^{\ell=3}$ $=\,B\left[6+14i\omega r_{s}-\omega^{2}r_{s}^{2}\left(\pi^{2}+\frac{76\,109}{2\,940}-\frac{13}{7}\log\mu r_{s}\right)\right],$ (C.18)
$\bar{B}_{\text{irr}}^{\ell=3}$ $=\,B\left[\frac{i\omega r_{s}^{8}}{6}-\omega^{2}r_{s}^{9}\left(\frac{139}{560}+\frac{1}{6}\log(\mu r_{s})\right)\right],$
$\bar{B}_{\text{reg}}^{\ell=4}$ $=\,B\left[28+i\omega r_{s}\frac{1\,246}{15}-\omega^{2}r_{s}^{2}\left(\frac{14\pi^{2}}{3}+\frac{1\,153\,179\,809}{6\,860\,700}-\frac{3\,142}{495}\log\mu r_{s}\right)\right],$
$\bar{B}_{\text{irr}}^{\ell=4}$ $=\,B\left[\frac{i\omega r_{s}^{10}}{28}-\omega^{2}r_{s}^{11}\left(\frac{23\,263}{323\,400}+\frac{1}{28}\log(\mu r_{s})\right)\right],$
The ratios of these fall-offs lead to the response functions (3.119).
Appendix D Tidal field in de Donder gauge
In Section 3.2, we made use of specific properties of the tidal field in de Donder gauge to simplify some expressions. Here, we present the details of those properties.
In particular, we will be interested in the last term in (3.63). Since it is already of $O(\omega)$, it will be sufficient for our purposes to work out the tidal field solution $\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{ij}$ in the static limit. We begin with the expression of the tidal field in Regge–Wheeler (RW) gauge [115], which we will transform to the de Donder (dD) gauge. In Minkowski space, this transformation can be written as
$h_{\mu\nu}^{\text{dD}}=h_{\mu\nu}^{\text{RW}}+\partial_{\mu}\xi_{\nu}+\partial_{\nu}\xi_{\mu},$ (D.1)
where $\xi^{\mu}$ defines the gauge parameter relating $h_{\mu\nu}^{\text{dD}}$ and $h_{\mu\nu}^{\text{RW}}$ (we omit for simplicity the bar over $h_{\mu\nu}$ to denote the background tidal field from now on). By definition, $h_{\mu\nu}^{\text{dD}}$ satisfies
$\partial^{\mu}h_{\mu\nu}^{\text{dD}}=\frac{1}{2}\partial_{\nu}h^{\text{dD}},$ (D.2)
and the equation of motion
$\square h_{\mu\nu}^{\text{dD}}=0.$ (D.3)
Denoting by $h_{\mu\nu}^{\text{odd}}$ ($h_{\mu\nu}^{\text{even}}$) the odd (even) components of $h_{\mu\nu}$, we have $h^{\text{odd}}=0$, [^38] This holds in any gauge—recall that the trace $h\equiv h^{\mu}_{\mu}$ is a scalar and vanishes, by definition, identically in the odd sector. and, from the trace of (D.1), $\partial^{\mu}\xi^{\text{odd}}_{\mu}=0$, with $\xi^{\text{odd}}_{\mu}$ the odd part of the gauge transformation. Since $\xi^{\text{odd}}_{0}=0$ identically, we then have
$\partial^{i}\xi^{\text{odd}}_{i}=0.$ (D.4)
This implies that $\xi^{\text{odd}}_{i}\equiv\epsilon_{ijk}\partial^{j}V^{k}$, for some vector $V^{k}$. Let us now recall that, after solving the constraint equations in the Regge–Wheeler gauge, one finds $h^{{\text{odd}},\text{RW}}_{ij}=0$. [^39] The only non-vanishing components of $h^{{\text{odd}},\text{RW}}_{\mu\nu}$ are $h^{{\text{odd}},\text{RW}}_{0i}$ in the static limit. See e.g., [35]. Then the previous conditions imply
$\vec{\nabla}^{2}\xi^{\text{odd}}_{i}=0.$ (D.5)
Expanding $\xi^{\text{odd}}_{i}$ in powers of $\vec{x}$, from (D.4) and (D.5) it follows that
$\xi^{\text{odd}}_{i}=\frac{1}{\ell+1}c_{i(j_{1}\cdots j_{\ell+1})_{T}}x^{j_{1}}\cdots x^{j_{\ell+1}},$ (D.6)
where $c_{i(j_{1}\cdots j_{\ell+1})_{T}}$ is a generic constant tensor, symmetric and traceless in its last $\ell+1$ indices, and satisfying ${c^{i}}_{(ij_{1}\cdots j_{\ell})_{T}}=0$.
Since $h^{{\text{odd}},\text{RW}}_{ij}=0$ in Regge–Wheeler gauge in the static regime, then
$h_{ij}^{{\text{odd}},\text{dD}}$ $=\left(c_{i(jj_{1}\cdots j_{\ell})_{T}}+c_{j(ij_{1}\cdots j_{\ell})_{T}}\right)x^{j_{1}}\cdots x^{j_{\ell}}$ (D.7)
$\equiv\mkern 2.5mu\overline{\mkern-2.5muc\mkern-2.5mu}\mkern 2.5mu_{(ijj_{1}\cdots j_{\ell})_{T}}x^{j_{1}}\cdots x^{j_{\ell}},$
where $c_{(ijj_{1}\cdots j_{\ell})_{T}}$ is—by construction—symmetric and traceless in all its indices. Plugging into (3.63), we can clearly see that
$\partial^{[k}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{\text{dD}}^{j]i}=\ell\left[\mkern 2.5mu\overline{\mkern-2.5muc\mkern-2.5mu}\mkern 2.5mu^{(jikj_{2}\cdots j_{\ell})_{T}}-\mkern 2.5mu\overline{\mkern-2.5muc\mkern-2.5mu}\mkern 2.5mu^{(kijj_{2}\cdots j_{\ell})_{T}}\right]x_{j_{2}}\cdots x_{j_{\ell}}=0,$ (D.8)
which is the property we used in the main text.

### v2

@@
 Interestingly, explicit calculations show that the nonlinear Love numbers of black holes continue to vanish in four-spacetime dimensions [14, 32, 33, 34, 35, 27, 36, 37, 38], hinting at a fully nonlinear symmetry structure underlying the static sector of general relativity [27, 38, 29].
-In this paper we study subleading in frequency tidal effects, namely the dynamical response.
+In this paper we study subleading-in-frequency tidal effects, namely the dynamical response.
 Specifically, we aim to compute the induced time-dependent tidal deformation of Schwarzschild black holes by matching results obtained in general relativity to a point-particle EFT description of black holes [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 6, 7, 13].
@@
 As such, the small frequency approximation is necessarily more complicated than simply neglecting the $r^{3}\omega^{2}/(r-r_{s})$ term.
-A natural way to proceed is to adopt an asymptotic expansion approach, reducing the problem to a simplified set of equations by making approximations which are valid in different regions with some overlap.
+A natural way to proceed is to adopt an asymptotic expansion approach, reducing the problem to a simplified set of equations by making approximations which are valid in different regions.
 The complete solution can then be reconstructed by matching the various solutions at the boundaries of their respective regions of validity.
@@
 A central conceptual point of the analysis is that in order to match to point-particle EFT, it suffices to know the solution in the Intermediate Zone, as this has an overlapping regime of validity with the EFT description, if one matches the one-point function.
+[^4] Note that what we are calling the “Intermediate Zone” is called the “Near Zone” in much of the literature.
+Perhaps what we call the Near Zone would better be called the “Very Near Zone”.
+We apologize for any confusion.
 In [74] they match to the EFT in the Far Zone, by matching scattering amplitudes to scattering computed in GR using the MST formalism.
@@
 $\frac{{\rm d}r_{\star}}{{\rm d}r}\equiv\left(1-\frac{r_{s}}{r}\right)^{-1},\qquad r_{\star}=r+r_{s}\log\left(\frac{r}{r_{s}}-1\right).$ (2.6)
-Imposing ingpoing boundary conditions at the horizon, the physical near-zone solution thus reads
+Imposing ingoing boundary conditions at the horizon, the physical near-zone solution thus reads
 $\phi_{\text{NZ}}(r)=B\frac{r_{s}}{r}\mathrm{e}^{-i\omega\left(r+r_{s}\log\left(\frac{r}{r_{s}}-1\right)\right)},$ (2.7)
@@
 These can be determined by matching the perturbative solutions $\phi_{\ell}^{(0)}(x)$ and $\phi_{\ell}^{(1)}(x)$ to the near-zone solution (2.7) across the boundary at $r=r_{s}$.
+[^7] As noted in [88], formally the approximations that we have made to define the Near and Intermediate zones have no region of overlap.
+Nevertheless, this does not preclude us from matching across their interface [87].
 To this end, it is useful to recall the asymptotic behavior of the Legendre polynomials near $x=1$ (see e.g., [93, 94]):
@@
 $\partial_{x}W=0$.
-In particular, it is straightforward to check that $W\equiv 1$.
+In particular, one can check that $W\equiv 1$.
 $\phi_{\ell,\text{p}}^{(2)}(x)=Q_{\ell}(x)\int^{x}_{x_{0}}{\rm d}y\,P_{\ell}(y)S_{\ell}(y)-P_{\ell}(x)\int^{x}_{x_{0}}{\rm d}y\,Q_{\ell}(y)S_{\ell}(y),$ (2.20)
@@
 This can be rectified by defining the projected covariant derivative $\nabla^{\bot}_{\mu}\equiv{\mathbb{P}^{\nu}}_{\mu}\nabla_{\nu}$, where ${\mathbb{P}^{\nu}}_{\mu}\equiv\delta_{\mu}^{\nu}+u_{\mu}u^{\nu}$ projects onto the plane orthogonal to the four velocity $u^{\mu}\equiv{\rm d}x^{\mu}/{\rm d}\tau$ (where $\tau$ is the proper time, and normalized such that $u^{\mu}u_{\mu}=-1$).
-The composite operator multipole operator $Q_{I}^{A_{\ell}}(\tau)$ is built from internal degrees of freedom $X_{\pm}$ localized on the worldline of the particle, which physically we can think of as modeling the degrees of freedom into which energy and other charges are dissipating.
+The composite multipole operator $Q_{I}^{A_{\ell}}(\tau)$ is built from internal degrees of freedom $X_{\pm}$ localized on the worldline of the particle, which physically we can think of as modeling the degrees of freedom into which energy and other charges are dissipating.
 [^12] Introducing the auxiliary worldline degrees of freedom $X$ is not strictly necessary.
@@
 $\Gamma^{\text{in-in}}_{\text{int}}=\int{\rm d}\tau_{1}{\rm d}\tau_{2}\sum_{\ell=0}^{\infty}K_{{IJ}}^{(\Phi)\,\ell}(\tau_{2}-\tau_{1})\nabla^{(\ell)}_{A_{\ell}}\Phi^{I}(\tau_{2})\nabla^{(\ell)}{}^{A_{\ell}}{\Phi}^{J}(\tau_{1})\,$, (2.35)
-where we have taken advantage of the fact that in the case of interest the system is rotationally symmetric.
+where we have used the fact that in the case of interest the system is rotationally symmetric.
 [^13] This means that the Green’s function (2.34) is diagonal in $\ell,\ell^{\prime}$ and its index structure can be written in terms of $\delta_{ij}$.
@@
 $\langle\Phi_{+}(t,\vec{x})\rangle_{\text{in-in}}=\operatorname{e}^{-i\omega t}\sum_{\ell=0}^{\infty}(-1)^{\ell}K^{(\Phi)}_{\ell}(\omega)\frac{2^{\ell-2}\ell!\,\Gamma\left(\frac{D-3}{2}\right)\Gamma\left(\frac{5-D}{2}\right)}{\pi^{\frac{D-1}{2}}\Gamma\left(\frac{5-D}{2}-\ell\right)}c_{i_{1}\cdots i_{\ell}}\frac{x^{i_{1}}\cdots x^{i_{\ell}}}{|\vec{x}|^{2\ell+D-3}}\,.$ (2.43)
-Simplifying the ratio of gamma functions and writing the tidal field in spherical coordinates as $c_{i_{1}\cdots i_{\ell}}x^{i_{1}}\cdots x^{i_{\ell}}=c_{\text{ext}}r^{\ell}Y_{\ell}^{m}$, we obtain
+Simplifying the ratio of gamma functions and writing the tidal field in spherical coordinates as $c_{i_{1}\cdots i_{\ell}}x^{i_{1}}\cdots x^{i_{\ell}}=c_{\text{ext}}r^{\ell}Y_{\ell}^{m}$, we obtain [^15] We suppress indices on $c_{\text{ext}}$ which depends implicitly on the spherical harmonic quantum numbers $\ell,m$.
 $\langle\Phi_{+}\rangle_{\text{in-in}}=\operatorname{e}^{-i\omega t}\sum_{\ell=0}^{\infty}K^{(\Phi)}_{\ell}(\omega)\frac{2^{\ell-2}\ell!\,\Gamma\left(\ell+\frac{D-3}{2}\right)}{\pi^{\frac{D-1}{2}}}\frac{c_{\text{ext}}Y_{\ell}^{m}}{r^{\ell+D-3}}\,.$ (2.44)
 Combining this with the external field and replacing $K_{\ell}^{(\Phi)}$ by $\mu^{2\varepsilon}K_{\ell}^{(\Phi)}$ (where $\mu$ is a dimensionful constant in order to keep the response function dimension independent) yields
-$\mkern 2.5mu\overline{\mkern-2.5mu\Phi\mkern-2.5mu}\mkern 2.5mu_{+}+\langle\Phi_{+}\rangle_{\text{in-in}}=c_{\text{ext}}\operatorname{e}^{-i\omega t}\sum_{\ell=0}^{\infty}Y_{\ell}^{m}\left(r^{\ell}+\mu^{2\varepsilon}K^{(\Phi)}_{\ell}(\omega)\frac{2^{\ell-2}\ell!\Gamma\left(\ell+\frac{D-3}{2}\right)}{\pi^{\frac{D-1}{2}}}r^{-\ell-D+3}\right).$ (2.45)
+$\mkern 2.5mu\overline{\mkern-2.5mu\Phi\mkern-2.5mu}\mkern 2.5mu_{+}+\langle\Phi_{+}\rangle_{\text{in-in}}=\operatorname{e}^{-i\omega t}\sum_{\ell=0}^{\infty}c_{\text{ext}}Y_{\ell}^{m}\left(r^{\ell}+\mu^{2\varepsilon}K^{(\Phi)}_{\ell}(\omega)\frac{2^{\ell-2}\ell!\Gamma\left(\ell+\frac{D-3}{2}\right)}{\pi^{\frac{D-1}{2}}}r^{-\ell-D+3}\right).$ (2.45)
 This result gives the scalar field response, generated by the interaction term of the action $S_{\text{int}}$, in the vicinity of the point particle.
@@
 with $B_{\text{reg}}$ and $B_{\text{irr}}$ the two arbitrary integration constants.
-[^14] For convenience, we extracted the constant prefactor $\frac{\mu^{2\epsilon}}{2\ell+1-2\epsilon}$ from the decaying branch.
+[^17] For convenience, we extracted the constant prefactor $\frac{\mu^{2\varepsilon}}{2\ell+1-2\varepsilon}$ from the decaying branch.
 Implicitly, the $r$-independent coefficients $B_{\text{reg}}$ and $B_{\text{irr}}$ should be thought of as expanded in powers of $G$ and $\omega$, as we discussed, and will see explicitly later on.
@@
 In practice, we determine the Love numbers $K_{\ell}^{(\Phi)}(\omega)$ by matching the homogeneous solution for $\Phi$
-$\Phi_{+}^{(h)}=\operatorname{e}^{-i\omega t}\sum_{\ell=0}^{\infty}Y_{\ell}^{m}\,B_{\rm reg}\left(r^{\ell}+\frac{B_{\rm irr}}{B_{\rm reg}}\frac{\mu^{2\epsilon}}{2\ell+1-2\varepsilon}r^{-\ell-1+2\varepsilon}\right),$ (2.52)
+$\Phi_{+}^{(h)}=\operatorname{e}^{-i\omega t}\sum_{\ell=0}^{\infty}Y_{\ell}^{m}\,B_{\rm reg}\left(r^{\ell}+\frac{B_{\rm irr}}{B_{\rm reg}}\frac{\mu^{2\varepsilon}}{2\ell+1-2\varepsilon}r^{-\ell-1+2\varepsilon}\right),$ (2.52)
 to the one-point function (2.45). This yields
@@
 We first construct perturbative-in-frequency solutions to the Regge–Wheeler and Zerilli equations, paralleling the scalar analysis of Section 2 (see also [60, 67] for related treatments).
-We then match these solutions to the worldline EFT in dimensional regularization, and explicitly compute the dynamical Love number coefficients, including, for the first time, the finite scheme-dependent terms.
+We then match these solutions to the worldline EFT in dimensional regularization, which allows us to explicitly compute the dynamical Love number coefficients, including, for the first time, the finite scheme-dependent terms.
 3.1 General relativistic solution via small-frequency expansion
@@
 We are interested in the dynamics of perturbations around a background Schwarzschild geometry (2.26).
-Denoting it with $\mkern 2.5mu\overline{\mkern-2.5mug\mkern-2.5mu}\mkern 2.5mu_{\mu\nu}$, we perturb the metric as ${g}_{\mu\nu}=\mkern 2.5mu\overline{\mkern-2.5mug\mkern-2.5mu}\mkern 2.5mu_{\mu\nu}+h_{\mu\nu}$.
+Denoting it with $\mkern 2.5mu\overline{\mkern-2.5mug\mkern-2.5mu}\mkern 2.5mu_{\mu\nu}$, we perturb the metric as ${g}_{\mu\nu}=\mkern 2.5mu\overline{\mkern-2.5mug\mkern-2.5mu}\mkern 2.5mu_{\mu\nu}+2h_{\mu\nu}/M_{\rm Pl}$.
 To make maximal use of the ${\rm SO}(D-1)$ background symmetry, it is convenient to decompose $h_{\mu\nu}$ into scalar, vector, and tensor spherical harmonics.
@@
 $K_{\ell}^{(B)}(\tau_{2}-\tau_{1})$ $=\int\frac{{\rm d}{\omega}}{2\pi}\operatorname{e}^{-i{\omega}(\tau_{2}-\tau_{1})}K_{\ell}^{(B)}(\omega)$ (3.53)
-$=\frac{1}{\ell!}\int\frac{{\rm d}{\omega}}{2\pi}\operatorname{e}^{-i{\omega}(\tau_{2}-\tau_{1})}\left[\frac{\lambda_{0,\ell}^{B}}{2}+i{\omega}r_{s}\frac{\lambda_{1,\ell}^{B}}{2}+({\omega}r_{s})^{2}\frac{\lambda_{2,\ell}^{B}}{2}+\cdots\right]\,$,
-where the parameters $\lambda_{i}$ are the response coefficients.
+$=\frac{1}{\ell!}\frac{\ell}{\ell+1}\int\frac{{\rm d}{\omega}}{2\pi}\operatorname{e}^{-i{\omega}(\tau_{2}-\tau_{1})}\left[\lambda_{0,\ell}^{B}+i{\omega}r_{s}\lambda_{1,\ell}^{B}+({\omega}r_{s})^{2}\lambda_{2,\ell}^{B}+\cdots\right]\,$,
+where the parameters $\lambda_{i}$ are the response coefficients and the overall normalization is chosen for later convenience.
 Terms with odd powers of $\omega$ (which are time-reversal odd) capture the dissipative response of the object, while those with even powers of $\omega$ parametrize the conservative tidal deformability.
@@
 A convenient choice is the de Donder gauge, defined by $\partial^{\mu}(h_{\mu\nu}-\frac{1}{2}\eta_{\mu\nu}h)=0$, with $h\equiv\eta^{\mu\nu}h_{\mu\nu}$.
-In this gauge, the linearized equations of motion are $\square h_{\mu\nu}=0$, and the on-shell magnetic component of the Weyl tensor reads
-$B_{abc}=C_{0abc}=2\left(\partial_{a}\partial_{[b}h_{c]0}-\partial_{0}\partial_{[b}h_{c]a}\right).$ (3.54)
+In this gauge, the linearized equations of motion are $\square h_{\mu\nu}=0$, and the on-shell magnetic component of the Weyl tensor reads [^28] Recall that $h_{\mu\nu}$ is the canonically normalized graviton field.
+$B_{abc}=C_{0abc}=\frac{2}{M_{\rm Pl}}\left(\partial_{a}\partial_{[b}h_{c]0}-\partial_{0}\partial_{[b}h_{c]a}\right).$ (3.54)
 Similarly, we can write the symmetrized derivative of the Weyl tensor as
-$B_{A_{\ell}\,j}=\partial_{i_{L}}h_{0j}-\partial_{j}\partial_{(i_{1}}\cdots\partial_{i_{\ell-1}}h_{i_{\ell})_{T}0}+\partial_{0}F_{A_{\ell}\,j}\,$, (3.55)
+$B_{A_{\ell}\,j}=\frac{1}{M_{\rm Pl}}\left[\partial_{i_{L}}h_{0j}-\partial_{j}\partial_{(i_{1}}\cdots\partial_{i_{\ell-1}}h_{i_{\ell})_{T}0}+\partial_{0}F_{A_{\ell}\,j}\right]\,$, (3.55)
 where we have defined the combination
@@
 Therefore, we are left with
-${B}_{-\,A_{\ell}j}\mkern 2.5mu\overline{\mkern-2.5mu{B}\mkern-2.5mu}\mkern 2.5mu_{+}^{A_{\ell}\,j}=$   $2\frac{\ell+1}{\ell}\partial_{A_{\ell}}h_{0j}\partial^{A_{\ell-1}}\partial^{[i_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{j]}{}_{0}+\partial_{A_{\ell}}h_{0j}\partial_{0}\mkern 2.5mu\overline{\mkern-2.5muF\mkern-2.5mu}\mkern 2.5mu^{A_{\ell}\,j}-(\partial_{j}\partial_{(A_{\ell-1}}h_{i_{\ell})_{T}0})\partial_{0}\mkern 2.5mu\overline{\mkern-2.5muF\mkern-2.5mu}\mkern 2.5mu^{A_{\ell}\,j}$ (3.58)
-$+\partial_{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{0j}\partial_{0}F^{A_{\ell}\,j}-(\partial_{j}\partial_{(A_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{j_{\ell})_{T}0})\partial_{0}F^{A_{\ell}\,j}+O(\omega^{2}).$
+${B}_{-\,A_{\ell}j}\mkern 2.5mu\overline{\mkern-2.5mu{B}\mkern-2.5mu}\mkern 2.5mu_{+}^{A_{\ell}\,j}=\frac{1}{M_{\rm Pl}^{2}}\bigg[$ $2\frac{\ell+1}{\ell}\partial_{A_{\ell}}h_{0j}\partial^{A_{\ell-1}}\partial^{[i_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{j]}{}_{0}+\partial_{A_{\ell}}h_{0j}\partial_{0}\mkern 2.5mu\overline{\mkern-2.5muF\mkern-2.5mu}\mkern 2.5mu^{A_{\ell}\,j}-(\partial_{j}\partial_{(A_{\ell-1}}h_{i_{\ell})_{T}0})\partial_{0}\mkern 2.5mu\overline{\mkern-2.5muF\mkern-2.5mu}\mkern 2.5mu^{A_{\ell}\,j}$ (3.58)
+$+\partial_{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{0j}\partial_{0}F^{A_{\ell}\,j}-(\partial_{j}\partial_{(A_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{j_{\ell})_{T}0})\partial_{0}F^{A_{\ell}\,j}+O(\omega^{2})\bigg].$
 Let us focus for a moment on the terms involving $F$ in (3.58).
@@
 As a result, (3.58) is
-${B}_{-\,A_{\ell}\,j}\mkern 2.5mu\overline{\mkern-2.5mu{B}\mkern-2.5mu}\mkern 2.5mu_{+}^{A_{\ell}\,j}=$   $2\frac{\ell+1}{\ell}\partial_{A_{\ell}}h_{0j}\partial^{A_{\ell-1}}\partial^{[i_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{j]}{}_{0}$ (3.59)
+${B}_{-\,A_{\ell}\,j}\mkern 2.5mu\overline{\mkern-2.5mu{B}\mkern-2.5mu}\mkern 2.5mu_{+}^{A_{\ell}\,j}=\frac{1}{M_{\rm Pl}^{2}}\bigg[$ $2\frac{\ell+1}{\ell}\partial_{A_{\ell}}h_{0j}\partial^{A_{\ell-1}}\partial^{[i_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{j]}{}_{0}$ (3.59)
 $-2\frac{\ell+1}{\ell}\left[\partial_{A_{\ell}}h_{0j}\left(\partial_{0}\partial^{A_{\ell-2}}\partial^{[i_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{j]i_{\ell}}\right)+(h\leftrightarrow\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu)\right]+O(\omega^{2}).$
 Combining this expression with (3.54), we thus have
-$\langle B_{+\,abc}{B}_{-\,A_{\ell}\,j}\rangle\mkern 2.5mu\overline{\mkern-2.5mu{B}\mkern-2.5mu}\mkern 2.5mu_{+}^{A_{\ell}\,j}=$ $\,\,\frac{4(\ell+1)}{\ell}\langle(\partial_{a}\partial_{[b}h_{+\,c]0})(\partial_{A_{\ell}}h_{-\,j0})\rangle\left[\partial^{A_{\ell-1}}\partial^{[i_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j]}{}_{0}-\partial_{0}\partial^{A_{\ell-2}}\partial^{[i_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j]i_{\ell}}\right]$ (3.60)
-$-\frac{4(\ell+1)}{\ell}\langle(\partial_{a}\partial_{[b}h_{+\,c]0})(\partial_{0}\partial_{A_{\ell-2}}\partial_{[i_{\ell-1}}h_{-\,j]i_{\ell}})\rangle\partial^{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j}{}_{0}$
-$-\frac{4(\ell+1)}{\ell}\langle(\partial_{0}\partial_{[b}h_{+\,c]a})(\partial_{A_{\ell}}h_{-\,0j})\rangle\partial^{A_{\ell-1}}\partial^{[i_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j]}{}_{0}+O(\omega^{2})$
+$\langle B_{+\,abc}{B}_{-\,A_{\ell}\,j}\rangle\mkern 2.5mu\overline{\mkern-2.5mu{B}\mkern-2.5mu}\mkern 2.5mu_{+}^{A_{\ell}\,j}=$ $\,\,\frac{4(\ell+1)}{\ell M_{\rm Pl}^{3}}\langle(\partial_{a}\partial_{[b}h_{+\,c]0})(\partial_{A_{\ell}}h_{-\,j0})\rangle\left[\partial^{A_{\ell-1}}\partial^{[i_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j]}{}_{0}-\partial_{0}\partial^{A_{\ell-2}}\partial^{[i_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j]i_{\ell}}\right]$ (3.60)
+$-\frac{4(\ell+1)}{\ell M_{\rm Pl}^{3}}\langle(\partial_{a}\partial_{[b}h_{+\,c]0})(\partial_{0}\partial_{A_{\ell-2}}\partial_{[i_{\ell-1}}h_{-\,j]i_{\ell}})\rangle\partial^{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j}{}_{0}$
+$-\frac{4(\ell+1)}{\ell M_{\rm Pl}^{3}}\langle(\partial_{0}\partial_{[b}h_{+\,c]a})(\partial_{A_{\ell}}h_{-\,0j})\rangle\partial^{A_{\ell-1}}\partial^{[i_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j]}{}_{0}+O(\omega^{2})$
 To compute this up to linear order in $\omega$, we can use the instantaneous propagator for the graviton in de Donder gauge [7]:
@@
 We are thus left with
-$\langle B_{+\,abc}{B}_{-\,A_{\ell}\,j}\rangle\mkern 2.5mu\overline{\mkern-2.5mu{B}\mkern-2.5mu}\mkern 2.5mu_{+}^{A_{\ell}\,j}=\frac{4(\ell+1)}{\ell}\langle(\partial_{a}\partial_{[b}h_{+\,c]0})(\partial_{A_{\ell}}h_{-\,j0})\rangle\left[\partial^{A_{\ell-1}}\partial^{[i_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j]}{}_{0}-\partial_{0}\partial^{A_{\ell-2}}\partial^{[i_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j]i_{\ell}}\right]+O(\omega^{2}).$ (3.63)
+$\langle B_{+\,abc}{B}_{-\,A_{\ell}\,j}\rangle\mkern 2.5mu\overline{\mkern-2.5mu{B}\mkern-2.5mu}\mkern 2.5mu_{+}^{A_{\ell}\,j}=\frac{4(\ell+1)}{\ell M_{\rm Pl}^{3}}\langle(\partial_{a}\partial_{[b}h_{+\,c]0})(\partial_{A_{\ell}}h_{-\,j0})\rangle\left[\partial^{A_{\ell-1}}\partial^{[i_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j]}{}_{0}-\partial_{0}\partial^{A_{\ell-2}}\partial^{[i_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j]i_{\ell}}\right]+O(\omega^{2}).$ (3.63)
 It is convenient at this point to use the explicit form of the external tidal field $\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}$.
@@
 In cartesian coordinates, the component $\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{0j}$ ca be written as
-$\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{0j}(t,\vec{x})=\operatorname{e}^{-i\omega t}\sum_{\ell}c_{j|j_{1}\cdots j_{\ell}}x^{j_{1}}\cdots x^{j_{\ell}}+O(\omega^{2}),$ (3.64)
+$\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{0j}(t,\vec{x})=M_{\rm Pl}\operatorname{e}^{-i\omega t}\sum_{\ell}c_{j|j_{1}\cdots j_{\ell}}x^{j_{1}}\cdots x^{j_{\ell}}+O(\omega^{2}),$ (3.64)
 with $c$ symmetric and traceless in its $j_{1}\cdots j_{\ell}$ indices (and vanishing when totally symmetrized).
@@
 As a result, (3.60) reduces to its first term, and the one-point function for the magnetic component of the Weyl tensor reads
-$\langle B_{+\,abc}(t,\vec{x})\rangle_{\text{in-in}}=4i\sum_{\ell=2}^{\infty}\frac{(-1)^{\ell}(\ell+1)}{\ell}\int$ ${\rm d}\tau_{1}{\rm d}\tau_{2}\int\frac{{\rm d}{\omega}}{2\pi}\operatorname{e}^{-i{\omega}(\tau_{2}-\tau_{1})}K^{(B)}_{\ell}({\omega})$ (3.65)
 […]
@@
 Plugging in the expressions for the propagator (3.61) and the external field (3.64), we obtain
-$\langle B_{+\,abc}(t,\vec{x})\rangle_{\text{in-in}}=-\operatorname{e}^{-i\omega t}\sum_{\ell=2}^{\infty}\frac{4(-i)^{\ell}(\ell+1)!}{\ell}K^{(B)}_{\ell}(\omega)c^{[j|i_{1}]\cdots i_{\ell}}\mathcal{P}_{0j0[c}^{\text{dD}}\partial_{b]}\partial_{a}\int\frac{{\rm d}^{D-1}\vec{p}}{(2\pi)^{D-1}}\operatorname{e}^{i\vec{p}\cdot\vec{x}}\frac{p_{i_{1}}\cdots p_{i_{\ell}}}{\vec{p}^{2}}\,.$ (3.66)
+$\langle B_{+\,abc}(t,\vec{x})\rangle_{\text{in-in}}=-\frac{\operatorname{e}^{-i\omega t}}{M_{\rm Pl}^{2}}\sum_{\ell=2}^{\infty}\frac{4(-i)^{\ell}(\ell+1)!}{\ell}K^{(B)}_{\ell}(\omega)c^{[j|i_{1}]\cdots i_{\ell}}\partial_{a}\mathcal{P}_{0j0[c}^{\text{dD}}\partial_{b]}\int\frac{{\rm d}^{D-1}\vec{p}}{(2\pi)^{D-1}}\operatorname{e}^{i\vec{p}\cdot\vec{x}}\frac{p_{i_{1}}\cdots p_{i_{\ell}}}{\vec{p}^{2}}\,.$ (3.66)
 We can Fourier transform this back into position space
-$\langle B_{+abc}(t,\vec{x})\rangle_{\text{in-in}}\!=\!\operatorname{e}^{-i\omega t}\sum_{\ell=2}^{\infty}\frac{(-1)^{\ell+1}(\ell+1)!}{\ell}K^{(B)}_{\ell}(\omega)\frac{2^{\ell-1}\Gamma\left(\frac{D-3}{2}\right)\Gamma\left(\frac{5-D}{2}\right)}{\pi^{\frac{D-1}{2}}\Gamma\left(\frac{5-D}{2}-\ell\right)}c_{[j|i_{1}]\cdots i_{\ell}}\delta^{j}_{[c}\partial_{b]}\partial_{a}\frac{x^{i_{1}}\cdots x^{i_{\ell}}}{|\vec{x}|^{2\ell+D-3}}\,$, (3.67)
-and simplify the tensor structure using
-$c_{[j|i_{1}]\cdots i_{\ell}}\delta^{j}_{[c}\partial_{b]}\partial_{a}\left(x^{i_{1}}\cdots x^{i_{\ell}}\right)=\partial_{a}\partial_{[b}\left(x^{i_{1}}\cdots x^{i_{\ell}}\right)c_{c]|A_{\ell}}.$ (3.68)
-Considering the $B_{rij}$ component of (3.66), and writing the tidal field in spherical coordinates as
+$\langle B_{+abc}(t,\vec{x})\rangle_{\text{in-in}}\!=\frac{\operatorname{e}^{-i\omega t}}{M_{\rm Pl}^{2}}\sum_{\ell=2}^{\infty}\frac{(-1)^{\ell+1}(\ell+1)!}{\ell}K^{(B)}_{\ell}(\omega)\frac{2^{\ell-1}\Gamma\left(\frac{D-3}{2}\right)\Gamma\left(\frac{5-D}{2}\right)}{\pi^{\frac{D-1}{2}}\Gamma\left(\frac{5-D}{2}-\ell\right)}c_{[j|i_{1}]\cdots i_{\ell}}\partial_{a}\delta^{j}_{[c}\partial_{b]}\frac{x^{i_{1}}\cdots x^{i_{\ell}}}{|\vec{x}|^{2\ell+D-3}}$ (3.67)
+and simplify the tensor structure using [^29] Note that in deriving (3.2.1) we assume that $\partial_{b}$ is a derivative with respect to coordinates on the (D2)(D-2)-sphere—which we make explicit in (3.70).
+This allows us to factor out $1/|\vec{x}|^{2\ell+D-3}$ in (3.2.1).
+$c_{[j|i_{1}]\cdots i_{\ell}}$ $\delta^{j}_{[c}\partial_{b]}\left(x^{i_{1}}\cdots x^{i_{\ell}}\right)=\frac{1}{4}\left(\partial_{b}c_{c|A_{\ell}}-\partial_{c}c_{b|A_{\ell}}-\partial_{b}c_{i_{\ell}|cA_{\ell-1}}+\partial_{c}c_{i_{\ell}|bA_{\ell-1}}\right)x^{i_{1}}\cdots x^{i_{\ell}}$
+$=\frac{1}{4}\left[\partial_{b}(x^{i_{1}}\cdots x^{i_{\ell}})c_{c|A_{\ell}}+\partial_{c}\left(x^{i}c_{i|A_{\ell}}\frac{1}{\ell}\partial_{b}(x^{i_{1}}\cdots x^{i_{\ell}})\right)-(c\leftrightarrow b)\right]$
+$=\frac{1}{4}\left[\partial_{b}(x^{i_{1}}\cdots x^{i_{\ell}})c_{c|A_{\ell}}+\frac{1}{\ell}\partial_{c}\partial_{b}\left(x^{i}c_{i|A_{\ell}}x^{i_{1}}\cdots x^{i_{\ell}}\right)-\frac{1}{\ell}\partial_{c}\left(c_{b|A_{\ell}}x^{i_{1}}\cdots x^{i_{\ell}}\right)-(c\leftrightarrow b)\right]$
+$=\frac{1}{4}\left[\partial_{b}(x^{i_{1}}\cdots x^{i_{\ell}})c_{c|A_{\ell}}+\frac{1}{\ell}\partial_{b}\left(x^{i_{1}}\cdots x^{i_{\ell}}\right)c_{c|A_{\ell}}-(c\leftrightarrow b)\right]$
+$=\frac{\ell+1}{2\ell}\partial_{[b}\left(x^{i_{1}}\cdots x^{i_{\ell}}\right)c_{c]|A_{\ell}}.$ (3.68)
+Considering the $B_{rAB}$ component of (3.66), where $AB$ denote angular coordinates, and writing the tidal field in spherical coordinates as
 $c_{c|i_{1}\cdots i_{\ell}}x^{i_{1}}\cdots x^{i_{\ell}}=c_{\text{ext}}r^{\ell+1}Y_{c}^{(T)}{}_{\ell}^{m}$, (3.69)
 with $Y_{c}^{(T)}{}_{\ell}^{m}$ a vector spherical harmonic [7], we finally obtain
-$\langle B_{+\,rij}(x)\rangle_{\text{in-in}}=-\operatorname{e}^{-i\omega t}\sum_{\ell=2}^{\infty}\frac{(\ell+1)!}{\ell}K^{(B)}_{\ell}(\omega)\frac{2^{\ell-1}\Gamma\left(\ell+\frac{D-3}{2}\right)}{\pi^{\frac{D-1}{2}}}\nabla_{[i}Y_{j]}^{(T)}{}_{\ell}^{m}r^{2}\partial_{r}\left(r^{-2}\frac{c_{\text{ext}}}{r^{\ell+D-4}}\right).$ (3.70)
+$\langle B_{+\,rAB}(x)\rangle_{\text{in-in}}=-\frac{\operatorname{e}^{-i\omega t}}{M_{\rm Pl}^{2}}\sum_{\ell=2}^{\infty}\frac{\ell+1}{2\ell}\frac{(\ell+1)!}{\ell}K^{(B)}_{\ell}(\omega)\frac{2^{\ell-1}\Gamma\left(\ell+\frac{D-3}{2}\right)}{\pi^{\frac{D-1}{2}}}\nabla_{[A}Y_{B]}^{(T)}{}_{\ell}^{m}r^{2}\partial_{r}\left(r^{-2}\frac{c_{\text{ext}}}{r^{\ell+D-4}}\right).$ (3.70)
 We can also write the background tidal field in spherical coordinates
-$\mkern 2.5mu\overline{\mkern-2.5muB\mkern-2.5mu}\mkern 2.5mu_{rij}$ $=2\nabla_{r}\nabla_{[i}h_{j]0}-2\partial_{t}\nabla_{[i}h_{j]r}$ (3.71)
-$=2\operatorname{e}^{-i\omega t}\sum_{\ell=2}^{\infty}c_{\text{ext}}r^{2}\partial_{r}\big(r^{\ell-1}\big)\big(1+O(\omega r)\big)\nabla_{[i}Y_{j]}^{(T)}{}_{\ell}^{m}$.
-The $O(\omega r)$ term in $\mkern 2.5mu\overline{\mkern-2.5muB\mkern-2.5mu}\mkern 2.5mu_{rij}$ represents a far-zone, frequency-dependent correction to the static tidal field.
+$\mkern 2.5mu\overline{\mkern-2.5muB\mkern-2.5mu}\mkern 2.5mu_{rAB}$ $=\frac{2}{M_{\rm Pl}}\left[\nabla_{r}\nabla_{[i}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{j]0}-\partial_{t}\nabla_{[i}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{j]r}\right]$ (3.71)
+$=2\operatorname{e}^{-i\omega t}\sum_{\ell=2}^{\infty}c_{\text{ext}}r^{2}\partial_{r}\big(r^{\ell-1}\big)\big(1+O(\omega r)\big)\nabla_{[A}Y_{B]}^{(T)}{}_{\ell}^{m}$.
+The $O(\omega r)$ term in $\mkern 2.5mu\overline{\mkern-2.5muB\mkern-2.5mu}\mkern 2.5mu_{rAB}$ represents a far-zone, frequency-dependent correction to the static tidal field.
 It exhibits a different scaling in $r$ compared to the standard $r^{\ell}$ and $r^{-\ell-1}$ falloffs.
@@
 […]
-$\times\left(r^{\ell}+\mu^{2\varepsilon}K^{(B)}_{\ell}(\omega)\frac{(\ell+1)!(D+\ell-2)}{\ell(\ell-1)}\frac{2^{\ell-2}\Gamma\left(\ell+\frac{D-3}{2}\right)}{\pi^{\frac{D-1}{2}}}r^{-\ell+3-D}\right),$
+$\times\left(r^{\ell}+\mu^{2\varepsilon}\frac{K^{(B)}_{\ell}(\omega)}{M_{\rm Pl}^{2}}\frac{\ell+1}{2\ell}\frac{(\ell+1)!(D+\ell-2)}{\ell(\ell-1)}\frac{2^{\ell-2}\Gamma\left(\ell+\frac{D-3}{2}\right)}{\pi^{\frac{D-1}{2}}}r^{-\ell+3-D}\right),$
 The result (3.72) gives the one-point function of the magnetic component of the Weyl tensor in the presence of a tidal field, up to quadratic order in $\omega$, which we will use to match to the full black hole solution.
@@
 $h_{1}$ $=i\omega q_{\ell}\operatorname{e}^{-i\omega t}\mu^{\varepsilon}r^{3-\frac{D}{2}}\Psi_{\text{RW}},$
-where $q_{\ell}\equiv 1/\sqrt{2(\ell-1)(D-2+\ell)}$. The components (3.75) can be used to evaluate the Weyl tensor in Regge–Wheeler gauge. In particular, we find $B_{rij}$ to be
-$B_{rij}^{\text{RW}}$ $=\sum_{\ell=2}^{\infty}\left[r^{2}\partial_{r}\left(r^{-2}h_{0}(t,r)\right)-\partial_{0}h_{1}(t,r)\right]\nabla_{[i}Y_{j]}^{(T)}{}_{\ell}^{m}$ (3.76)
-$=-\sum_{\ell=2}^{\infty}\frac{q_{\ell}\operatorname{e}^{-i\omega t}}{4}\mu^{\varepsilon}r^{1-\frac{D}{2}}\left[\left(2D-D^{2}+4r^{2}\omega^{2}\right)\Psi_{\text{RW}}(r)+4r^{2}\Psi_{\text{RW}}^{\prime\prime}(r)\right]\nabla_{[i}Y_{j]}^{(T)}{}_{\ell}^{m}$.
+where $q_{\ell}\equiv 1/\sqrt{2(\ell-1)(D-2+\ell)}$. The components (3.75) can be used to evaluate the Weyl tensor in Regge–Wheeler gauge. In particular, we find $B_{rAB}$ to be
+$B_{rAB}^{\text{RW}}$ $=\frac{1}{M_{\rm Pl}}\sum_{\ell=2}^{\infty}\left[r^{2}\partial_{r}\left(r^{-2}h_{0}(t,r)\right)-\partial_{0}h_{1}(t,r)\right]\nabla_{[A}Y_{B]}^{(T)}{}_{\ell}^{m}$ (3.76)
+$=-\frac{1}{M_{\rm Pl}}\sum_{\ell=2}^{\infty}\frac{q_{\ell}\operatorname{e}^{-i\omega t}}{4}\mu^{\varepsilon}r^{1-\frac{D}{2}}\left[\left(2D-D^{2}+4r^{2}\omega^{2}\right)\Psi_{\text{RW}}(r)+4r^{2}\Psi_{\text{RW}}^{\prime\prime}(r)\right]\nabla_{[A}Y_{B]}^{(T)}{}_{\ell}^{m}$.
 Plugging in the homogeneous solution (3.74) for $\Psi_{\text{RW}}$, we get
-$B_{rij}^{\text{RW}}=-\sum_{\ell=2}^{\infty}q_{\ell}\operatorname{e}^{-i\omega t}\left[(\ell-1)(\ell+2-2\varepsilon)+r^{2}\omega^{2}\right]\left(B_{\text{reg}}r^{\ell}+\frac{\mu^{2\varepsilon}B_{\text{irr}}}{2\ell+1-2\varepsilon}r^{-\ell-1+2\varepsilon}\right)\nabla_{[i}Y_{j]}^{(T)}{}_{\ell}^{m}$.
+$B_{rAB}^{\text{RW}}=-\frac{1}{M_{\rm Pl}}\sum_{\ell=2}^{\infty}q_{\ell}\operatorname{e}^{-i\omega t}\left[(\ell-1)(\ell+2-2\varepsilon)+r^{2}\omega^{2}\right]\left(B_{\text{reg}}r^{\ell}+\frac{\mu^{2\varepsilon}B_{\text{irr}}}{2\ell+1-2\varepsilon}r^{-\ell-1+2\varepsilon}\right)\nabla_{[A}Y_{B]}^{(T)}{}_{\ell}^{m}$.
 (3.77)
-Since the linearized Weyl tensor is gauge invariant in flat space, we can directly compare $B_{rij}^{\text{RW}}$ with (3.72), obtained in de Donder gauge.
+Since the linearized Weyl tensor is gauge invariant in flat space, we can directly compare $B_{rAB}^{\text{RW}}$ with (3.72), obtained in de Donder gauge.
 By matching the $r^{\ell}$ and $r^{-\ell-1+2\varepsilon}$ coefficients in the two expressions, we can identify
-$\frac{(\ell+1)!(D+\ell-2)}{\ell(\ell-1)}\frac{2^{\ell-1}\Gamma\left(\ell+\frac{D-1}{2}\right)}{\pi^{\frac{D-1}{2}}}K^{(B)}_{\ell}(\omega)=\frac{B_{\text{irr}}}{B_{\text{reg}}}.$ (3.78)
+$\frac{\ell+1}{2\ell}\frac{(\ell+1)!(D+\ell-2)}{\ell(\ell-1)}\frac{2^{\ell-1}\Gamma\left(\ell+\frac{D-1}{2}\right)}{\pi^{\frac{D-1}{2}}}\frac{K^{(B)}_{\ell}(\omega)}{M_{\rm Pl}^{2}}=\frac{B_{\text{irr}}}{B_{\text{reg}}}.$ (3.78)
 This formula relates the response $K^{(B)}(\omega)$ to the coefficients of the homogeneous solution to the Regge–Wheeler equation.
@@
 (Note that we rescaled $B\to Br_{s}^{2}$ so that $B_{\text{reg}}$ is a polynomial in $r_{s}$.) Using (3.78), we finally obtain
-$\frac{45}{\pi}K^{(B)}_{\ell=2}(\omega)=i\omega r_{s}^{6}+\omega^{2}r_{s}^{7}\left(\frac{787}{2520}-\log(\mu r_{s})\right).$ (3.87)
+$\frac{3}{2}\frac{45}{2\pi M_{\rm Pl}^{2}}K^{(B)}_{\ell=2}(\omega)=i\omega r_{s}^{6}+\omega^{2}r_{s}^{7}\left(\frac{787}{2520}-\log(\mu r_{s})\right).$ (3.87)
 Repeating the same procedure for $\ell=3$ and $\ell=4$ (see Appendix C for details), we get:
-$\frac{18900}{\pi}K^{(B)}_{\ell=3}(\omega)$ $=i\omega r_{s}^{8}+\omega^{2}r_{s}^{9}\left(\frac{1\,727}{2\,520}-\log(\mu r_{s})\right),$ (3.88) $\frac{11113200}{\pi}K^{(B)}_{\ell=4}(\omega)$ $=i\omega r_{s}^{10}+\omega^{2}r_{s}^{11}\left(\frac{237\,529}{277\,200}-\log(\mu r_{s})\right).$
+$\frac{4}{3}\frac{18900}{2\pi M_{\rm Pl}^{2}}K^{(B)}_{\ell=3}(\omega)$ $=i\omega r_{s}^{8}+\omega^{2}r_{s}^{9}\left(\frac{1\,727}{2\,520}-\log(\mu r_{s})\right),$ (3.88) $\frac{5}{4}\frac{11113200}{2\pi M_{\rm Pl}^{2}}K^{(B)}_{\ell=4}(\omega)$ $=i\omega r_{s}^{10}+\omega^{2}r_{s}^{11}\left(\frac{237\,529}{277\,200}-\log(\mu r_{s})\right).$
 We see that, as expected, the magnetic dynamical love numbers are nonzero.
@@
 $K_{+-}^{(E)}(\tau_{2}-\tau_{1})=-G_{R}^{(Q_{E})}(\tau_{2}-\tau_{1})\equiv i\langle\left[Q_{E\,+}(\tau_{2}),Q_{E\,-}(\tau_{1})\right]\rangle\theta(\tau_{2}-\tau_{1}).$ (3.93)
-As in the magnetic sector, we can expand the response kernel in fourier space similarly to (3.53), where
+As in the magnetic sector, we can expand the response kernel in fourier space similarly to (3.53),
 $K_{{+-},\ell}^{(E)}(\tau_{2}-\tau_{1})$ $=\int\frac{{\rm d}{\omega}}{2\pi}\operatorname{e}^{-i{\omega}(\tau_{2}-\tau_{1})}K_{\ell}^{(E)}(\omega)$ (3.94)
-$=\frac{1}{\ell!}\int\frac{{\rm d}{\omega}}{2\pi}\operatorname{e}^{-i\tilde{\omega}(\tau_{2}-\tau_{1})}\left[\lambda_{0,\ell}^{E}+i{\omega}r_{s}\lambda_{1,\ell}^{E}+({\omega}r_{s})^{2}\lambda_{2,\ell}^{E}+\cdots\right],$
-$\lambda_{0,\ell}^{E}$ are the static Love numbers, $\lambda_{1,\ell}^{E}$ capture the linear-in-frequency dissipative response of the object, and $\lambda_{2,\ell}^{E}$ are the dynamical Love numbers at order $O(\omega^{2})$.
+$=\frac{1}{\ell!}\int\frac{{\rm d}{\omega}}{2\pi}\operatorname{e}^{-i\omega(\tau_{2}-\tau_{1})}\left[\lambda_{0,\ell}^{E}+i{\omega}r_{s}\lambda_{1,\ell}^{E}+({\omega}r_{s})^{2}\lambda_{2,\ell}^{E}+\cdots\right],$
+where $\lambda_{0,\ell}^{E}$ are the static Love numbers, $\lambda_{1,\ell}^{E}$ capture the linear-in-frequency dissipative response of the object, and $\lambda_{2,\ell}^{E}$ are the dynamical Love numbers at order $O(\omega^{2})$.
 Since we are interested in computing $\lambda_{2,\ell}^{E}$ for Schwarzschild black holes in four-dimensional general relativity, we will set $\lambda_{0,\ell}^{E}=0$ in the following.
@@
 From the definition of the Weyl tensor, we have on shell:
-$E_{ab}$ $=C_{0a0b}=2\partial_{a}\partial_{[0}h_{b]0}-2\partial_{0}\partial_{[0}h_{b]a}=-\partial_{a}\partial_{b}h_{00}+2\partial_{0}\partial_{(a}h_{b)0}+O(\omega^{2}),$ (3.95)
-$E_{A_{\ell}}$ $=-\partial_{A_{\ell}}h_{00}+2\partial_{0}\partial_{(A_{\ell-1}}h_{i_{\ell})_{T}0}+O(\omega^{2}).$
+$E_{ab}$ $=C_{0a0b}=\frac{2}{M_{\rm Pl}}\left[\partial_{a}\partial_{[0}h_{b]0}-\partial_{0}\partial_{[0}h_{b]a}\right]=\frac{1}{M_{\rm Pl}}\left[-\partial_{a}\partial_{b}h_{00}+2\partial_{0}\partial_{(a}h_{b)0}+O(\omega^{2})\right],$ (3.95)
+$E_{A_{\ell}}$ $=\frac{1}{M_{\rm Pl}}\left[-\partial_{A_{\ell}}h_{00}+2\partial_{0}\partial_{(A_{\ell-1}}h_{i_{\ell})_{T}0}+O(\omega^{2})\right].$
 As in the odd sector, we fix de Donder gauge and expand $\langle E^{+}_{ab}(t,\vec{x})E^{-}_{A_{\ell}}\left(\tau_{2}\right)\rangle\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu^{+\,A_{\ell}}\left(\tau_{1}\right)$ in (3.92) up to order $O(\omega)$.
@@
 Then, from the contraction of the Weyl $E$-operators we find
-$E_{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu^{A_{\ell}}=\partial_{A_{\ell}}h_{00}\partial^{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{00}-2\partial_{A_{\ell}}h_{00}\partial^{0}\partial^{A_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{i_{\ell}0}-2\partial_{0}\partial_{A_{\ell-1}}h_{i_{\ell}0}\partial^{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{00}+O(\omega^{2}).$ (3.96)
+$E_{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu^{A_{\ell}}=\frac{1}{M_{\rm Pl}^{2}}\left[\partial_{A_{\ell}}h_{00}\partial^{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{00}-2\partial_{A_{\ell}}h_{00}\partial^{0}\partial^{A_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{i_{\ell}0}-2\partial_{0}\partial_{A_{\ell-1}}h_{i_{\ell}0}\partial^{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{00}+O(\omega^{2})\right].$ (3.96)
 Putting this together, we find the contraction
-$\langle E_{+\,ab}(t,\vec{x})E_{-\,A_{\ell}}\left(\tau_{2}\right)\rangle\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu_{+}^{A_{\ell}}\left(\tau_{1}\right)=-\langle\partial_{a}\partial_{b}h_{+\,00}\partial_{A_{\ell}}h_{-\,00}\rangle\left(\partial^{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{00}-2\partial^{0}\partial^{A_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{i_{\ell}0}\right)+O(\omega^{2}),$ (3.97)
+$\langle E_{+\,ab}(t,\vec{x})E_{-\,A_{\ell}}\left(\tau_{2}\right)\rangle\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu_{+}^{A_{\ell}}\left(\tau_{1}\right)=-\frac{1}{M_{\rm Pl}^{3}}\langle\partial_{a}\partial_{b}h_{+\,00}\partial_{A_{\ell}}h_{-\,00}\rangle\left(\partial^{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{00}-2\partial^{0}\partial^{A_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{i_{\ell}0}\right)+O(\omega^{2}),$ (3.97)
 where we dropped terms that are either of order $O(\omega^{2})$ or proportional to $\mathcal{P}^{\rm dD}_{0a00}=0$, with $\mathcal{P}_{\mu\nu\rho\sigma}^{\text{dD}}$ defined as in (3.62).
@@
 where $\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{00}^{\text{RW}}$ is the $tt$-component of the tidal field metric perturbation in Regge–Wheeler gauge. We can express this as [2]
-$\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{00}^{\text{RW}}(t,\vec{x})=\operatorname{e}^{-i\omega t}\sum_{\ell}c_{\rm ext}Y_{\ell}^{m}r^{\ell}=\operatorname{e}^{-i\omega t}\sum_{\ell}c_{j_{1}\cdots j_{\ell}}x^{j_{1}}\cdots x^{j_{\ell}}+O(\omega^{2}),$ (3.99)
+$\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{00}^{\text{RW}}(t,\vec{x})=M_{\rm Pl}\operatorname{e}^{-i\omega t}\sum_{\ell}c_{\rm ext}Y_{\ell}^{m}r^{\ell}=M_{\rm Pl}\operatorname{e}^{-i\omega t}\sum_{\ell}c_{j_{1}\cdots j_{\ell}}x^{j_{1}}\cdots x^{j_{\ell}}+O(\omega^{2}),$ (3.99)
 where $c^{i_{1}\cdots i_{\ell}}$ is a traceless symmetric tensor. Plugging this into (3.92) for the one-point function and using the expression (3.61) for the graviton propagator, we obtain
-$\langle E_{+\,ab}(t,\vec{x})\rangle_{\text{in-in}}=-\operatorname{e}^{-i\omega t}\frac{D-3}{D-2}\sum_{\ell=2}^{\infty}\ell!(-i)^{\ell}K_{\ell}^{(E)}(\omega)c^{i_{1}\cdots i_{\ell}}\partial_{a}\partial_{b}\int\frac{{\rm d}^{D-1}\vec{p}}{(2\pi)^{D-1}}\operatorname{e}^{i\vec{p}\cdot\vec{x}}\frac{p_{i_{1}}\cdots p_{i_{\ell}}}{\vec{p}^{2}}$ (3.100)
+$\langle E_{+\,ab}(t,\vec{x})\rangle_{\text{in-in}}=-\frac{\operatorname{e}^{-i\omega t}}{M_{\rm Pl}^{2}}\frac{D-3}{D-2}\sum_{\ell=2}^{\infty}\ell!(-i)^{\ell}K_{\ell}^{(E)}(\omega)c^{i_{1}\cdots i_{\ell}}\partial_{a}\partial_{b}\int\frac{{\rm d}^{D-1}\vec{p}}{(2\pi)^{D-1}}\operatorname{e}^{i\vec{p}\cdot\vec{x}}\frac{p_{i_{1}}\cdots p_{i_{\ell}}}{\vec{p}^{2}}.$ (3.100)
 Going back to position space, this is
-$\langle E_{+\,ab}(t,\vec{x})\rangle_{\text{in-in}}=-\operatorname{e}^{-i\omega t}\sum_{\ell=2}^{\infty}K_{\ell}^{(E)}(\omega)\frac{2^{\ell-2}\ell!\,\Gamma\left(\ell+\frac{D-3}{2}\right)}{\pi^{\frac{D-1}{2}}}\frac{D-3}{D-2}c_{i_{1}\cdots i_{\ell}}\partial_{a}\partial_{b}\frac{x^{i_{1}}\cdots x^{i_{\ell}}}{|\vec{x}|^{2\ell+D-3}}.$ (3.101)
+$\langle E_{+\,ab}(t,\vec{x})\rangle_{\text{in-in}}=-\frac{\operatorname{e}^{-i\omega t}}{M_{\rm Pl}^{2}}\sum_{\ell=2}^{\infty}K_{\ell}^{(E)}(\omega)\frac{2^{\ell-2}\ell!\,\Gamma\left(\ell+\frac{D-3}{2}\right)}{\pi^{\frac{D-1}{2}}}\frac{D-3}{D-2}c_{i_{1}\cdots i_{\ell}}\partial_{a}\partial_{b}\frac{x^{i_{1}}\cdots x^{i_{\ell}}}{|\vec{x}|^{2\ell+D-3}}.$ (3.101)
 We can similarly write the background expression of the Weyl tensor. Focusing on the $rr$-component,
-$\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu_{rr}$ $=-\partial_{r}^{2}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{00}+2\partial_{0}\partial_{r}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{0r}+O(\omega^{2})$ (3.102)
-$=-\partial_{r}^{2}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{00}^{\text{RW}}+O(\omega^{2})=-\operatorname{e}^{-i\omega\tau}\sum_{\ell=2}^{\infty}c_{\rm ext}\,\ell(\ell-1)r^{\ell-2}Y_{\ell}^{m}+O(\omega^{2}).$
+$\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu_{rr}$ $=\frac{1}{M_{\rm Pl}^{2}}\left(-\partial_{r}^{2}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{00}+2\partial_{0}\partial_{r}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{0r}\right)+O(\omega^{2})$ (3.102)
+$=-\frac{1}{M_{\rm Pl}^{2}}\partial_{r}^{2}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{00}^{\text{RW}}+O(\omega^{2})=-\operatorname{e}^{-i\omega\tau}\sum_{\ell=2}^{\infty}c_{\rm ext}\,\ell(\ell-1)r^{\ell-2}Y_{\ell}^{m}+O(\omega^{2}).$
 Using $c_{j_{1}\cdots j_{\ell}}x^{j_{1}}\cdots x^{j_{\ell}}=c_{\rm ext}r^{\ell}Y_{\ell}^{m}$ with $r\equiv|\vec{x}|$, and $\nabla_{r}\nabla_{r}=\partial_{r}\partial_{r}$ in (3.101), and adding it to (3.102), we get
-$\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu_{rr}+\langle E_{+\,rr}\rangle_{\text{in-in}}$ $=-c_{\text{ext}}\operatorname{e}^{-i\omega t}\sum_{\ell=2}^{\infty}\ell(\ell-1)Y_{\ell}^{m}$ (3.103)
-$\!\!\!\times\left(r^{\ell-2}+\mu^{2\varepsilon}K^{(E)}_{\ell}(\omega)\frac{D-3}{D-2}\frac{(\ell+D-3)(\ell+D-2)}{\ell(\ell-1)}\frac{2^{\ell-2}\ell!\,\Gamma\left(\ell+\frac{D-3}{2}\right)}{\pi^{\frac{D-1}{2}}}r^{-\ell-D+1}\right),$
+$\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu_{rr}+\langle E_{+\,rr}\rangle_{\text{in-in}}$ $=-\operatorname{e}^{-i\omega t}\sum_{\ell=2}^{\infty}c_{\rm ext}\ell(\ell-1)Y_{\ell}^{m}$ (3.103)
+$\!\!\!\times\left(r^{\ell-2}+\mu^{2\varepsilon}\frac{K^{(E)}_{\ell}(\omega)}{M_{\rm Pl}^{2}}\frac{D-3}{D-2}\frac{(\ell+D-3)(\ell+D-2)}{\ell(\ell-1)}\frac{2^{\ell-2}\ell!\,\Gamma\left(\ell+\frac{D-3}{2}\right)}{\pi^{\frac{D-1}{2}}}r^{-\ell-D+1}\right),$
 where we replaced $K^{(E)}_{\ell}$ by $\mu^{2\varepsilon}K^{(E)}_{\ell}$. As before, we need to include gravity, renormalize, and match.
@@
 In particular, to leading order in the flat-space limit,
-$E_{rr}=-\partial_{r}^{2}H_{0}-2i\omega\partial_{r}H_{1}.$ (3.106)
+$E_{rr}=-\frac{1}{M_{\rm Pl}}\left(\partial_{r}^{2}H_{0}+2i\omega\partial_{r}H_{1}\right).$ (3.106)
 Using the constraint equations (see eq. (3.54) of [7]),
@@
 Comparing the resulting Weyl tensor of the homogeneous solution (3.105) with (3.103), we can extract $K^{(E)}_{\ell}$ as a function of the ratio $B_{\text{irr}}/B_{\text{reg}}$:
-$\frac{D-3}{D-2}\frac{(\ell+D-3)(\ell+D-2)}{\ell(\ell-1)}\frac{2^{\ell-1}\ell!\,\Gamma\left(\ell+\frac{D-1}{2}\right)}{\pi^{\frac{D-1}{2}}}K^{(E)}_{\ell}(\omega)=\frac{B_{\text{irr}}}{B_{\text{reg}}}.$ (3.109)
+$\frac{D-3}{D-2}\frac{(\ell+D-3)(\ell+D-2)}{\ell(\ell-1)}\frac{2^{\ell-1}\ell!\,\Gamma\left(\ell+\frac{D-1}{2}\right)}{\pi^{\frac{D-1}{2}}}\frac{K^{(E)}_{\ell}(\omega)}{M_{\rm Pl}^{2}}=\frac{B_{\text{irr}}}{B_{\text{reg}}}.$ (3.109)
 This relates the worldline response to the parameters of the Zerilli solution.
@@
 and therefore, from (3.103), this implies the following value for the $\ell=2$ electric type Love number
-$\frac{45}{2\pi}K^{(E)}_{\ell=2}(\omega)=i\omega r_{s}^{6}-\omega^{2}r_{s}^{7}\left(-\frac{797}{1260}+\log(\mu r_{s})\right).$ (3.118)
+$\frac{45}{2\pi M_{\rm Pl}^{2}}K^{(E)}_{\ell=2}(\omega)=i\omega r_{s}^{6}-\omega^{2}r_{s}^{7}\left(-\frac{797}{1260}+\log(\mu r_{s})\right).$ (3.118)
+Note that the linear term in the frequency and the coefficient of the logarithm agree with [49] (see also [40]) after accounting for the different conventions.
+[^33] In particular, in the notation of [49], $\tilde{F}(\omega)=4K^{(E)}_{\ell=2}(\omega)$, where the factor of $4$ follows from canonical normalization of the graviton field.
 Repeating the same procedure for $\ell=3$ and $\ell=4$ (see Appendix C for details) we get:
-$\frac{18900}{2\pi}K^{(E)}_{\ell=3}(\omega)$ $=i\omega r_{s}^{8}-\omega^{2}r_{s}^{9}\left(-\frac{709}{840}+\log(\mu r_{s})\right),$ (3.119) $\frac{11113200}{2\pi}K^{(E)}_{\ell=4}(\omega)$ $=i\omega r_{s}^{10}-\omega^{2}r_{s}^{11}\left(-\frac{5\,501}{5\,775}+\log(\mu r_{s})\right).$
+$\frac{18900}{2\pi M_{\rm Pl}^{2}}K^{(E)}_{\ell=3}(\omega)$ $=i\omega r_{s}^{8}-\omega^{2}r_{s}^{9}\left(-\frac{709}{840}+\log(\mu r_{s})\right),$ (3.119) $\frac{11113200}{2\pi{M_{\rm Pl}^{2}}}K^{(E)}_{\ell=4}(\omega)$ $=i\omega r_{s}^{10}-\omega^{2}r_{s}^{11}\left(-\frac{5\,501}{5\,775}+\log(\mu r_{s})\right).$
 We see that the dynamical Love numbers are nonzero, and their logarithmic running is given by the dissipative response coefficients.
 3.3 Symmetries
-Aside from the fact that they are nonzero, the dynamical Love numbers that we have computed have some properties worth remarking on.
+Aside from the fact that they are nonzero, the dynamical Love numbers have some properties worth remarking on.
 Most notably, if we expand the response kernels $K$ as in (3.53) and (3.94) and then use (3.87) and (3.118) along with (3.88) and (3.119) to determine the effective worldline couplings, we see that $\lambda_{1,\ell}^{E}=\lambda_{1,\ell}^{B}$ and the coefficients of the logs in $\lambda_{2,\ell}^{E}$ are exactly the same as those in $\lambda_{2,\ell}^{B}$.
-[^28] In making this comparison, there is a relative normalization factor of $1/2$ in the magnetic $\lambda$ coefficients coming from the relation between the Weyl tensor and gravito-magnetic field.
-That is, the dynamical Love numbers in the even and odd sectors are equal—this can be understood as a consequence of Chandrasekhar’s duality.
+[^34] In making this comparison, there is a relative normalization factor of $\ell/(\ell+1)$ in the definition of the magnetic $\lambda$ coefficients coming from the relation between the Weyl tensor and gravito-magnetic field.
+In addition, recall that in flat space Chandrasekhar’s duality acts by rotating the Riemann tensor and its dual, so that in particular $\delta E_{ab}=-\frac{1}{2}{\epsilon_{0a}}^{cd}B_{bcd}$ and $\delta B_{abc}={\epsilon^{0d}}_{bc}E_{ad}$ (see e.g.
+(3.72) and (3.103)).
+That is, the dynamical Love numbers in the even and odd sectors are equal.
+This can be understood as a consequence of Chandrasekhar’s duality.
 As we mentioned in Section 3.1.2, Chandrasekhar discovered a mapping between the Regge–Wheeler and Zerilli equations [76, 77, 78].
 There we used this fact as a solution-generating technique to obtain solutions in the Zerilli sector from those in the Regge–Wheeler sector.
-Here we note that this relation can also be read as a symmetry [7].
+Here we note that this relation can also be read as a symmetry [7, 126].
 Specifically, both the Regge–Wheeler potential and Zerilli potential in $D=4$ can be written as
@@
 At large distances it interchanges the electric and magnetic parts of the Weyl tensor, and so is a Schwarzschild generalization of electric-magnetic duality.
-[^29] Note that it keeps the background fixed, it does not map Schwarzschild into Taub–NUT.
+[^35] Note that this transformation keeps the background fixed, it does not map Schwarzschild into Taub–NUT.
 It is therefore directly a symmetry of linearized perturbations around Schwarzschild.
@@
 However, we computed these quantities via dimensional regularization, and the symmetry (3.121) exists only in $D=4$.
-Since we have chosen a regularization scheme that breaks the symmetry, it is not surprising that the final answers do not respect it.
-Instead, the existence of the symmetry suggests that a scheme exists that preserves the symmetry.
+Since we have chosen a regularization scheme that breaks the symmetry, it is not surprising that the scheme-dependent parts of the final answers do not respect it.
+The existence of the symmetry implies the existence of a scheme that preserves the symmetry.
 Indeed, an important consistency check is that the scheme-independent parts of the response—the imaginary dissipative response and the coefficient of the logarithms—match exactly between the two sectors.
@@
 We dealt with the nonlinearities of gravitational interactions in the EFT by utilizing the Born series approach of [75, 74], and we find also that the field profiles in the EFT must be regularized and renormalized before they can be matched to the general relativity solution.
-What we find after matching to define the dynamical Love numbers of black holes is that they have several interesting properties.
+After matching to define the dynamical Love numbers of black holes, we find they have several interesting properties.
 They are nonzero, and have a characteristic running, where the coefficient of the logarithm in all cases is $\omega r_{s}$ times the corresponding dissipative response coefficient.
 One way to read this relation is that the beta function for the dynamical Love numbers is determined by the dissipative response.
-However, this appears to become more complicated at subleading order [74].
-Another way to read this relation is that the discontinuity of the conservative part of $K$ is $\beta\omega/2$ times the dissipative part of $K$, where $\beta=4\pi r_{s}$ is the inverse black hole temperature.
+[^36] This is rather surprising because the long-distance running of the conservative response is sourced by the Schwarzschild potential and hence is universal, while the dissipative response is a finite-size effect particular to black holes.
+This relation appears to become more complicated at subleading order [74].
+Another reading is that the discontinuity of the conservative part of $K$ is $\beta\omega/2$ times the dissipative part of $K$, where $\beta=4\pi r_{s}$ is the inverse black hole temperature.
 [^37] This looks suspiciously like a fluctuation–dissipation relation.
 However, it is a relation involving just the causal Green’s function.
-(Interestingly this relation continues to hold for the most transcendental parts of the response kernel at subleading orders in frequency, at least in the examples of [74].) It would be nice to understand the features of these logarithmic terms better.
+(Interestingly this relation continues to hold for the most transcendental parts of the response kernel at subleading orders in frequency, at least in the examples of [74].) It would be nice to understand these features of the logarithmic terms better.
 Another interesting feature is that, as expected from the symmetries of black hole perturbations, there exists a scheme in which the electric and magnetic dynamical Love numbers (along with the dissipative responses) are equal.
@@
 In principle there are only technical complications to doing so, and it would be very interesting to see if the patterns seen in the four-dimensional case survive there.
-Relatedly, one could carry out the computation of dynamical responses to external electromagnetic fields and consider the generalization to charged black hole responses.
+Relatedly, one could compute dynamical responses to external electromagnetic fields, and consider the generalization to charged black hole responses.
 Perhaps the most phenomenologically interesting generalization would be to compute the full dynamical Love numbers for a spinning Kerr black hole.
-To do this, the main advances required are technical.
-In particular, re-summing and regulating gravitational effects in the EFT is most simply done in dimensional regularization, but gravitational couplings to spinning black holes in higher dimensions are somewhat complicated.
-At leading order, we can extract a Kerr black hole’s dynamical Love number from those of a Schwarzschild black hole by exploiting the fact that a Kerr black hole is a spinning Schwarzschild black hole.
-The coordinate transformation between the (rotating) black hole’s rest frame and the laboratory frame captures the leading-in-spin effect (much in the same way that a Kerr black hole’s zero frequency dissipation is related to the finite frequency Schwarzschild dissipation) [46, 19].
-Nevertheless the computation of the full response for Kerr is a problem that we hope to return to.
+The main advances required are technical.
+In particular, resumming and regulating gravitational effects in the EFT is simplest in dimensional regularization, but gravitational couplings to spinning black holes in higher dimensions are complicated.
+At leading order, we can extract a Kerr black hole’s dynamical Love number by exploiting the fact that a Kerr black hole is a spinning Schwarzschild black hole.
+The coordinate transformation between the (rotating) black hole’s rest frame and the laboratory frame captures the leading-in-spin effect (similar to how a Kerr black hole’s zero frequency dissipation is related to the finite frequency Schwarzschild dissipation [46, 19]).
+The computation of the full response for Kerr is a problem that we hope to return to.
 One of the most compelling reasons to study the material properties of black holes is that these quantities are in principle measurable.
@@
 Acknowledgements:
-Thanks to John Joseph Carrasco, Lam Hui, Riccardo Penco, and Mikhail Solon for helpful conversations AJ and DG are supported in part by DOE (HEP) Award DE-SC0025323.
+Thanks to John Joseph Carrasco, Lam Hui, Riccardo Penco, Chia-Hsien Shen, Mikhail Solon, and Zihan Zhou for helpful conversations, and to Giulia Isabella, Massimiliano M.
+Riva, and Jan Steinhoff for comments on the draft.
+AJ and DG are supported in part by DOE (HEP) Award DE-SC0025323.
 The work of MJR is supported through the NSF grant PHY-2309270.
@@
 The work is supported by a PhD Joint Program of the International Research Center for Fundamental Scientific Discovery (IRC Discovery).
+Note added:
+Shortly after this paper, [128] appeared, which computes the odd-parity magnetic dynamical tidal response, with a slightly different perspective and formalism.
+Where our results overlap, they are in agreement.
 Appendix A Scalar dynamical response running
@@
 In conclusion, no $\log(x)/x^{\ell+1}$ term results from the $x\rightarrow\infty$ expansion of the $I_{n}$ integrals (A.6), for all $n$.
-The same conclusion is trivially true for the second integral in (A.4), since both $P_{\ell}(y)$ and $W_{\ell-1}(y)$ are polynomials in (y1)(y-1).
+The second integral in (A.4) can also be computed:
+$\int^{x}{\rm d}y\frac{(1+y)^{3}}{y-1}P_{\ell}(y)W_{\ell-1}(y)=-\sum_{k=0}^{2\ell-1}d_{k}\,q_{k}(x)-\sum_{k=1}^{2\ell-1}\frac{8}{k}d_{k}\,(x-1)^{k}+8\,c_{0}\log(x-1),$ (A.9)
+where $q_{k}$ denotes the polynomial
+$q_{k}(y)\equiv\left(\frac{12(y-1)}{1+k}+\frac{6(y-1)^{2}}{2+k}+\frac{(y-1)^{3}}{3+k}\right)(y-1)^{k}\,$, (A.10)
+and the coefficients in the series expansion are
+$d_{k}=\frac{1}{2^{k}}\sum_{n=0}^{k}\frac{\Gamma(\ell+1+n)\,\Gamma(\ell+1+k-n)}{\Gamma(\ell+1-n)\,\Gamma(\ell+1-(k-n))\,(n!)^{2}\,((k-n)!)^{2}}\bigl(\psi(\ell+1)-\psi(k-n+1)\bigr).$ (A.11)
+From this expression, we can now make the following consideration.
+As noted before, for the first integral we argue that the same conclusion is trivially true for the second integral in (A.4), since both $P_{\ell}(y)$ and $W_{\ell-1}(y)$ are polynomials in (y1)(y-1).
 Therefore, there is no $\log(x)/x^{\ell+1}$ type of term resulting from the first integral in (2.21).
@@
 Explicitly,
-$\frac{B}{4}\int^{x}{\rm d}y\frac{(1+y)^{3}}{y-1}P_{\ell}(y)^{2}=2B\log(x-1)+\text{pol}(x).$ (A.10)
+$\frac{B}{4}\int^{x}{\rm d}y\frac{(1+y)^{3}}{y-1}P_{\ell}(y)^{2}$ $=\frac{B}{4}\left(\sum_{m=0}^{2\ell}f_{m}\,q_{m}(x)+\sum_{m=1}^{2\ell}\frac{8}{m}f_{m}(x-1)^{m}+8\log(x-1)\right)$
+$=2B\log(x-1)+\text{pol}(x),$ (A.13)
+where the coefficients $f_{m}$ from the convolution of the series expansions of $P_{\ell}^{2}$ take the form
+$f_{m}=\sum_{n=0}^{m}a_{n}\,a_{m-n}.$ (A.14)
 From the large-$x$ expansion of the $Q_{\ell}(x)$ functions,
@@
 where $\tilde{F}$ denotes the regularized function, defined by
-$_{p}\tilde{F}_{q}\left[\begin{array}[]{c}a_{1}\,,\,\cdots\,,\,a_{p}\\[-2.0pt] b_{1}\,,\,\cdots\,,\,b_{q}\end{array}\bigg\rvert\,x\,\right]\equiv\frac{\,{}_{p}F_{q}\left[\begin{array}[]{c}a_{1}\,,\,\cdots\,,\,a_{p}\\[-2.0pt] b_{1}\,,\,\cdots\,,\,b_{q}\end{array}\bigg\rvert\,x\,\right]}{\Gamma(b_{1})\cdots\Gamma(b_{q})}.$ (B.17)
+$_{p}\tilde{F}_{q}\left[\begin{array}[]{c}a_{1}\,,\,\cdots\,,\,a_{p}\\[-2.0pt] b_{1}\,,\,\cdots\,,\,b_{q}\end{array}\bigg\rvert\,x\,\right]\equiv\Big[\Gamma(b_{1})\cdots\Gamma(b_{q})\Big]^{-1}\,{}_{p}F_{q}\left[\begin{array}[]{c}a_{1}\,,\,\cdots\,,\,a_{p}\\[-2.0pt] b_{1}\,,\,\cdots\,,\,b_{q}\end{array}\bigg\rvert\,x\,\right].$ (B.17)
 To match with the intermediate-zone solution, we need the small-$z$ expansion of the solution (B.14). First, let us use the series representation of the generalized hypergeometric functions,
@@
 $g_{1,k}=\frac{(-1)^{k}(\ell)_{k}}{\Gamma(\ell+k+\frac{3}{2})\Gamma(2\ell+k+2)k!},\qquad g_{2,k}=\frac{(-1)^{k}((\ell+1)_{k})^{2}}{\Gamma(\ell+k+\frac{3}{2})\Gamma(2\ell+k+2)\Gamma(\ell+k+2)k!}.$ (B.22)
-To expand the Coulomb functions, we can then use their relations to spherical Bessel functions, i.e.,
+To expand the Coulomb functions, we can then use their relations to spherical Bessel functions:
 $F_{\ell}(0,z)=zj_{\ell}(z),\qquad\qquad G_{\ell}(0,z)=-zy_{\ell}(z),$ (B.23)
@@
 $+\left[c_{2}^{(1)}z^{-\ell}-\frac{c_{1}^{(0)}\sqrt{\pi}}{4}\left(\sum_{k=1}^{\infty}\gamma_{k}^{\ell}z^{\ell+2k}+(\ell+1)!g_{1,0}z^{\ell}\right)\right]\sum_{n=0}^{\infty}y_{n}^{\ell}z^{2n}.$
-Plugging this into the definition of $\phi^{(1)}_{\ell,\text{FZ}}=\frac{\omega}{z}(\xi^{(1)}_{\ell}+\frac{1}{2z}\xi^{(0)}_{\ell})$, we obtain the following $z\rightarrow 0$ behavior for $\phi_{\ell,\text{FZ}}^{(1)}$ [^36] We use $y_{0}^{\ell}=(2\ell-1)!!=\frac{(2\ell+1)!!}{2\ell+1}$ to simplify the expressions.
+Plugging this into the definition of $\phi^{(1)}_{\ell,\text{FZ}}=\frac{\omega}{z}(\xi^{(1)}_{\ell}+\frac{1}{2z}\xi^{(0)}_{\ell})$, we obtain the following $z\rightarrow 0$ behavior for $\phi_{\ell,\text{FZ}}^{(1)}$ [^43] We use $y_{0}^{\ell}=(2\ell-1)!!=\frac{(2\ell+1)!!}{2\ell+1}$ to simplify the expressions.:
 $\phi_{\ell,\text{FZ}}^{(1)}(z)\underset{z\to 0}{\sim}$ $\,\omega\Bigg[c_{2}^{(1)}z^{-\ell-1}\left(y_{0}^{\ell}+\sum_{n=1}^{\ell}y_{n}^{\ell}z^{2n}\right)-c_{1}^{(0)}z^{\ell-1}\left(\frac{\ell(\ell+1)}{2\ell+1}j_{0}^{\ell}+\frac{\sqrt{\pi}(\ell+1)!}{4}g_{1,0}y_{0}^{\ell}-\frac{1}{2(2\ell+1)!!}\right)$
-$\quad+c_{1}^{(1)}j_{0}^{\ell}z^{\ell}+O(z^{\ell+1})\Bigg]$
-$=$ $\,\,\omega c_{2}^{(1)}\frac{(2\ell+1)!!}{2\ell+1}z^{-\ell-1}\left(1+\sum_{n=1}^{\ell-1}\frac{y_{n}^{\ell}}{y_{0}^{\ell}}z^{2n}\right)-\omega c_{1}^{(0)}\frac{\ell}{2(2\ell+1)!!}z^{\ell-1}$
-$+\frac{\omega c_{1}^{(1)}}{(2\ell+1)!!}z^{\ell}+O(z^{\ell+1}).$ (B.27)
+$\quad+c_{1}^{(1)}j_{0}^{\ell}z^{\ell}+O(z^{\ell+1})\Bigg]$ (B.27)
+$=$ $\,\,\omega c_{2}^{(1)}\frac{(2\ell+1)!!}{2\ell+1}z^{-\ell-1}\left(1+\sum_{n=1}^{\ell-1}\frac{y_{n}^{\ell}}{y_{0}^{\ell}}z^{2n}\right)-\omega c_{1}^{(0)}\frac{\ell}{2(2\ell+1)!!}z^{\ell-1}+\frac{\omega c_{1}^{(1)}}{(2\ell+1)!!}z^{\ell}+O(z^{\ell+1}).$
 The integration constants $c_{1}^{(1)}$ and $c_{2}^{(1)}$ can be determined by comparing $\phi_{\ell,\text{FZ}}^{(1)}$ with the intermediate-zone solution:
@@
 Even sector: Similarly we tabulate the solutions to the Zerilli equation in the even sector. The solution for $\ell=3$ is
+$\Psi_{\text{Z,IZ}}^{\ell=3}(r)\underset{\frac{r}{r_{s}}\to\infty}{=}\frac{C}{50}\bigg(300\frac{r^{4}}{r_{s}^{4}}-190\frac{r^{3}}{r_{s}^{3}}-93\frac{r^{2}}{r_{s}^{2}}+\frac{279}{10}\frac{r}{r_{s}}+\frac{663}{100}-\frac{1\,989}{1\,000}\frac{r_{s}}{r}+\frac{5\,967}{10\,000}\frac{r_{s}^{2}}{r^{2}}-\frac{17\,901}{100\,000}\frac{r_{s}^{3}}{r^{3}}\bigg)$
 […]
+$\penalty 10000\ \hskip 2.84544pt+C\omega^{2}r_{s}^{2}\Bigg[-\frac{1}{3}\frac{r^{6}}{r_{s}^{6}}-\frac{11}{10}\frac{r^{5}}{r_{s}^{5}}-\frac{r^{4}}{r_{s}^{4}}\!\left(\pi^{2}+\frac{51\,917}{2\,100}-\frac{13}{7}\log\frac{r_{s}}{r}\right)+\frac{r^{3}}{r_{s}^{3}}\!\left(\frac{19\pi^{2}}{30}+\frac{399\,451}{21\,000}-\frac{247}{210}\log\frac{r_{s}}{r}\right)$
+$\penalty 10000\ \hskip 52.63777pt+\frac{r^{2}}{r_{s}^{2}}\!\left(\frac{31\pi^{2}}{100}+\frac{1\,634\,147}{210\,000}-\frac{403}{700}\log\frac{r_{s}}{r}\right)-\frac{r}{10^{3}r_{s}}\!\left(93\pi^{2}+\frac{2\,039\,147}{700}-\frac{1\,209}{7}\log\frac{r_{s}}{r}\right)$
 […]
-$\penalty 10000\ \hskip 78.24507pt+\frac{13\,923}{500\,000}\frac{r_{s}^{2}}{r^{2}}+\frac{1\,622\,851}{105\,000\,000}\frac{r_{s}^{3}}{r^{3}}\bigg)$
-$\penalty 10000\ \hskip 28.45274pt+C\omega^{2}r_{s}^{2}\bigg[-\frac{1}{3}\frac{r^{6}}{r_{s}^{6}}-\frac{11}{10}\frac{r^{5}}{r_{s}^{5}}-\frac{r^{4}}{r_{s}^{4}}\!\left(\pi^{2}+\frac{51\,917}{2\,100}-\frac{13}{7}\log\frac{r_{s}}{r}\right)$
-$\penalty 10000\ \hskip 78.24507pt+\frac{r^{3}}{r_{s}^{3}}\!\left(\frac{19\pi^{2}}{30}+\frac{399\,451}{21\,000}-\frac{247}{210}\log\frac{r_{s}}{r}\right)$
-$\penalty 10000\ \hskip 78.24507pt+\frac{r^{2}}{r_{s}^{2}}\!\left(\frac{31\pi^{2}}{100}+\frac{1\,634\,147}{210\,000}-\frac{403}{700}\log\frac{r_{s}}{r}\right)$
-$\penalty 10000\ \hskip 78.24507pt-\frac{r}{10^{3}r_{s}}\!\left(93\pi^{2}+\frac{2\,039\,147}{700}-\frac{1\,209}{7}\log\frac{r_{s}}{r}\right)$
-$\penalty 10000\ \hskip 78.24507pt-\frac{221\pi^{2}}{10\,000}-\frac{10\,422\,677}{21\,000\,000}+\frac{2\,873}{70\,000}\log\frac{r_{s}}{r}$
-$\penalty 10000\ \hskip 78.24507pt+\frac{r_{s}}{10^{5}r}\!\left(663\pi^{2}+\frac{21\,972\,677}{700}-\frac{8\,619}{7}\log\!\left(\frac{r_{s}}{r}\right)\right)$
-$\penalty 10000\ \hskip 78.24507pt+\frac{r_{s}^{2}}{10^{6}r^{2}}\!\left(-1\,989\pi^{2}+\frac{220\,495\,907}{2\,100}+\frac{25\,857}{7}\log\frac{r_{s}}{r}\right)$
-$\penalty 10000\ \hskip 78.24507pt+\frac{r_{s}^{3}}{10^{7}r^{3}}\!\left(5\,967\pi^{2}+\frac{58\,238\,757\,859}{44\,100}-\frac{5\,232\,713}{21}\log\frac{r_{s}}{r}\right)\bigg]+O\!\left(\frac{r_{s}^{4}}{r^{4}}\right),$
+$\penalty 10000\ \hskip 52.63777pt+\frac{r_{s}^{2}}{10^{6}r^{2}}\!\left(-1\,989\pi^{2}+\frac{220\,495\,907}{2\,100}+\frac{25\,857}{7}\log\frac{r_{s}}{r}\right)$
+$\penalty 10000\ \hskip 52.63777pt+\frac{r_{s}^{3}}{10^{7}r^{3}}\!\left(5\,967\pi^{2}+\frac{58\,238\,757\,859}{44\,100}-\frac{5\,232\,713}{21}\log\frac{r_{s}}{r}\right)\Bigg]+O\!\left(\frac{r_{s}^{4}}{r^{4}}\right),$ (C.3)
 and the solution for $\ell=4$ is
+$\Psi^{\ell=4}_{\text{Z,IZ}}(r)\underset{\frac{r}{r_{s}}\to\infty}{=}C\bigg(28\frac{r^{5}}{r_{s}^{5}}-\frac{119}{3}\frac{r^{4}}{r_{s}^{4}}+\frac{179}{18}\frac{r^{3}}{r_{s}^{3}}+\frac{361}{108}\frac{r^{2}}{r_{s}^{2}}-\frac{361}{648}\frac{r}{r_{s}}$
+$\penalty 10000\ \hskip 44.10185pt-\frac{287}{3\,888}+\frac{287}{23\,328}\frac{r_{s}}{r}-\frac{287}{139\,968}\frac{r_{s}^{2}}{r^{2}}+\frac{287}{839\,808}\frac{r_{s}^{3}}{r^{3}}-\frac{287}{5\,038\,848}\frac{r_{s}^{4}}{r^{4}}\bigg)$
+$\penalty 10000\ \hskip 2.84544pt+Ci\omega r_{s}\Bigg[\frac{1\,246}{15}\frac{r^{5}}{r_{s}^{5}}-\frac{10\,591}{90}\frac{r^{4}}{r_{s}^{4}}+\frac{15\,931}{540}\frac{r^{3}}{r_{s}^{3}}+\frac{32\,129}{3\,240}\frac{r^{2}}{r_{s}^{2}}-\frac{32\,129}{19\,440}\frac{r}{r_{s}}$
 […]
-$\penalty 10000\ \hskip 44.10185pt-\frac{287}{3\,888}+\frac{287}{23\,328}\frac{r_{s}}{r}-\frac{287}{139\,968}\frac{r_{s}^{2}}{r^{2}}+\frac{287}{839\,808}\frac{r_{s}^{3}}{r^{3}}-\frac{287}{5\,038\,848}\frac{r_{s}^{4}}{r^{4}}\bigg)$
-$\penalty 10000\ \hskip 28.45274pt+Ci\omega r_{s}\bigg[\frac{1\,246}{15}\frac{r^{5}}{r_{s}^{5}}-\frac{10\,591}{90}\frac{r^{4}}{r_{s}^{4}}+\frac{15\,931}{540}\frac{r^{3}}{r_{s}^{3}}+\frac{32\,129}{3\,240}\frac{r^{2}}{r_{s}^{2}}-\frac{32\,129}{19\,440}\frac{r}{r_{s}}$
-[…]
-$\penalty 10000\ \hskip 28.45274pt+C\omega^{2}r_{s}^{2}\bigg[-\frac{14}{11}\frac{r^{7}}{r_{s}^{7}}-\frac{931}{330}\frac{r^{6}}{r_{s}^{6}}-\frac{r^{5}}{r_{s}^{5}}\!\left(\frac{14\pi^{2}}{3}+\frac{2\,798\,434}{17\,325}-\frac{3\,142}{495}\log\frac{r_{s}}{r}\right)$
-$\penalty 10000\ \hskip 78.24507pt+\frac{r^{4}}{r_{s}^{4}}\!\left(\frac{119\pi^{2}}{18}+\frac{202\,591\,567}{831\,600}-\frac{26\,707}{2\,970}\log\frac{r_{s}}{r}\right)$
-$\penalty 10000\ \hskip 78.24507pt-\frac{r^{3}}{r_{s}^{3}}\!\left(\frac{179\pi^{2}}{108}+\frac{332\,182\,607}{4\,989\,600}-\frac{281\,209}{124\,740}\log\frac{r_{s}}{r})\right)$
-$\penalty 10000\ \hskip 78.24507pt-\frac{r^{2}}{r_{s}^{2}}\!\left(\frac{361\pi^{2}}{648}+\frac{603\,088\,753}{29\,937\,600}-\frac{567\,131}{748\,440}\log\frac{r_{s}}{r}\right)$
+$\penalty 10000\ \hskip 2.84544pt+C\omega^{2}r_{s}^{2}\Bigg[-\frac{14}{11}\frac{r^{7}}{r_{s}^{7}}-\frac{931}{330}\frac{r^{6}}{r_{s}^{6}}-\frac{r^{5}}{r_{s}^{5}}\!\left(\frac{14\pi^{2}}{3}+\frac{2\,798\,434}{17\,325}-\frac{3\,142}{495}\log\frac{r_{s}}{r}\right)$
 […]